### v1

# Online Optimization Algorithms in Repeated Price Competition: Equilibrium Learning and Algorithmic Collusion

Abstract
This paper addresses the question of whether or not uncoupled online learning algorithms converge to the Nash equilibrium in pricing competition or whether they can learn to collude. Algorithmic collusion has been debated among competition regulators, and it is a highly relevant phenomenon for buyers and sellers on online retail platforms. We analyze formally if mean-based algorithms, a class of bandit algorithms relevant to algorithmic pricing, converge to the Nash equilibrium in repeated Bertrand oligopolies. Bandit algorithms only learn the profit of the agent for the price set in each step. In addition, we provide results of extensive experiments with different types of multi-armed bandit algorithms used for algorithmic pricing. In a mathematical proof, we show that mean-based algorithms converge to correlated rational strategy profiles, which coincide with the Nash equilibrium in versions of the Bertrand competition. Learning algorithms do not converge to a Nash equilibrium in general, and the fact that Bertrand pricing games are learnable with bandit algorithms is remarkable. Our numerical results suggest that wide-spread bandit algorithms that are not mean-based also converge to equilibrium and that algorithmic collusion only arises with symmetric implementations of UCB or Q-learning, but not if different algorithms are used by sellers. In addition, the level of supra-competitive prices decreases with increasing numbers of sellers. Supra-competitive prices decrease consumer welfare. If algorithms lead to algorithmic collusion, this is important for consumers, sellers, and regulators to understand. We show that for the important class of multi-armed bandit algorithms such fears are overrated unless all sellers agree on a symmetric implementation of certain collusive algorithms. .
Keywords price competition, algorithmic collusion, bandit algorithms, rationalizability
1 Introduction
Algorithmic pricing, where prices are determined by software agents, is becoming increasingly prevalent in online retail markets. Chen et al. (2016) estimated for 2015 that algorithms were involved in the pricing of approximately one-third of the roughly 1600 best-selling products on the Amazon marketplace. In 2018, the average product price on Amazon was reported to change once every ten minutes. [^1] https://www.businessinsider.com/amazon-price-changes-2018-8 Since then, an industry has emerged that specializes in automated pricing software with over 130 companies offering a variety of 185 pricing algorithms in 2021 (Calzolari and Hanspach 2024).
From the point of view of a single seller, algorithmic pricing aims to solve an online optimization problem, where a firm’s actions are the prices it sets, and the objective is the profit it wants to maximize over time. When entering the market, sellers have little information about competitors’ costs or the demand of buyers at different prices and need to find out over time which prices maximize profit.
A key challenge is deciding whether to prioritize short-term revenue (by exploiting a known price with a high payoff) or invest in discovering better long-term pricing strategies (through exploration of different prices). Online optimization algorithms are designed to effectively balance this trade-off, and they scale effectively across large sets of products (Bubeck 2011). On online exchanges, such optimization algorithms have access to bandit feedback. This means that after performing an action (setting a price), an agent learns the value of the objective function value (his profit) for this specific action. This (multi-armed) bandit model in online optimization is particularly suitable for algorithmic pricing applications, which was recognized early on. Bandit algorithms were already suggested for pricing by Rothschild (1974) long before digital platform markets emerged. Today, there is an extensive academic literature on multi-armed bandit algorithms for pricing (Trovo et al. 2015, den Boer 2015, Bauer and Jannach 2018, Mueller et al. 2019, Elreedy et al. 2021, Taywade et al. 2023, Qu 2024) and there are numerous resources by practitioners on how to implement bandit algorithms for pricing. [^2] https://towardsdatascience.com/dynamic-pricing-with-multi-armed-bandit-learning
-by-doing-3e4550ed02ac
, https://www.griddynamics.com/blog/dynamic-pricing-algorithms
Online optimization algorithms target online problems where agents optimize against a stochastic process that is unknown and independent of their actions. Game-theoretical problems are different because the actions of one player impact the objective of the others. In games, the Nash equilibrium (NE) describes a state where no agent wants to deviate unilaterally. Unlike the optimum in single-player optimization, this state may bring suboptimal payoffs to all agents. In spite of a long literature on learning in games, we don’t have a comprehensive theory about which algorithms converge to a Nash equilibrium in which types of games (Fudenberg and Levine 1999, Young 2010, Cesa-Bianchi and Lugosi 2006).
Algorithmic pricing is a game-theoretical setting. More specifically, an online retail market where multiple sellers compete via prices in a market for homogeneous goods is naturally modeled as a Bertrand competition (Bertrand 1883a). This established oligopoly model allows for different assumptions about consumer demand (e.g., all-or-nothing, linear, or logit demand models) and there is a long literature characterizing the prices that emerge in equilibrium. In a Bertrand competition with homogeneous products and all-or-nothing demand, the equilibrium price equals marginal cost, which leads to maximization of consumer welfare.
Equilibrium analysis assumes rational bidders that play their equilibrium strategy from the start. In practice, however, competitors lack information about each other’s costs and the demand model, which are both needed to calculate such a price in advance. Additionally, changes in demand and supply over time would require players to recalculate their equilibrium strategy.
Unfortunately, a growing number of articles showed that optimization and learning algorithms can lead to supra-competitive prices higher than the Nash equilibrium of the stage game Waltman and Kaymak (2008), Calvano et al. (2019), Abada and Lambin (2023), Klein (2021), Abada et al. (2024b), Brown and MacKay (2023). [^3] Note that by the Folk Theorem, such a supra-competitive price could still be an equilibrium in an infinitely repeated game, if players have sufficient patience (Fudenberg and Tirole 1991). The phenomenon is commonly referred to as algorithmic collusion (den Boer 2023), and it raised substantial concerns among regulators (OECD 2017) as it decreases consumer welfare. Abada et al. (2024a) provide an up-do-date treatment of the subject, an overview of relevant literature, and policy implications. They define algorithmic collusion as persistent supra-competitive outcomes produced by learning algorithms without human design to produce those outcomes, and we adhere to this definition.
The discussion on algorithmic collusion draws largely on numerical experiments of certain learning algorithms in a repeated Bertrand competition with a fixed set of sellers and a specific demand model. While real-world environments might be more complex, this model enables an analysis of convergence to the static Nash equilibrium. In particular, analysts can perform comparative statics based on the knowledge of costs and demand functions that are not available in empirical work. One might argue that if algorithmic collusion does not arise in repeated play in a Bertrand pricing game, it is unlikely to emerge in more complex environments where demand and supply change over time. There are a number of environments where algorithmic collusion was shown experimentally. Most articles focus on Q-learning (Calvano et al. 2020a, 2021, den Boer et al. 2022, Klein 2021). But Hansen et al. (2021) recently found evidence for supra-competitive prices with Upper Confidence Bound (UCB, Auer et al. (2002)), a multi-armed bandit algorithm, as well. Reward-punishment schemes that might be the result of reinforcement learning algorithms with large state spaces can be ruled out in these environments (Lambin 2024). We argue that it is important to study not only properties of algorithms but also the games that possibly lead to algorithmic collusion.
Decades of research on learning in games (Young 2010, Foster and Vohra 1997) has shown that convergence of learning algorithms to a Nash equilibrium typically requires strong assumptions. It is well-known that there are games that cannot be learned via any uncoupled learning dynamics (Hart and Mas-Colell 2006, Milionis et al. 2022). Also, random normal-form games can lead to cycles or even chaotic dynamics (Sanders et al. 2018). The Bertrand competition has more structure that we exploit in this paper to show the convergence of large classes of algorithms relevant to algorithmic pricing. The utility functions of the sellers are symmetric, and they all face the same trade-off between high payoff and high market share. It turns out that the resulting structure of these games is such that uncoupled learning dynamics do converge to equilibrium.
1.1 Contributions
Prior studies on algorithmic pricing and algorithmic collusion analyzed specific algorithms in a repeated Bertrand competition with specific demand models. We make two main contributions: First, we prove that an important class of online optimization algorithms, mean-based algorithms, converge to correlated rationalizable strategies (Brandenburger and Dekel 1987), a solution concept that contains correlated equilibria (Aumann 1987) and Nash equilibria. Then, we show that in symmetric Bertrand competition with all-or-nothing or linear demand and discrete actions, mean-based algorithms converge almost surely to actions that are close to their Nash equilibria. This is because the set of correlated rationalizable strategies is equivalent or close to the set of Nash equilibria of these games. The connection between mean-based algorithms and correlated rationalizable strategies was unknown and the finding adds to the literature on learning in games. The result has real-world relevance because it provides proof that widely used algorithms such as Exponential Weights (Exp3) converge to a Nash equilibrium in important models of pricing competition. Note that such convergence results of bandit algorithms to Nash equilibrium are rare in the literature on learning in games.
Not all online optimization algorithms suitable for algorithmic pricing are mean-based. In a second contribution, we provide extensive experiments with a variety of multi-armed bandit algorithms that have been suggested for algorithmic pricing and show that algorithmic collusion is an exception limited to symmetric installations of UCB algorithms and Q-learning among a small number of sellers. We report experiments on Exp3, $\varepsilon$-greedy, UCB, and Thompson sampling, showing that UCB leads to high prices while the other algorithms converge to the Nash equilibrium. This holds for all-or-nothing, linear, and logit demand models for symmetric and asymmetric model parameterizations. In experiments in which two or more different algorithms (including UCB) are combined, we find that prices converge to the Nash equilibrium quickly. Algorithmic collusion with UCB is also much reduced if there are more than three sellers. Overall, these results indicate that for the important class of online optimization algorithms, algorithmic collusion is less of a concern.
1.2 Organization of the Paper
In the next section, we discuss related work on online optimization, algorithmic collusion, and learning in games. In Section 3, we introduce the relevant game-theoretical solution concepts and the Bertrand competition. Section 4 provides the central theoretical result of our paper on the convergence of mean-based algorithms to equilibrium, before we describe our experimental results in Section 5. Section 6 provides conclusions and an outlook.
2 Related Work
In what follows, we provide an overview of the relevant literature. We briefly discuss online optimization algorithms, summarize the literature on algorithmic collusion, and the key results from the literature on learning in games relevant to this paper.
2.1 Online Optimization
Online optimization is concerned with making sequential decisions in an unknown environment with the goal of optimizing a performance metric over time. Let us briefly introduce the basic model of online optimization. At each stage $t=1,2,\dots$ the agent chooses a strategy $x_{t}\in\mathcal{X}$ from some set and gets a utility $u_{t}(x_{t})$. In algorithmic pricing, this action could be the price or the quantity. It is usually assumed that $\mathcal{X}$ is a closed convex subset of some vector space and $u_{t}$ are convex functions (Shalev-Shwartz 2011). In what follows, we are going to focus on the learning algorithms with performance guarantees.
Algorithms are commonly analyzed in two models, the adversarial model and the stochastic model. In the stochastic model, the objective goal is to minimize the expected regret over the distribution of the input, which is drawn independently and identically from some underlying distribution. In the adversarial model, the input can be chosen by an adversary. A standard performance measure of algorithms generating a sequence of strategies $x_{t}$ in both models is the notion of (external) regret.
Definition 2.1 (No-regret Algorithm).
The (external) regret denotes the difference between the aggregated utilities after $T$ stages and the best fixed action in hindsight and is defined by
$\text{Reg}(T)=\max_{x\in\mathcal{X}}\sum_{t=1}^{T}u_{t}(x)-u_{t}(x_{t}).$ (1)
We say that an algorithm has no regret if the regret $\text{Reg}(T)$ grows sublinearly. This means that an algorithm performs at least as well as the best fixed action in hindsight.
There are different types of no-regret online algorithms. One can distinguish such algorithms based on the feedback available to the agents: zeroth-order (i.e., bandit) or first-order (i.e., gradient-based) algorithms.
  • Bandit feedback: The algorithm only receives partial feedback about the performance of its decisions. Typically, the feedback consists of a scalar reward or cost signal associated with the chosen action, but it does not reveal the potential rewards or costs of the other actions. Real-world implementations of learning agents are best modeled with bandit oracles. For example, after setting a price in an oligopoly competition, an agent observes a specific reward in the next period.
  • Gradient feedback: The algorithm receives feedback about the first-order derivatives (e.g., gradients) of the cost or payoff function with respect to the chosen action. Algorithms that have access to first-order oracles are often more efficient and allow for the implementation of fast algorithms. While such feedback is typically not available to learning algorithms in the field, one can compute gradient feedback when the goal is to develop equilibrium solvers rather than to mimic real-world interactions.
Apart from these feedback or oracle models, there can also be full information feedback, where agents have access to the reward of all their possible actions, or higher-order feedback. We will largely focus on bandit feedback, which is a natural algorithm design principle for algorithmic pricing. Sellers set a price, and they learn their profit for this price in the next period. They often have no or only incomplete information about the demand model or the strategies used by all other sellers. This is particularly true on large online retail platforms where there are many substitutes for a good. Note that the notion of no regret does not make any assumptions about the feedback or the distribution of the utilities.
Exponential Weights, Follow-the-Perturbed-Leader (FTPL), Online Gradient Descent, or Online Mirror Descent are all algorithms that satisfy the no-regret property. Exponential Weights (or the Exp3 variant) uses bandit feedback and updates the weights associated with each action based on the cumulative reward observed for that action. The algorithm then chooses actions with probabilities proportional to their weights.
Exp3 is also a mean-based algorithm (Braverman et al. 2018), a property of online optimization algorithms that will play an important role in our analysis. Informally, if the mean reward of action $a$ is significantly larger than the mean reward of action $b$, the learning algorithm will choose action $b$ with negligible probability. Apart from Exp3, also Multiplicative Weights Update (MWU), and FTPL are mean-based. We will show that such algorithms converge to a Nash equilibrium in versions of the Bertrand competition.
There are academic articles using (deep) reinforcement learning algorithms for algorithmic pricing (Rana and Oliveira 2014, Kastius and Schlosser 2022, Deng et al. 2024). Such algorithms allow the agent to take into account the state of a system. This state could be historical prices, the day of the week, or other variables that are potentially relevant. With an increasing state and action space, learning requires a lot of training rounds and doesn’t scale well. Compared to online learning algorithms without states, reinforcement learning algorithms are sample-inefficient, which makes them less appropriate for dynamic pricing, where agents need to adapt rapidly to changing market conditions. Q-learning is a reinforcement learning technique that was used in the literature on algorithmic collusion that we discuss next. However, the state space is often limited to the current price such that it resembles an $\varepsilon$-greedy online optimization algorithm. Lambin (2024) shows that algorithmic collusion can arise with Q-learning even in a stateless version of Q-learning. Based on these arguments and discussions with practitioners, we focus on online optimization algorithms with bandit feedback in our work. However, we report versions of Q-learning with restricted state space in the online supplement B.2, providing a comparison to prior research using this algorithm.
2.2 Algorithmic Collusion
Algorithmic collusion has drawn substantial attention from the research community and from policymakers, as we outlined in the introduction. Most of this literature analyzes specific algorithms such as Q-learning for specific model variations, i.e., Bertrand oligopolies with standard all-or-nothing demand, with linear, or logit demand (Bertrand 1883a). Calvano et al. (2020b) analyze a Bertrand competition with logit demand and constant marginal cost. They find that when playing against copies of a Q-learning agent, the competition of these agents can lead to supra-competitive prices higher than the Nash equilibrium. In line with the literature, we refer to such supra-competitive prices as algorithmic collusion. A related sequential move pricing duopoly environment with linear demand (instead of the simultaneous move Bertrand model in Calvano et al. (2020b)) was analyzed by Klein (2021), who also found collusion with Q-learning agents. Asker et al. (2024) detect in their experiments on Bertrand competition with standard (all-or-nothing) demand that collusion depends on specifics of the Q-learning algorithm (e.g., synchronous vs. asynchronous updating). While there are measures that can be implemented against collusion, e.g., on a platform-level (Johnson et al. 2023), these articles agree on the potential risk to consumer welfare posed by algorithmic sellers.
In contrast, Abada et al. (2024b) analyze Q-learning in Bertrand oligopolies and show that Q-learning algorithms with sufficiently large $\epsilon$-greedy exploration show no collusion. den Boer et al. (2022) provide a detailed analysis of the inner workings of Q-learning and reveal that there is no immediate reason to believe that Q-learning would lead to collusion easily. In addition, Eschenbaum et al. (2022) criticize the claim that algorithms can be trained offline to successfully collude online in different market environments. The authors find that collusion breaks down when collusive reinforcement learning policies are extrapolated from a training environment to the market.
Most of this experimental literature on algorithmic collusion is based on Q-learning, although there is no evidence that this algorithm is particularly important or widespread for algorithmic pricing. More recently, Hansen et al. (2021) analyzes the price levels that arise in a duopoly setting where UCB ("Upper Confidence Bound") agents determine prices. They run a series of experiments where a variant of symmetric UCB algorithms interact simultaneously in a Bertrand economy competition with linear demand functions. The agents observe a perturbed estimate of their revenues which are a result of their prices and the corresponding demand. Hansen et al. (2021) find that, under sufficiently small reward noise, agents eventually explore prices in a correlated manner, giving rise to supra-competitive outcomes. [^4] Note that in this literature review, we focus on algorithms that were not designed to collude (Meylahn and V. den Boer 2022). A more extensive discussion of the literature is provided by Abada et al. (2024a). They find that in the settings studied in the literature, sets that include Q-learning and non-Q-learning algorithms do not seem to constitute persistent sets of supracompetitive algorithms. We provide a rationale for this by analyzing the properties of these games and analyze a variety of bandit algorithms that have been reported for algorithmic pricing.
In summary, the literature on algorithmic collusion largely focuses on specific algorithms (mostly Q-learning) that are employed in specific types of oligopoly models. The same type of algorithm is used in a symmetric model. We analyze a variety of multi-armed bandit algorithms competing against the same but also against different algorithms. We analyze standard all-or-nothing demand, linear, and logit demand models in a Bertrand competition.
2.3 Learning in Games
The literature on algorithmic collusion is inherently tied to that on learning in games. The latter has a long history and asks what type of equilibrium behavior (if any) may arise in the long run of a process of learning and adaptation, in which agents are trying to maximize their payoff while learning about the actions of other agents through repeated interactions (Fudenberg and Levine 1999). Cournot’s study of duopoly (Cournot 1838) already introduced a Nash equilibrium and a particular learning process. Cournot’s model of best reply dynamics appears unrealistic as a model for algorithmic pricing, because players would need a lot of information about competitors. Later, Brown (1951) suggested fictitious play, which converges to equilibrium distributions for any two-player, finite-strategy, zero-sum game. However, Shapley (1964) established that it can lead to cycles and that the frequency distribution does not necessarily converge. Many learning algorithms have been developed, ranging from iterative best-response to first-order online optimization algorithms in which agents follow their utility gradient in each step (Mertikopoulos and Zhou 2019, Bichler et al. 2023).
In this general context, it is well known that the dynamics of learning agents do not always converge to a Nash equilibrium (Milionis et al. 2022, Vlatakis-Gkaragkounis et al. 2020): they may cycle, diverge, or be chaotic, even in zero-sum games, where the Nash equilibrium is tractable (Mertikopoulos et al. 2018, Bailey and Piliouras 2018). While there is no comprehensive characterization of games that are “learnable” and one cannot expect that uncoupled dynamics lead to Nash equilibrium in all games (Hart and Mas-Colell 2003, Milionis et al. 2022), there are some important results regarding learners. A classical result is that the class of no-regret learning algorithms converges to the so-called [coarse] correlated equilibrium ([C]CE) of a game Fudenberg and Levine (1999). [C]CEs are superclasses of Nash equilibria. However, CCEs can also contain dominated strategies and are a rather weak solution concept.
Less is known about conditions of games in which learning algorithms converge to a Nash equilibrium. In a landmark paper, Monderer and Shapley (1996) introduced the class of potential games and they showed that Cournot oligopolies with linear price or cost functions are potential games. Potential games are guaranteed to have at least one pure Nash equilibrium. Importantly, it was shown that algorithms such as best and better response dynamics or fictitious play converge to a Nash equilibrium in these games. More recently, convergence of MWU (Palaiopanos et al. 2017) and Exp3 (Cohen et al. 2017) in potential games was also shown.
Another important property of a game is that of a game being supermodular (Vives 2001, Topkis 1998, Milgrom and Roberts 1990). In such games, the best response of each player is positively correlated with the strategies chosen by other players. Milgrom and Roberts (1990) showed that Bertrand oligopoly games are log-supermodular if each firm’s elasticity of demand is a decreasing function of its competitors’ prices. Later, Milgrom and Roberts (1991) established that adaptive learning algorithms converge to the unique NE in a Cournot duopoly model and Bertrand oligopoly models with linear and logit demand. Adaptive learning algorithms include best-response dynamics and fictitious play. In the bandit algorithms that we analyze, an agent can only observe his profit for a particular action, but not necessarily the strategies of all other players as is necessary for these earlier equilibrium finding algorithms.
3 Model
In the following, we will first introduce the necessary notation and definitions before we define different types of Bertrand competition models.
3.1 Basic Definitions and Notation
We begin with the concepts of a finite normal-form game, a Nash equilibrium, and a symmetric game. A normal-form game is a representation in game theory that defines the strategies available to each player, their corresponding payoffs, and the resulting outcomes in a simultaneous and strategic interaction. Formally, we have the following definition.
Definition 3.1 ((Finite) Normal-form Game).
A normal-form game with $n$ players can be described as a tuple $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$, with a finite set of players $i\in\mathcal{N}=\{1,\dots,n\}$, a action space $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ for all $i\in\mathcal{N}$, and payoff or utility functions $u=(u_{1},\dots,u_{n})$ with $u_{i}:\mathcal{A}\rightarrow\mathbb{R}$. We call the game a finite normal-form game if the action spaces are also finite, i.e., $|\mathcal{A}_{i}|<\infty$.
All agents $i=1,\dots,n$ submit their actions $a_{i}$ to form an action profile $a=(a_{1},\dots,a_{n})$. We often abbreviate the profile by $(a_{i},a_{-i})$ where $a_{-i}=(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n})$. Similarly, we will sometimes index parts of combined spaces as follows: $\mathcal{A}_{-i}=\mathcal{A}_{1}\times\dots\times\mathcal{A}_{i-1}\times\mathcal{A}_{i+1}\times\dots\times\mathcal{A}_{n}$. The agents may also play a distribution over actions, also known as a mixed strategy, denoted by $x_{i}\in\Delta(\mathcal{A}_{i})$, where $\Delta(\mathcal{A}_{i})$ is the probability simplex over $\mathcal{A}_{i}$. For any joint distribution $x\in\Delta(\mathcal{A})$ over action tuples, the expected utility of player $i$ is $u_{i}(x)=\operatorname{\mathbb{E}}_{a\sim x}\left[u_{i}(a_{i},a_{-i})\right]$.
For many of our theoretical results, we assume that the game is symmetric, meaning that every player is identical with respect to the game rules. Textbook versions of the Bertrand competition assume symmetry and this allows for clean results.
Definition 3.2 (Symmetric Game).
A normal-form game is symmetric if all players have identical action spaces, i.e., $\mathcal{A}_{1}=\dots=\mathcal{A}_{n}$, and anonymous utility functions in the sense that $u_{i}(a_{i},a_{-i})=u_{j}(a_{j},a_{-j})$ for $a_{i}=a_{j}$ and $a_{-i}=a_{-j}$ for all $i,j\in\mathcal{N}$.
A Nash equilibrium is a situation in a strategic interaction where each player’s strategy is optimal given the strategies chosen by all other players, and no player has an incentive to unilaterally deviate from their chosen strategy.
Definition 3.3 (Nash Equilibrium (NE)).
In a normal-form game $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$, a strategy profile $x^{*}=(x_{1}^{*},\ldots,x_{n}^{*})$ is a Nash equilibrium if, for every player $i\in\mathcal{N}$, we have:
$u_{i}(x_{i}^{*},x_{-i}^{*})\geq u_{i}(x_{i},x_{-i}^{*}),\quad\forall x_{i}\in\Delta\mathcal{A}_{i}.$ (2)
Note that every symmetric game has a symmetric Nash equilibrium in mixed strategies (Cheng et al. 2004). It is also well-known that computing Nash equilibria is PPAD-hard in the worst case (Daskalakis et al. 2009).
There are specific subsets of normal-form games where the utilities have more structure. This allows us to find an equilibrium with faster iterative algorithms. Potential games (Monderer and Shapley 1996) and supermodular games (Topkis 1979, Milgrom and Roberts 1990) have received significant attention.
Definition 3.4 (Potential Game).
A game $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$ is a potential game if there exists a potential function $\phi:\mathcal{A}\rightarrow\mathbb{R}$ such that for every player $i\in\mathcal{N}$ and every pair of actions $a_{i},a^{\prime}_{i}\in\mathcal{A}_{i}$
$u_{i}(a_{i},a_{-i})-u_{i}(a^{\prime}_{i},a_{-i})=\phi(a_{i},a_{-i})-\phi(a^{\prime}_{i},a_{-i}),\quad\forall a_{-i}\in\mathcal{A}_{-i}.$ (3)
It is known that best response dynamics, better response dynamics, Fictitious Play, Replicator Dynamics, and simultaneous gradient ascent all converge in potential games (Fudenberg and Levine 1999, Monderer and Shapley 1996, Sandholm 2010, Swenson et al. 2018). Cohen et al. (2017) could also show convergence of bandit algorithms such as Exp3 to a Nash equilibrium in potential games. Exp3 is a bandit algorithm relevant to algorithmic pricing.
Bertrand competitions with some demand models are supermodular games, which were introduced by Topkis (1979). Supermodular games are defined on complete lattices and require supermodular utility functions (see, for instance, (Milgrom and Roberts 1990)). Since we only consider one-dimensional action spaces $\mathcal{A}\subset\mathbb{R}$, the utility functions trivially satisfy the supermodularity condition, and we can use the componentwise ordering to define a complete lattice on $\mathcal{A}_{i}$ and $\mathcal{A}_{-i}$. Then, the definition of supermodular games reduces to the following.
Definition 3.5 (Supermodular Game).
A normal-form game $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$ with $\mathcal{A}_{i}\subset\mathbb{R}$ is called a supermodular game if it satisfies the following conditions for each player $i\in\mathcal{N}$:
  1. (1)
    $u_{i}$ is order upper semi-continuous in $a_{i}$ and continuous in $x_{-i}$ and has a finite upper bound
  2. (2)
    $u_{i}$ has increasing differences in $a_{i}$ and $a_{-i}$, i.e., for any two action profiles $a=(a_{i},a_{-i})$ and $a^{\prime}=(a_{i}^{\prime},a_{-i}^{\prime})$ such that $a_{i}^{\prime}\geq a_{i}$ and $a_{j}^{\prime}\geq a_{j}$for all $j\neq i$, the following inequality holds:
    $u_{i}(a_{i}^{\prime},a_{-i}^{\prime})-u_{i}(a_{i},a_{-i}^{\prime})\geq u_{i}(a_{i}^{\prime},a_{-i})-u_{i}(a_{i},a_{-i})$ (4)
If we restrict ourselves to finite normal-form games, i.e., discrete actions, then we only have to check the increasing differences property (2) and can ignore the continuity requirements (1).
It is known that pure strategy Nash equilibria exist in supermodular games. The set of actions surviving iterated strict dominance has a greatest and least element, and both are Nash equilibria (Levin 2006). Consequently, if a game has a unique Nash equilibrium, then strict elimination of dominated strategies finds this equilibrium. However, the elimination of dominated strategies is not a learning method that can be used independently by all players of the game because it requires complete information about the payoff matrix of a game. Such information is not available in algorithmic pricing. Therefore, we will focus on uncoupled algorithms that can be used independently by the agents and that only learn bandit feedback after each round.
3.2 Bertrand Competition
The Bertrand pricing game or Bertrand competition (Bertrand 1883b) is an economic model that describes the revenue of a firm according to the price it sets for its product and the resulting demand. The firms’ action $a_{i}\in\mathcal{A}_{i}$ is the price for a good it wants to produce. All firms produce the same good, and they all take action simultaneously. Firms compete for demand with their prices and affect each others’ revenues: The demand depends on all agents’ actions and is decreasing in the agent’s own price. As is common, we assume that the cost functions are linear in the demand. In summary, the utility can be described by
$u_{i}(a_{i},a_{-i})=d_{i}(a)\cdot(a_{i}-c_{i}).$ (5)
3.2.1 Demand Models
We consider different demand models. The standard or all-or-nothing demand (also called Bertrand demand) is similar to an auction, where only the highest bidder gets the item. In this case, the firm with the lowest price gets all the demand. Different from auctions where the valuation of an item does not depend on its price, the demand in this competition is additionally linearly decreasing in the agent’s price. If multiple firms offer the lowest price, the demand is shared equally. This leads to a demand function given by
$d_{i}(a_{i},a_{-i})=\begin{cases}\frac{D}{n_{min}}(1-a_{i})&\text{if }i\in\arg\min_{j\in\mathcal{N}}a_{j}\\ 0&\text{else}\end{cases},$ (standard)
where $D>0$ is the maximum total demand, and $n_{min}:=\lvert\arg\min_{j\in\mathcal{N}}a_{j}\rvert$ is the number of firms with the lowest price.
Another demand model that was used in Calvano et al. (2020b), is the multinomial or logit demand. In this setting, the demand is split between the $n$ agents/goods and some outside good (indexed with 00) according to
$d_{i}(a_{i},a_{-i})=\dfrac{\exp\left(\frac{\alpha_{i}-a_{i}}{\mu}\right)}{\exp\left(\frac{\alpha_{0}}{\mu}\right)+\sum_{j=1}^{n}\exp\left(\frac{\alpha_{j}-a_{j}}{\mu}\right)}.$ (logit)
The parameters $\alpha_{i}>0$ capture different product quality indices and $\mu>0$ models a product differentiation between the goods, i.e., if $\mu\rightarrow 0$, the goods are perfect substitutes.
And finally, we also introduce the linear demand that was considered in Hansen et al. (2021). The agents’ demand function is given by
$d_{i}(a_{i},a_{-i})=\alpha_{i}-\beta_{i}a_{i}+\tfrac{\gamma}{n-1}\sum_{j\neq i}a_{j}.$ (linear)
Similar to the standard model, the demand decreases with rate $\beta_{i}>0$ in the price of the agent, but it also increases with rate $\gamma>0$ in the sum of the prices of all other agents. We assume that the effect of the own price is greater than the effect of the average opponents’ prices, i.e., $\beta_{i}>\gamma$.
In a standard Bertrand model with homogeneous products and symmetric firms, prices equal marginal costs in equilibrium. However, this changes with asymmetries or product differentiation. With asymmetric costs between firms, the Nash equilibrium typically involves the low-cost firm pricing at or just below the marginal cost of the high-cost firm. With product differentiation (as in logit demand), equilibrium prices are typically above marginal costs due to firms having some market power. Independent of the demand model, we will use the Nash equilibrium as a baseline for the learning outcome.
3.2.2 Game-Theoretical Properties
Let us now discuss some properties that are already known about Bertrand pricing models. The referenced propositions can be found in the appendix (Section A.1). Milgrom and Roberts (1991) showed that Bertrand competitions with linear and logistic demand and continuous actions are supermodular games. As indicated earlier, iterated elimination of dominated strategies converges to a unique Nash equilibrium in this model. However, this is not the case for standard all-or-nothing demand, as we show (see Proposition A.1).
Potential games have even stronger properties, as we have seen. Apart from best- and better-response dynamics, and fictitious play, the class of no-regret learning converges to equilibrium in potential games. We show that the Bertrand competition with a linear demand function is not only supermodular, it is also a potential game (see Proposition A.2). For Bertrand competitions with standard all-or-nothing and logit demand, we can show that these games are, in general, no potential games (see Proposition A.3 and A.4). An overview of these concepts for the different utility models is given in Table 1.
3.3 Alternative Solution Concepts
We explore alternative solution concepts to the Nash equilibrium. While the latter is computationally hard to find, there exist simple iterative algorithms that can identify strategy profiles corresponding to these solution concepts. In some games, the alternatives coincide with the Nash equilibrium, which is what we will leverage in our convergence proof for mean-based algorithms.
3.3.1 (Coarse) Correlated Equilibria
Apart from Nash equilibria (NE), Correlated Equilibria (CE) and Coarse Correlated Equilibria (CCE) have received significant attention in the literature on learning in games. It is well known that $NE\subseteq CE\subseteq CCE$.
Definition 3.6 (Correlated Equilibrium).
A joint distribution $\sigma\in\Delta(\mathcal{A})$ is a correlated equilibrium if for every player $i\in\mathcal{N}$ and for all actions $a_{i},a_{i}^{\prime}\in\mathcal{A}_{i}$,
$\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i},a_{-i})\mid a_{i}]\geq\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i}^{\prime},a_{-i})\mid a_{i}].$
Definition 3.7 (Coarse Correlated Equilibrium).
A joint distribution $\sigma\in\Delta(\mathcal{A})$ is a coarse correlated equilibrium if for every player $i\in\mathcal{N}$ and for all actions $a_{i}^{\prime}\in\mathcal{A}_{i}$,
$\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i},a_{-i})]\geq\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i}^{\prime},a_{-i})].$
The CCE constraints form a polytope, and therefore individual CCEs can be computed via a linear program. It was shown that algorithms with no internal regret converge to a CE and those with no external regret to a CCE (Foster and Vohra 1997). Popular bandit algorithms such as Exp3 are no-external-regret algorithms. However, CCEs are a weak solution concept, and they can contain dominated strategies (Viossat and Zapechelnyuk 2013). Also, the set of CCEs can be a very large superset of the set of Nash equilibria. In Example 3.1, we computed different CCEs for Bertrand competition models with linear and all-or-nothing demand.
Example 3.1 (Coarse Correlated Equilibria in Bertrand Competitions).
We consider Bertrand competitions with 2 players, action spaces $\mathcal{A}=\{0.1,0.2,\dots,0.9\}$, cost parameters $c_{1}=c_{2}=0$, and two different demand functions, namely linear demand ($\alpha_{i}=0.48,\,\beta_{i}=0.9,\,\gamma=0.6$ for $i\in\mathcal{N}$) and all-or-nothing demand ($D=1$). The computed CCEs are visualized in Figure 1.
(a) Linear Demand
(b) All-or-Nothing Demand
Figure 1: Coarse Correlated Equilibria for Different Bertrand Competitions.
In the figures, we visualize exemplary coarse correlated equilibria and highlight the Nash equilibria (boxes with black frames). The CCEs were computed by solving a linear program (LP), where the constraints come from the definition of a CCE, and the objective was chosen such that the Wasserstein distance to the Nash equilibria is maximized. This approach places more weight on action profiles further away from the equilibrium profiles.
For the linear demand model, the CCEs contain only actions near the Nash equilibrium. However, for the all-or-nothing demand model, we can find CCEs that contain dominated actions. This is why the guarantee that no-regret learners converge to a CCE is not sufficient in a Bertrand competition with all-or-nothing demand.
3.3.2 Serially Undominated Set
Another important solution concept that can be computed by means of a simple, iterative algorithm is that of a serially undominated set. For this, let us introduce different types of dominance relationships.
Definition 3.8 (Dominated Actions).
An action $a_{i}\in\mathcal{A}_{i}$ is
  • strongly dominated, if there exists another action $\hat{a}\in\mathcal{A}_{i}$ such that $u_{i}(\hat{a},a_{-i})>u_{i}(a_{i},a_{-i})$
  • strictly dominated, if there exists a mixed strategy $\hat{x}\in\Delta(\mathcal{A}_{i})$ such that $u_{i}(\hat{x},a_{-i})>u_{i}(a_{i},a_{-i})$
for all opponents’ action profiles $a_{-i}\in\mathcal{A}_{-i}$.
If we iteratively remove dominated actions, we end up with a serially undominated set. From (Milgrom and Roberts 1990), we know that, in supermodular games, the iterative procedure of removing strongly dominated strategies leads to such a serially undominated set where the maximum and minimum elements are component strategies of pure Nash equilibria. This set includes all Nash equilibria, correlated equilibria, and rationalizable actions (Bernheim 1984, Pearce 1984).
Iteratively crossing out strictly dominated actions, also known as Iterated Strict Dominance (ISD), leads to a smaller set of actions, which we will call strictly serially undominated set (SSU). A game is said to be strict dominance solvable if ISD yields a unique strategy profile. In finite games, this strategy profile is necessarily the unique Nash equilibrium (Fudenberg and Tirole 1991, Kolpin 2009). Note that in order to perform ISD, we need complete information about the payoff matrix. This assumption is too strong for algorithmic pricing, which is why we will focus on bandit algorithms later. In finite games, there is a close relationship between actions that survive ISD and correlated rationalizable actions, which we will explore in the following subsection.
3.3.3 Correlated Rationalizable Set
Our main convergence statement is centered on the concept of correlated rationalizability (Brandenburger and Dekel 1987). These are actions that are a best response against any joint mixed strategy of the opponents that is also correlated rationalizable.
Definition 3.9 (Correlated Rationalizable Set (CR)).
The product set of actions $\bar{\mathcal{A}}=\bar{\mathcal{A}}_{1}\times\dots\times\bar{\mathcal{A}}_{n}$, $\bar{\mathcal{A}}_{i}\subseteq\mathcal{A}_{i}$, is correlated rationalizable if for each $i\in\mathcal{N}$ and $a_{i}\in\bar{\mathcal{A}}_{i}$
$\exists~{}\hat{x}_{-i}\in\Delta(\bar{\mathcal{A}}_{-i}),~{}\forall~{}a_{i}^{\prime}\in\mathcal{A}_{i}:~{}u_{i}(a_{i},\hat{x}_{-i})\geq u_{i}(a_{i}^{\prime},\hat{x}_{-i}).$
This means that an action $a_{i}\in\mathcal{A}_{i}$ is correlated rationalizable if it maximizes the player’s utility for some probability distribution $\hat{x}_{-i}$ with support on $\bar{\mathcal{A}}_{-i}$. Note the recursive nature of the definition, which ensures that only those actions are in the support of $\hat{x}_{-i}$ that are correlated rationalizable themselves.
Different from the definition of rationalizable actions (Bernheim (1984), Pearce (1984)), correlated rationalizability allows the mixed strategies $\hat{x}_{-i}\in\hat{\mathcal{A}}_{-i}$ to be dependent/correlated probability distributions. The set of rationalizable actions is a subset of the set of correlated rationalizable actions, but the subset relation may not be strict. Bernheim (1984) has shown that all Nash equilibria are rationalizable. As a consequence, all Nash equilibria are also correlated rationalizable. In finite games, iterated strict dominance and correlated rationalizability give the same solution set (Levin 2006). While this set (SSU) can be computed with ISD having the complete payoff matrix available, we will show that uncoupled mean-based algorithms also find specific instances of the correlated rationalizable set.
3.3.4 Summary
We summarize the relation between different solution concepts for games in Figure 2. It is well known that $NE\subseteq CE\subseteq CCE$ in finite games. The relationship between correlated rationalizable sets (or strictly serially undominated sets) and [coarse] correlated equilibria is more complex. First, we can show that, similar to rationalizable actions, actions contained in the support of correlated equilibria are also correlated rationalizable.
Proposition 3.1.
Given a finite normal-formal game, any action profile that is in the support of a correlated equilibrium is also in the correlated rationalizable set.
While it is known that CCE may contain dominated strategies (Viossat and Zapechelnyuk 2013), we can show that actions supported by a CCE are not a superset of correlated rationalizable actions. To show this, we use the equivalence of CR and SSU.
Proposition 3.2.
There exists a finite-normal form game where an action from the strictly serially undominated set is not in the support of any CCE.
Please refer to the appendix (Section A.2) for both proofs.
NECECR=SSUCCE
Figure 2: Relation Between Solution Concepts.
The displayed relations are between the support sets of the solution concepts.
CCE = Coarse Correlated Equilibrium, SSU = Strictly Serially Undominated, CR = Correlated Rationalizable, CE = Correlated Equilibrium, NE = Nash Equilibrium.
As indicated earlier, algorithms that have no (external) regret (see Equation (1)) converge to a CCE, while algorithms with no internal regret (such as regret matching) converge to CEs (Foster and Vohra 1997). ISD computes the strictly serially undominated set (SSU) and, thus, also the correlated rationalizable set (CR), but this algorithm requires full access to the payoff matrix.
Table 1 summarizes properties of games that are satisfied by Bertrand competition with different demand models (all-or-nothing, linear, and logit demand) and algorithms that are known to converge to a Nash equilibrium if certain of the game properties hold.
Table 1: Overview of Games and Properties.
Game Properties Demand Model Convergence to NE
All-or-Nothing Linear Logit
Potential ✗ (Prop. A.3) ✓ (Prop. A.2) ✗ (Prop. A.4) FP, BR, Exp3
Supermodular & Unique NE ✗ (Prop. A.1) ✓ (MR90) ✓ (MR90) ISD, FP, BR (MR90)
[…]
Fictitious Play (FP), Best Response (BR), Exponential Weights (Exp3), Iterated Strict Dominance (ISD), Mean-Based (MB); (MR90) refers to Milgrom and Roberts (1990).
1 The set of CR is almost unique. We show that only two neighboring actions per agent can be in CR. Depending on the discretization, either one or both are NE.
2 Numerically, we observe two neighboring or one unique NE in the game, depending on the discretization. Only the actions of the NE are contained in CR.
The Bertrand competition with linear demand is a potential game for which we know that even bandit algorithms such as Exp3 converge. However, other demand types are no potential games as we showed earlier. Milgrom and Roberts (1990) showed that Bertrand competitions with continuous actions and linear demand are supermodular and have a unique Nash equilibrium, while with logit demand, they are log-supermodular. We don’t know of convergence proofs for bandit algorithms, but we know that ISD, Fictitious Play, and best response dynamics find the unique Nash equilibrium in such supermodular games. ISD also finds correlated rationalizable sets, because they are equivalent to serially undominated sets. Our main analytical contribution is to show that mean-based algorithms, with Exp3 as the leading example, converge to the correlated rationalizable set. Importantly, in a Bertrand competition with all-or-nothing or linear demand, this set is either a singleton, meaning it must be the Nash equilibrium, or consists of only two adjacent actions.
4 Convergence of Mean-Based Algorithms
Let us now focus on mean-based algorithms. These are online optimization algorithms that pick actions with low average rewards with low probability (Braverman et al. 2018).
Definition 4.1 (Mean-based Algorithm).
Let $\alpha_{t}(a)$ be the average reward of action $a\in\mathcal{A}$ in the first $t-1$ rounds: $\alpha_{t}(a)=\frac{1}{t-1}\cdot\sum_{s=1}^{t-1}u_{s}(a)$. An algorithm is $\gamma_{t}$-mean-based if, for any $a\in\mathcal{A}$, whenever there exists $a^{\prime}\in\mathcal{A}$ such that $\alpha_{t}(a^{\prime})-\alpha_{t}(a)>\gamma_{t}$, the probability that the algorithm picks $a$ at round $t$ is at most $\gamma_{t}$. An algorithm is mean-based if it is $\gamma_{t}$-mean-based for some decreasing sequence $(\gamma_{t})_{t=1}^{\infty}$ such that $\gamma_{t}\rightarrow 0$ as $t\rightarrow\infty$.
There are many no-regret algorithms, such as Exp3, Multiplicative Weights, and the Follow-the-Perturbed-Leader algorithm, which are also mean-based. We know from Kolumbus and Nisan (2022) that if a mean-based algorithm converges to a CCE, then the CCE is co-undominated, which avoids outcomes with dominated strategies as illustrated in Figure 1(b). Deng et al. (2022) analyzed the convergence of mean-based algorithms in first-price auctions, and they found convergence in the complete-information game. We focus on Bertrand pricing competitions, which are different due to the various demand models. In what follows, we provide proof that mean-based algorithms converge to strategy profiles in the correlated rationalizable set.
We analyze convergence in a time-average and in a last-iterate sense (Anagnostides et al. 2024). While the former measures the fraction of times that all players have chosen actions according to a reference profile or reference set of actions, the latter makes a statement about the current strategy of the players, i.e., their current distribution over action selections.
Definition 4.2 (Time-average Convergence).
A sequence of action profiles $\{a_{s}\}_{s=1}^{t}$ converges in time-average to a set of actions $\bar{\mathcal{A}}\subseteq\mathcal{A}$ if
$\lim_{t\to\infty}\frac{1}{t}\sum_{s=1}^{t}\mathbb{I}[\forall i:a_{is}\in\bar{\mathcal{A}}]=1.$
The sequence converges in time-average to a single action profile $\bar{a}$ if $\bar{\mathcal{A}}=\{\bar{a}\}$.
With this definition at hand, our central convergence result is as follows:
Theorem 4.1.
In finite normal-form games, mean-based algorithms converge almost surely to a set of correlated rationalizable actions in time-average.
The complete proof can be found in the appendix (Section A.4). Let us briefly describe a proof sketch. We want to show that all agents play only correlated rationalizable actions with probability close to 1 as time goes to infinity. Towards this result, we do the following:
  • Step 1 (Proposition A.8):
    We show that every mean-based agent tends to choose correlated rational responses to their competitors’ past actions with high probability. In the context of Bertrand competition games, this means that an individual agent undercuts the opponents’ prices if the prices are too high and increases prices if the prices are too low.
  • Step 2. (Proposition A.9):
    By an inductive argument, we find that agents will settle on correlated rationalizable actions. Assume that all players have been playing a subset of the action space for a prolonged time. Every mean-based player will thus select actions that are correlated rational responses to their competitors’ past actions. This will continue for an extended number of steps, which changes the set of past actions. As a result, the players adapt their strategy again by selecting correlated rational responses to the new history. By noting that such a behavior of repeated correlated rational responses will lead to a set of correlated rationalizable responses, we can establish our result. Since mean-based algorithms do not "exploit" all of the time but sometimes choose "exploratory" actions at random, we must rely on probabilistic arguments to show that this happens with probability one.
So far, we have shown that players will eventually play correlated rationalizable actions in almost every round. But we can also say something about the actual (mixed) strategy profile of the players, i.e., about last-iterate convergence. In the general setting of the previous theorem, we can only show that the support of the mixed strategies is within the correlated rationalizable actions in the limit. In the special case, where the correlated rationalizable actions correspond to the unique Nash equilibrium, e.g., in the all-or-nothing demand setting as described in Proposition 4.1, we get last-iterate convergence to the Nash equilibrium.
Corollary 4.1.
In games with a unique pure Nash equilibrium $a^{*}$ being the only correlated rationalizable action, mean-based algorithms converge almost surely to the Nash equilibrium in last-iterate, i.e., $\lim_{t\to\infty}x_{t}=\bar{x}$, where $x_{t}$ is the sequence of mixed strategy profiles and $\bar{x}_{i}=\mathbf{1}_{a_{i}^{*}}$.
Proof.
From Lemma A.3 we get that $P_{t-1}(\mathcal{A}^{K-1})\geq\tilde{p}$ for sufficiently large $t$. According to Proposition A.8, a player $i$ plays outside of $\bar{\mathcal{A}}=\mathcal{A}^{K}$ with probability at most $\gamma_{t}$. This means that the player plays inside $\bar{\mathcal{A}}$ with probability at least $1-\gamma_{t}$. As $t$ approaches infinity, this probability approaches 1, showing that the player converges to rationalizable actions in last-iterate, which corresponds by assumption to the action of the Nash equilibrium $a^{*}$. The same statement holds for all players. ∎
Our analysis leads to a corollary that is of independent interest and beyond the analysis of Bertrand competitions.
Corollary 4.2.
Mean-based algorithms converge to strategy profiles in the serially undominated set.
This follows from the fact that serially undominated sets are equivalent to correlated rationalizable sets of strategy profiles. Next, we focus on the class of Bertrand competition games. Let us present our second main result informally.
Informal theorem 1 (Informal theorem).
In symmetric Bertrand competition games with standard or linear demand and discrete actions, mean-based algorithms converge almost surely to actions that are close to their Nash equilibria.
For this, we just need to show that the correlated rationalizable sets are close to the set of Nash equilibria for the two demand models.
This leads to a first result on Bertrand competition with all-or-nothing demand.
Proposition 4.1.
In the symmetric Bertrand competition game with all-or-nothing demand and action space $\mathcal{A}_{i}=\{a^{1},\dots,a^{m}\}$, $i=1,\dots,n$, with $c<a^{1}<\dots<a^{m}\leq 1$ and $a^{k-1}-(a^{k}-a^{k-1})\geq c$, the set of correlated rationalizable actions is $\bar{\mathcal{A}}_{i}=\{a^{1}\}$ for every player, which is equivalent to the unique Nash equilibrium of the game.
We show this statement in the appendix (Section A.3). One example of a discretization that satisfies this assumption is an action space with equidistant actions, i.e., $\mathcal{A}=\{c+\varepsilon,c+2\varepsilon,\dots,c+K\varepsilon\}$ with $\varepsilon>0$, $c+K\varepsilon\leq 1$, and $K\geq 2$. In a similar attempt, we can make a statement about Bertrand competitions with linear demand.
Proposition 4.2.
For a symmetric Bertrand competition game with linear demand, $\gamma<\beta$, and action space $\mathcal{A}=\{a^{1},\dots,a^{m}\}$ where $0\leq a^{1}\leq\dots,<a^{k^{*}}\leq p^{*}<a^{k^{*}+1}\dots<a^{m}$ with $p^{*}=\frac{\alpha+\beta c}{2\beta-\gamma}$, all correlated rationalizable sets are subsets of $\{a^{k^{*}},a^{k^{*}+1}\}$ for every player $i$.
Again, we provide the proof in the appendix (Section A.3). Depending on the discretization, the Bertrand competition with linear demand can have more than one pure Nash equilibrium. Recall that all Nash equilibria are correlated rationalizable.
5 Experimental Analysis
In our experimental analysis, we analyze the behavior of various bandit algorithms in Bertrand oligopolies with different demand models. The focus variables of our experiments are the converged prices and the convergence behavior. We focus on scenarios that our theoretical contribution does not cover. We use algorithms that are well-known but (with the exception of Exp3) not known to be mean-based and simulate their behavior in a variety of Bertrand oligopolies.
In our numerical experiments, most algorithms consistently reach Nash equilibrium prices within a few iterations, indicating that even without tens of thousands of repeated interactions and without the mean-based property, algorithmic collusion is rarely established. By including different demand functions and symmetric as well as asymmetric environments, we show that our findings are robust across a wide range of treatments.
5.1 Selected Algorithms
In our study, we include $\epsilon$-greedy, UCB, Thompson Sampling, and Exp3, which are widely used and cited in the literature (Bubeck 2011). All bandit algorithms share similarities. First, an action is selected according to some potentially random action-selection rule that is based on past experiences. Then, the agent receives a corresponding reward feedback with which it subsequently updates its internal state. We describe this generic scheme in the Algorithm 1 below. All algorithms that we analyze follow this generic scheme while implementing distinct rules for selecting actions and updating their beliefs. Details of the algorithms can be found in online supplement B.1. Typically, the algorithms allow for different variations, and each description describes a family of algorithms.
Algorithm 1 Bandit Algorithm Scheme
Initialization: Algorithm parameters $\theta$
for $t=1,2,\dots$ do
     $a_{t}\leftarrow$ get_action($\theta$, $t$) // Agents choose action.
     $u_{t}\leftarrow$ environment($a_{t}$) // Agents submit actions and observe own utility.
     $\theta\leftarrow$ update($\theta$, $a_{t}$, $u_{t}$, $t$) // Agents update their strategy.
end for
5.2 Experiment Setup
Algorithmic collusion can lead to supra-competitive prices between the Nash equilibrium prices and the monopoly prices. Prices that are not supra-competitive will be called competitive in our results. We demonstrate that even in treatments where no analytical convergence results are available, convergence to competitive prices can be observed consistently with many bandit algorithms. We characterize supra-competitive prices via the following metrics. Let $(\bar{a}_{i})_{i=1}^{n}$ be the largest component strategies of any pure Nash equilibrium, and let $(a^{M}_{i})_{i=1}^{n}$ be the actions played in the monopoly. We define the price-collusion index of player $i$
$CI_{price,i}(a_{i})=\frac{a_{i}-\bar{a}_{i}}{a^{M}_{i}-\bar{a}_{i}},~{}i\in[n]$ (6)
and the profit-collusion index of all players
$CI_{profit}(a)=\frac{\sum_{i=1}^{n}(u_{i}(a)-u_{i}(\bar{a}))}{\sum_{i=1}^{n}(u_{i}(a^{M})-u_{i}(\bar{a}))}.$ (7)
A price-collusion index of zero captures that player $i$ plays $\bar{a}_{i}$, and thus non-collusive. Similarly, an index value of one means that the agent plays its monopoly price, indicating perfect collusion. Analogously, the profit-collusion index represents the fraction of monopolistic payoff surplus that all agents could achieve in total.
In the following, we introduce the treatment variables of our experiments, such as the oligopoly configurations and the bandit algorithms, as well as the experiment setup. We then state the conclusions drawn from our experiments and present the empirical evidence supporting them.
In our experiments, we vary the games that the agents compete in, the algorithms they use, and the number of competitors they face. An experiment is formed by a unique combination of these factors. Table 2 provides an overview of the treatment and focus variables. Below, we describe the individual treatments in more detail.
Table 2: Treatment and Focus Variables of the Experiments.
Treatment Variables Variations
Oligopoly model Demand models (standard, linear, logit), Symmetric/asymmetric parameterization
Algorithms UCB-T, Exp3-$\epsilon$, $\epsilon$-Greedy, TS, (UCB1, Q-Learning, MWU)
Number of competitors 2 - 10
Focus Variables Description
Prices Prices after convergence
Convergence speed Time to convergence and convergence behavior
Oligopoly Model.
Our oligopolies model the demand and the costs that firms face. They are based on the games described in section 3.2. We implemented models with standard, linear, and logit demand functions. Depending on the parameter choices, the utility functions are the same for all agents (symmetric game) or not (asymmetric game). Our experiments are based on the following instances of Bertrand competitions. The first three configurations are symmetric games with standard demand (O1, $c_{i}=c=0.0$), linear demand (O2, $c_{i}=c=0$, $\alpha=0.48$, $\beta=0.9$, $\gamma=0.6$), and logit demand (O3, $c_{i}=c=0$, $a_{i}=a_{0}=a=2.0$, $\mu=0.25$). Likewise, we define asymmetric counterparts for duopolies with linear demand (O2’, $c_{1}=0$, $c_{2}=0.2$) and logit demand (O3’, $c_{1}=0.5$, $c_{2}=1.0$, $a_{1}=1.5$, $a_{2}=2.0$, $\mu=0.25$). The asymmetric settings are a step towards more realistic scenarios where firms might face different costs and demand structures. In oligopolies with all-or-nothing demand, agents can choose actions within $[0.05,1.0]$; in oligopolies with standard and linear demand, they can choose from actions in $[0,1]$ and $[1,2]$, respectively.
Algorithms.
We investigate the behavior of four widespread bandit algorithms. First, we evaluate a variant of the UCB-Tuned (UCB-T) algorithm that resolves ties randomly and does not eliminate actions. Next, we run experiments with the simple $\epsilon$-greedy algorithm ($\epsilon$-Greedy) and an Exp3 (Exp3-$\epsilon$) algorithm with fixed exploration rate $\epsilon$. We note that this version of Exp3 is not mean-based. Known mean-based versions of Exp3 suffer from a very slowly decaying exploration rate, which makes them impractical for experiments and applications. Finally, we implemented an algorithm (TS) that performs Thompson Sampling with Gaussian priors, likelihoods, and posteriors. We note that our experiment results are stable against reasonable modifications of the parameters. In one result on mean-based algorithms, we also analyze Multiplicative Weights Update (MWU) and the mean-based version of Exp3 introduced by Braverman et al. (2018). In the appendix, we also report results on Q-learning to provide a comparison. However, given the multitude of hyperparameters available in this algorithm, we restrict our analysis to a few experiments replicating prior results. We also compare these results to those of the algorithms in the main part of the article.
Number of Competitors.
Some of our empirical results are based on the number of firms that compete in an oligopoly. The symmetric games (O1, O2, O3) can be easily extended to an arbitrary number of players. We investigate numbers between two and ten for an oligopoly based on O2.
Throughout our experiments, we keep certain experimental settings constant. We discretize the action interval for each game into 21 equidistant points, including the boundaries of the intervals described above. We run each experiment ten times with random seeds from 0 to 9. Fixing the seeds allows us to reproduce results despite the random nature of the algorithms. Each run consists of 250,000 iteration steps during which the agents first submit their actions simultaneously and independently to the market. Then, the demands are evaluated, and the agents observe their respective rewards. The agents update their beliefs at the end of each step. We found that almost all of our experiments converged within the provided time frame.
5.3 Results
Let us now report the main results of our experimental analysis.
5.3.1 Convergence Speed with Mean-based Algorithms
Let us first focus on markets with purely mean-based algorithms, for which we have a convergence guarantee. Speed of convergence is unknown, however. We consider the mean-based versions of Multiplicative Weights and Exp3 that have been described by Braverman et al. (2018). Figure 3 shows the evolution of the price collusion index over the course of the experiments for both algorithms. We observe that both converge to very low index values. While the MWU algorithm reaches the final outcome within a few iterations, the Exp3 algorithm requires significantly more iterations. In the remainder, we use a different version of Exp3, which is even faster (but not mean-based in a strict sense).
(a) Multiplicative Weights
(b) Exp3 (mean-based)
Figure 3: Evolution of Prices During Training for Mean-based Algorithms.
The experiments are based on the symmetric environment with linear demand (O2). The figures display the collusion indices based on the charged prices (running medians over 1000 steps) for ten runs of the experiment in thin lines. The thick line shows the average of all runs and thus indicates the overall trend.
5.3.2 Algorithmic Collusion in Duopolies
Next, we analyze duopolies with different algorithms that are not mean-based anymore. We first report the prices at the end of the experiments in two-player games. After 250,000 iterations, the algorithms consistently settle at a price level (see next section), so we can assume convergence. Our first result makes a statement about the collusion levels agents eventually reach.
Result 1.
Supra-competitive pricing only evolves with the UCB algorithm in symmetric environments. Other algorithms consistently price close to the Nash equilibrium prices. In particular, combinations of diverse algorithms rarely display collusive behavior. This happens for all demand models (standard, linear, logit) and for symmetric and asymmetric settings.
We visualize these results in Figure 4, which displays the profit- and price-collusion indices found in our experiments. We average the indices over settings, runs, and - if applicable - over agents. In settings with two identical competing algorithms (diagonal entries), we only observe collusive behavior with the UCB-T algorithm. The combined settings of any two different algorithms show collusion indices close to zero, with a combination of UCB-T and $\epsilon$-Greedy resulting in only slightly supra-competitive prices. We observe collusion both in terms of prices and profit, meaning that colluding algorithms indeed benefit from their non-competitive play.
Our experiments also demonstrate that collusion with Bandit algorithms does not develop in a stable manner. While we could observe collusion with UCB-T for all demand functions, and in symmetric as well as asymmetric games, these experiments usually suffer from a high variance between runs, making the result unpredictable in advance. In contrast, non-colluding algorithms performed consistently in all environments and all runs.
(a) Price-collusion indices
(b) Profit-collusion indices
Figure 4: Collusion in Duopolies.
We visualize the collusion indices, based on (a) the median prices and (b) the mean profits in the last 1000 steps, of several combinations of algorithms. The displayed numbers are the numeric values of the collusion indices; the values in brackets display their corresponding standard deviation. Averages and standard deviations were computed over two agents, 10 runs, and the five settings O1 - O3’.
5.3.3 Algorithmic Collusion with More Sellers
There may be more than two competitors in a market. There are two major questions we answer in this context: 1) Do colluding algorithms such as UCB-T eventually reach competitive prices if the number of competitors grows large enough? 2) How many non-colluding agents must be inserted in a system of colluding agents to break supra-competitive pricing? Our second result sheds light on these questions.
Result 2.
With an increasing number of competitors, collusion indices decrease. Pure UCB-T oligopolies can still achieve supra-competitive prices but at a lower level. Adding other bandit algorithms reduces algorithmic collusion.
To answer the first question, we ran experiments with identical algorithms in the linear demand setting (O2). Figure 5(a) shows that with an increasing number of prices, UCB-T comes closer to the Nash equilibrium prices. However, price levels remain collusive in oligopolies of up to ten players. In line with our previous experiments, the other algorithms remain competitive for any number of players.
We answer the second question by configuring oligopolies where all players commit to the UCB-T algorithm except for one. The results are displayed in Figure 5(b). While the collusion levels decrease compared to the pure UCB-T oligopolies, the UCB-T algorithms are still able to coordinate on supra-competitive prices.
(a) Identical algorithms
(b) UCB-T majority
Figure 5: Profit Collusion with Increasing Number of Players.
The plots visualize the price-collusion indices at the end of training for different numbers of players in a Bertrand economy with linear demand. a) shows oligopolies with identical algorithms while b) shows the combination of $n-1$ UCB-T algorithms with one different algorithm. The mean profit of the last 1000 iteration steps is averaged over all runs to form the index.
Of course, this coordination is only possible if there are sufficiently many players to support it. We analyze the competition of varying shares of UCB-T agents in a 10-player oligopoly with competing Exp3-$\epsilon$ algorithms with linear demand (O2). The result is displayed in Figure 6. We only observe collusion if the majority of agents play UCB-T. With increasing numbers, collusion increases, too.
Figure 6: Different Numbers of UCB-T Agents Against Exp3-$\epsilon$ Competitors in a 10-player Oligopoly.
The plot visualizes the price-collusion index in 10-player oligopoly experiments with linear demand (O2) for an increasing share of UCB-T players. The remaining agents act according to Exp3-$\epsilon$. The mean profit of the last 1000 iteration steps is averaged over all runs to form the index.
5.3.4 Convergence Speed
Aside from the converged prices, we also analyzed the training behavior of the algorithms. As with the final prices, we find significant differences between the UCB-T algorithms and the other algorithms, which we summarize as our third result.
Result 3.
Exp3-$\epsilon$, TS, and $\epsilon$-Greedy converge quickly and reliably to the Nash equilibrium prices. UCB-T combinations erratically learn to charge higher prices with large differences within and between runs.
In Figures 7 and 8, we provide the price-collusion indices for all four algorithms in symmetric duopolies with linear demand (O2). Indeed, we find that the convergence of Exp3-$\epsilon$ and TS is particularly consistent and quick. The former algorithm sets prices close to the competitive Nash equilibrium after roughly 50,000 steps, while the latter requires even fewer interactions between the agents. With $\epsilon$-Greedy, the agents arrive at competitive prices quickly, with small variations between runs. We attribute this to the fixed exploration rate of our algorithm configuration for $\epsilon$-Greedy.
(a) Exp3-$\epsilon$
(b) TS
Figure 7: Evolution of Prices During Training.
The figures display the collusion indices based on the charged prices (running medians over 1000 steps) for ten runs of the experiment in thin lines. The thick line shows the average of all runs and thus indicates the overall trend.
(a) $\epsilon$-Greedy
(b) UCB-T
Figure 8: Evolution of Prices During Training.
The figures display the collusion indices based on the charged prices (running medians over 1000 steps) for ten runs of the experiment in thin lines. The thick line shows the average of all runs and thus indicates the overall trend.
In stark contrast to this, the UCB-T experiments exhibit a larger variation of prices, both within runs and between runs. The variation within runs indicates that agents do not simply "agree" on a supra-competitive price level. Instead, they erratically raise and lower prices in a seemingly random pattern. Since these agents choose actions by maximizing their heuristic (which is a non-smooth operation), their actions can vary drastically if multiple actions experience values of the heuristics at around the same level. Each run of our UCB-T experiments is different, and we observed a strong variation between the runs. We note that the only source of randomness comes from the random action selection if one or more actions have the same heuristic value. The varying outcomes make it hard for firms to reliably predict their returns when employing a UCB-T algorithm in this scenario. On average, though, we observe collusion, starting from an early point in the training.
6 Conclusions
The discussion of algorithmic collusion is largely based on the experimental results of specific algorithms and versions of the Bertrand competition. In general, learning algorithms do not converge to an equilibrium. However, we prove that important classes of algorithms do so in repeated Bertrand pricing competition with all-or-nothing and linear demand models. The convergence of mean-based algorithms to correlated rationalizable strategies is of independent interest and might be useful for the analysis of different games as well. The result closes a gap in the literature on learning in games, which largely focused on algorithms with no internal or external regret, which converge to correlated or coarse correlated equilibria, resp. Mean-based algorithms such as Exp3 only require bandit feedback after each round, a realistic assumption for algorithms used in algorithmicp pricing. The fact that correlated rationalizable strategy profiles coincide with the set of Nash equilibria in Bertrand competition games is an important insight and it shows that such games can be learned even with very simple algorithms relevant in practice.
We also show experimentally that when different (not mean-based) online optimization algorithms are used in combination, algorithmic collusion is not sustained in experiments. If sellers use UCB or Q-learning symmetrically, they might learn prices higher than the Nash equilibrium. However, this can be exploited by agents who deliberately undercut those Q-learning agents and increase their profit unilaterally. For example, den Boer et al. (2022) show that Exp3 outperforms Q-learning in terms of profit, meaning that Q-learning is not a good choice in the meta-game. Therefore, incentives to use such algorithms might be low unless sellers actively collude and agree on the use of algorithms with a tendency for supra-competitive prices. Such insights are important for regulators when they want to identify collusion in online markets. While multi-armed bandit algorithms are an important and widely used class of algorithms used for algorithmic pricing, one cannot guarantee that sellers don’t use other algorithms. In future research, it will be valuable to analyze and classify alternative algorithms and models. The properties that we showed for Bertrand competition will be useful also for subsequent studies.
Acknowledgments
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - GRK 2201/2 - Project Number 277991500 and BI 1057/9.
Appendix A Proofs
A.1 Proofs of Game-Theoretical Properties
Proposition A.1.
The Bertrand competition with all-or-nothing demand is not supermodular.
Proof.
To prove that the Bertrand-competition with standard all-or-nothing demand is supermodular, we show that the property of increasing differences, i.e., $u_{i}(a_{i}^{\prime},a_{-i}^{\prime})-u_{i}(a_{i},a_{-i}^{\prime})\geq u_{i}(a_{i}^{\prime},a_{-i})-u_{i}(a_{i},a_{-i})$ with $a_{i}^{\prime}\geq a_{i}$ and $a_{j}^{\prime}\geq a_{j}$ for all $j\neq i$, is not satisfied. To that end, we consider two cases in the simple $n=2$ player case:
For $a_{2}<a_{1}<a_{2}^{\prime}<a_{1}^{\prime}$, the inequality is violated:
$u_{1}(a_{1}^{\prime},a_{2}^{\prime})-u_{1}(a_{1},a_{2}^{\prime})$ $=0-u_{1}(a_{1},a_{2}^{\prime})<0$
$u_{1}(a_{1}^{\prime},a_{2})-u_{1}(a_{1},a_{2})$ $=0-0=0$
For $a_{1}<a_{2}<a_{1}^{\prime}<a_{2}^{\prime}$ with $1-(a_{1}-c_{1})>a_{1}^{\prime}$, the inequality is satisfied:
$u_{1}(a_{1}^{\prime},a_{2}^{\prime})-u_{1}(a_{1},a_{2}^{\prime})$ $=(1-a_{1}^{\prime})(a_{1}^{\prime}-c_{1})-(1-a_{1})(a_{1}-c_{1})>0$
$u_{1}(a_{1}^{\prime},a_{2})-u_{1}(a_{1},a_{2})$ $=0-u_{1}(a_{1},a_{2})<0$
The condition $1-(a_{1}-c_{1})>a_{1}^{\prime}$ simply means that $a_{1}$ is closer to $c_{1}$ than $a_{1}^{\prime}$ is to 1, which is necessary to ensure that $u_{1}(a_{1}^{\prime},a_{2}^{\prime})>u_{1}(a_{1},a_{2}^{\prime})$. Since the inequalities is strictly satisfied and violated, we cannot define a different order on the opponent’s action set such that the increasing differences property is satisfied. Therefore, the game cannot be supermodular. Note that this argument also works for continuous actions. ∎
Proposition A.2.
The Bertrand competition with linear demand is a potential game.
Proof.
To show that the Bertrand competition with linear demand is a potential game, we prove that a function $\phi:\mathcal{X}\rightarrow\mathbb{R}$ satisfies the definition of a potential. Our choice of $\phi$ is similar to the potential function of the Cournot setting in Slade (1994) and is given by
$\phi(a)=\sum_{k}\alpha_{k}a_{k}-\sum_{k}\beta_{k}a_{k}(a_{k}-c_{k})+\tfrac{\gamma}{n-1}\sum_{k<l}a_{k}a_{l}.$
It is easy to check that this function satisfies the condition of an exact potential:
$\phi(a_{i},a_{-i})-\phi(a^{\prime}_{i},a_{-i})$ $=\alpha_{i}[a_{i}-a^{\prime}_{i}]-\beta_{i}[a_{i}(x_{i}-c_{i})-a^{\prime}_{i}(a^{\prime}_{i}-c_{i})]$
$\quad+\tfrac{\gamma}{n-1}\sum_{l>i}[a_{i}-a^{\prime}_{i}]a_{l}+\tfrac{\gamma}{n-1}\sum_{k<i}a_{k}(a_{i}-a^{\prime}_{i})$
Note that in the sum over $k,l$, we have the two cases where either $k=i$ or $l=i$, which leads to the two sums above. We can combine them and get:
$=\alpha_{i}(a_{i}-a^{\prime}_{i})-\beta_{i}[a_{i}(a_{i}-c_{i})-a^{\prime}_{i}(a^{\prime}_{i}-c_{i})]+\tfrac{\gamma}{n-1}\sum_{j\neq i}a_{j}(a_{i}-a^{\prime}_{i})$
$=d_{i}(a_{i},a_{-i})\cdot(a_{i}-c_{i})-d_{i}(a^{\prime}_{i},a_{-i})\cdot(a^{\prime}_{i}-c_{i})=u_{i}(a_{i},a_{-i})-u_{i}(a^{\prime}_{i},a_{-i}).$
This argument works for both continuous and discrete actions. ∎
Proposition A.3.
The Bertrand competition with all-or-nothing demand is not a potential game.
Proof.
To show that the Bertrand competition with standard demand is not a potential game, it is sufficient so show that there exists an example, where the closed path property is not satisfied (Monderer and Shapley 1996, Theorem 2.8). For simplicity, we focus on the two player version, but the argument can be extended to more agents. Consider the actions $10<a_{1}<a_{2}<a_{1}^{\prime}<a_{2}^{\prime}<$ 1 and and the closed path $\gamma$ of length 4 given by the action profiles $(a_{1},a_{2})\rightarrow(a_{1}^{\prime},a_{2})\rightarrow(a_{1}^{\prime},a_{2}^{\prime})\rightarrow(a_{1},a_{2}^{\prime})\rightarrow(a_{1},a_{2})$. We know that our game is a potential game if and only if $I(\gamma,u)=0$. In this example, $I(\gamma,u)$ is given by
$I(\gamma,u)$ $:=[u_{1}(a_{1}^{\prime},a_{2})-u_{1}(a_{1},a_{2})]+[u_{2}(a_{1}^{\prime},a_{2}^{\prime})-u_{2}(a_{1}^{\prime},a_{2})]$
$\quad+[u_{1}(a_{1},a_{2}^{\prime})-u_{1}(a_{1}^{\prime},a_{2}^{\prime})]+[u_{2}(a_{1},a_{2})-u_{2}(a_{1},a_{2}^{\prime})]$
Using the definition of the standard demand model, we get
$=[0-D(1-a_{1})(a_{1}-c_{1})]+[0-D(1-a_{2})(a_{2}-c_{2})]$
$\quad+[D(1-a_{1})(a_{1}-c_{1})-D(1-a_{1}^{\prime})(a_{1}^{\prime}-c_{1})]+[0-0]$
$=-D(1-a_{2})(a_{2}-c_{2})-D(1-a_{1}^{\prime})(a_{1}^{\prime}-c_{1})$
If we further assume that both agents bid above their marginal costs, i.e., $c_{1}<a_{1}<1$ and $c_{2}<a_{2}$ the last line is strictly negative and thereby $I(\gamma,u)\neq 0$, which implies that the game cannot be a potential game. ∎
Proposition A.4.
The Bertrand competition with logit demand is in general not a potential game.
Proof.
Since the utility function in the Bertrand competition with logit demand is twice differentiable, we can use that the game is a potential game if and only if $\tfrac{\partial^{2}u_{i}}{\partial a_{i}\partial a_{j}}=\tfrac{\partial^{2}u_{j}}{\partial a_{i}\partial a_{j}}$ (Monderer and Shapley 1996, Theorem 4.5). Again, we focus on a simplified version to show that the game is in general not a potential game. Let us assume that we have $n=2$ agents with costs $c_{1},c_{2}>0$ and $\alpha_{0}=\alpha_{1}=\alpha_{2}=\alpha$, and $\mu_{0}=\mu_{1}=\mu_{2}=\mu$. If we compute the second derivatives we get
$\quad\frac{\partial^{2}u_{1}(a_{1},a_{2})}{\partial a_{1}\partial a_{2}}=\frac{\partial^{2}u_{2}(a_{1},a_{2})}{\partial a_{1}\partial a_{2}}$
[…]
$\qquad+2\left(a_{1}-c_{1}\right)e^{\frac{-2a_{1}-a_{2}+3\alpha}{\mu}}$
$=\mu\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+\left(-a_{2}+c_{2}\right)\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}$
$\qquad+2\left(a_{2}-c_{2}\right)e^{\frac{-a_{1}-2a_{2}+3\alpha}{\mu}}$
[…]
$=\left(a_{2}-c_{2}\right)\left(-\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+2e^{\frac{-a_{1}-2a_{2}+3\alpha}{\mu}}\right)$
It is easy to see that if both agents bid for instance $c_{1}$, only the left side is equal to zero while the right side is generally not. And even if agents are symmetric with respect to the costs, i.e., $c_{1}=c_{2}=c$, and assume that agents bid $a_{1}=2c$ and $a_{2}=3c$ we get
[…]
Therefore the condition from the theorem is not satisfied and the Bertrand compition with logit demand is in general not a potential game. ∎
A.2 Remarks on Correlated Rationalizability
In this section, we define a series of rational response sets, $(\mathcal{A}^{i})_{i=1}^{K}$, together with a game-dependent number, the competition constant $\delta$, that will play an essential role in our proof of convergence.
Definition A.1 (Correlated Rational Responses).
Given a product set $\hat{\mathcal{A}}=\hat{\mathcal{A}}_{1}\times\dots\times\hat{\mathcal{A}}_{n}$ of action sets $\hat{\mathcal{A}}_{i}\subseteq\mathcal{A}_{i}$, we define player $i$’s correlated rational responses $R_{i}(\hat{\mathcal{A}})$ as
$R_{i}(\hat{\mathcal{A}})=\{a_{i}\in\mathcal{A}_{i}~{}|~{}\exists~{}\hat{x}_{-i}\in\Delta(\hat{\mathcal{A}}_{-i}),~{}\forall~{}a_{i}^{\prime}\in\mathcal{A}_{i}:~{}u_{i}(a_{i},\hat{x}_{-i})\geq u_{i}(a_{i}^{\prime},\hat{x}_{-i})\}$
In other words, the correlated rational responses are the actions which maximize the utility for some probability distribution $\hat{x}_{-i}$ with support on $\hat{\mathcal{A}}_{-i}$. For all actions $a_{i}\not\in R_{i}(\hat{\mathcal{A}})$, we find that there always exists a superior action for every probability distribution $\hat{x}_{-i}\in\Delta(\hat{\mathcal{A}}_{-i})$.
$\forall~{}\hat{x}_{-i}\in\Delta(\hat{\mathcal{A}}_{-i}),~{}\exists~{}a_{i}^{\prime}\in\mathcal{A}_{i}:\quad u_{i}(a_{i},\hat{x}_{-i})<u_{i}(a_{i}^{\prime},\hat{x}_{-i})$ (8)
Based on the concept of correlated rational responses from Definition A.1, we can define an iteration of product sets as follows. Let $\mathcal{A}^{0}=\mathcal{A}_{1}\times\dots\times\mathcal{A}_{n}$ be the cartesian product of the action sets of all players and let $R(\hat{\mathcal{A}})=R_{1}(\hat{\mathcal{A}})\times\dots\times R_{n}(\hat{\mathcal{A}})$ be an operator on the sets of all players’ actions that combines the correlated rational responses of all players. For $k\geq 1$, we iteratively define $\mathcal{A}^{k}=R(\mathcal{A}^{k-1})$. We say that an action is stable under $R$ or $R_{i}$ for a given $\hat{\mathcal{A}}$ if it is contained in $R_{i}(\hat{\mathcal{A}})$. Furthermore, we denote the competition constant $\delta>0$ of the game as the largest number for which the utility difference for non-rational responses to $\mathcal{A}^{k}$ is at least $\delta$ for all $k$:
$\forall k\geq 0,~{}a_{i}\not\in R_{i}(\mathcal{A}^{k}):\quad\forall~{}\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i}),~{}\exists~{}a_{i}^{\prime}\in\mathcal{A}_{i}:~{}u_{i}(a_{i}^{\prime},\hat{x}_{-i})-u_{i}(a_{i},\hat{x}_{-i})\geq\delta.$
Remark A.1.
The sets of correlated rationalizable actions $\bar{\mathcal{A}}$ are fix points under $R(\cdot)$ which follows directly from Definition 3.9.
Remark A.2.
The operator $R(\cdot)$ may cycle for some games and some inputs, meaning that after some number of iterations greater than 1, the same set is obtained: $(R\circ\dots\circ R)(\hat{\mathcal{A}})=\hat{\mathcal{A}}$. However, this can only happen if there are some correlated rational responses to $\hat{\mathcal{A}}$ that are not within $\hat{\mathcal{A}}$. If we start applying $R$ at the full action sets $\mathcal{A}^{0}$ as defined above, we may never end up in such a cyclic relation. Instead, we will reach a fixed point of $R$ that is a correlated rationalizable set. The following proposition makes this statement more formal.
Proposition A.5.
The sets $\mathcal{A}^{k}$ are (non-strict) subsets of their predecessors $\mathcal{A}^{k-1}$: $\mathcal{A}^{k}\subseteq\mathcal{A}^{k-1}$.
Proof.
We prove this statement by induction.
  • Induction start, $k=1$: Since $\mathcal{A}^{0}=\mathcal{A}$, we get $\mathcal{A}^{1}\subseteq\mathcal{A}^{0}$.
  • Induction step, $k\to k+1$: By induction, we know that $\mathcal{A}^{k^{\prime}}\subseteq\mathcal{A}^{k^{\prime}-1}$ for all $k^{\prime}\leq k$. Assume, towards contradiction, that $\mathcal{A}^{k+1}_{i}$ contains an element $a\not\in\mathcal{A}^{k}_{i}$. For this element, there is a mixed opponent strategy $\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i})$ such that $a$ maximizes $u_{i}(a,\hat{x}_{-i})$. Since $\mathcal{A}^{k}_{-i}\subseteq\mathcal{A}^{k-1}_{-i}$, $\hat{x}_{-i}\in\Delta(\mathcal{A}^{k-1}_{-i})$, and thus $a$ must also be in $\mathcal{A}^{k}_{i}$, leading to contradiction.
Finally, we also provide proof for Proposition 3.1 and 3.2.
Proof of Proposition 3.1.
Consider an action $a_{i}\in\mathcal{A}_{i}$ that is supported in some CE, meaning that there is some probability distribution $p$ on the cartesian product of action sets $\mathcal{A}=\mathcal{A}_{1}\times\dots\times\mathcal{A}_{n}$ such that for all $a_{i}^{\prime}\in\mathcal{A}_{i}$
$\operatorname{\mathbb{E}}_{a\sim p}[u_{i}(a)|a_{i}]\geq\operatorname{\mathbb{E}}_{a\sim p}[u_{i}(a_{i}^{\prime},a_{-i})|a_{i}].$
The existence of $p$ implies that there exists a joint distribution $\hat{x}_{-i}$ on the opponents’ actions $\mathcal{A}_{-i}$, namely the posterior distribution of $p$ given $a_{i}$, such that for all $a_{i}^{\prime}$
$u_{i}(a_{i},\hat{x}_{-i})\geq u_{i}(a_{i}^{\prime},\hat{x}_{-i}).$
This shows that $a_{i}$ is a correlated rational response to $\hat{x}_{-i}$. Because this holds for all players in a CE, all players can only choose correlated rational responses. Mutually correlated rational responses show that the actions are indeed correlated rationalizable. ∎
Proof of Proposition 3.2.
We show that not every strictly serially undominated action is supported by some CCE. To that end we consider the following matrix game with two agents $\mathcal{N}=\{1,2\}$ and two actions $\mathcal{A}_{1}=\mathcal{A}_{2}=\{a_{1},a_{2}\}$ each. The utilities are given by the payoff matrices
$P_{1}=\begin{pmatrix}0&2\\ 2&1\end{pmatrix},\quad P_{2}=\begin{pmatrix}2&0\\ 0&0\end{pmatrix},$
where $u_{i}(a_{k},a_{l})=(P_{i})_{kl}$ with $i\in\mathcal{N}$ and $k,l\in\{1,2\}$. It is easy to see that no action is strictly dominated. Therefore the set of serially undominated actions is the whole action set. We will now show that there is no CCE $\sigma=(\sigma_{11},\sigma_{12},\sigma_{21},\sigma_{22})\in\Delta(\mathcal{A})$ where the action $a_{1}$ of agent 1 has a positive probability, i.e., we show that $\sigma_{11}=\sigma_{12}=0$, even though $a_{1}$ is undominated. The constraints for the CCE are given by the following constraints (see Definition 3.7):
$\sigma_{21}(2-0)+\sigma_{22}(1-2)$ $\geq 0$
$\sigma_{11}(0-2)+\sigma_{12}(2-1)$ $\geq 0$
$\sigma_{12}(0-2)+\sigma_{22}(0-0)$ $\geq 0$
$\sigma_{11}(2-0)+\sigma_{21}(0-0)$ $\geq 0$
Since $\sigma$ has to be a probability measure over $\mathcal{A}$, we also have $\sigma_{11}+\sigma_{12}+\sigma_{21}+\sigma_{22}=1$ and $\sigma_{11},\sigma_{12},\sigma_{21},\sigma_{22}\geq 0$ From the third constraint we immediately get $\sigma_{12}\leq 0$ and thereby $\sigma_{12}=0$. Using this in the second constraint gives us $\sigma_{12}\geq 2\sigma_{11}$. Therefore we have $\sigma_{11}=\sigma_{12}=0$, which means that action 1 of agent 1 is not support in any CCE even though it is a undominated action $\Rightarrow$ $SSU\not\subseteq CCE$. ∎
A.3 Correlated Rationalizable Set in Bertrand Oligopolies
Proposition A.6.
In the symmetric Bertrand competition game with standard demand and action space $\mathcal{A}_{i}=\{a^{1},\dots,a^{m}\}$, $i=1,\dots,n$, with $c<a^{1}<\dots<a^{m}\leq 1$ and $a^{k-1}-(a^{k}-a^{k-1})\geq c$, the set of correlated rationalizable actions is $\bar{\mathcal{A}}_{i}=\{a^{1}\}$ for every player, which is equivalent to the unique Nash equilibrium of the game.
Proof.
First, we show that $a_{i}^{*}=a^{1},\,\forall i\in\mathcal{N}$ is the unique pure Nash equilibrium. It is easy to see that it is a NE, since it is only possible to deviate to a higher action which yields a utility of 00 instead of $\tfrac{D}{n}(1-a^{1})(c-a^{1})>0$. It remains to show that there is no other Nash equilibrium. Other candidates have to be symmetric, i.e., $a_{i}^{*}=a^{k},\,\forall i\in\mathcal{N}$ for some $k\in\{2,\dots,m\}$, since otherwise there is at least one agent pricing higher than some other agents, which means that the player has a utility of zero which could be strictly improved by deviating to $a^{1}$. Now assuming that all agents play $a^{k}$, we want to show that deviating to $a^{k-1}$ strictly increases an agent’s utility, showing that $a_{i}^{*}=a^{k}$ for all $i\in\mathcal{N}$ cannot be a NE.
$u_{i}(a^{k-1},a^{k}_{-i})-u_{i}(a^{k},a^{k}_{-i})$ $=D(1-a^{k-1})(a^{k-1}-c)-\tfrac{D}{n}(1-a^{k})(a^{k}-c)$
$\geq\tfrac{D}{n}\left[(1-a^{k-1})(2a^{k-1}-2c)-(1-a^{k-1})(a^{k}-c)+(a^{k}-a^{k-1})(a^{k}-c)\right]$
$=\tfrac{D}{n}\left((1-a^{k-1})(a^{k-1}-(a^{k}-a^{k-1})-c)+(a^{k}-a^{k-1})(a^{k}-c)\right)$
$\geq\tfrac{D}{n}(a^{k}-a^{k-1})(a^{k}-c)>0.$
In the last step, we use the assumption on the discretization, i.e., $(a^{k-1}-(a^{k}-a^{k-1})-c)\geq 0$. Therefore, we have uniqueness of our NE.
Next, we want to show that no other action than $a^{1}$ can be correlated rationalizable. Towards contradiction, assume we have supports $\hat{\mathcal{A}}_{1},\dots,\hat{\mathcal{A}}_{n}$ for all players where at least one player can play an action unequal $a^{1}$. We show that there must be some player $i$ and one action for which Equation (8) holds, meaning that the support will change under $R_{i}$. W.l.o.g. let player $i$ support actions up to $a^{k},k\in\{2,\dots,m\}$, the highest action of any support. (There could be multiple such players.) The other players have a joint distribution $x_{-i}$ with support only on actions up to $a^{max}=a^{k}$. Let us define $\Delta:=\max_{a}u_{i}(a)$ and $p\in[0,1]$, which denotes the probability that all opponents play $a^{k}$ under $x_{-i}$. Depending on $x_{-i}$ (in particular $p$), we will show that either $a^{1}$ or $a^{k-1}$ can be chosen to get inequality (8).
  1. (1)
    If $p\leq\frac{1}{2\Delta}(1-a^{1})(a^{1}-c)$, we choose $a_{i}^{\prime}=a^{1}$:
    $\operatorname{\mathbb{E}}_{-i}\left[u_{i}(a^{1},a_{-i})-u_{i}(a^{k},a_{-i})\right]$ $=p\left[u_{i}(a^{1},a^{k}_{-i})-u_{i}(a^{k},a^{k}_{-i})\right]+(1-p)\left[u_{i}(a^{1},a^{1}_{-i})+\dots\right]$
    $\geq p\tfrac{D}{n}\left[(n-1)(1-a^{1})(a^{1}-c)-(1-a^{k})(a^{k}-c)\right]+\tfrac{D}{n}(1-a^{1})(a^{1}-c)$
    $>-p\tfrac{D}{n}\Delta+\tfrac{D}{n}(1-a^{1})(a^{1}-c)$
    $>\tfrac{D}{2n}(1-a^{1})(a^{1}-c)=\frac{1}{2}u_{i}(a^{1},a^{1}_{-i})>0$
    For the last line we used that $(n-1)(1-a^{1})(a^{1}-c)>0$, $-(1-a^{k})(a^{k}-c)\geq-\Delta$, and $p<\frac{1}{2\Delta}(1-a^{1})(a^{1}-c)$.
  2. (2)
    If $p>\frac{1}{2\Delta}(1-a^{1})(a^{1}-c)$, we choose $a_{i}^{\prime}=a^{k-1}$:
    $\operatorname{\mathbb{E}}_{-i}\left[u_{i}(a^{k-1},a_{-i})-u_{i}(a^{k},a_{-i})\right]$ $\geq p\left[u_{i}(a^{k-1},a^{k}_{-i})-u_{i}(a^{k},a^{k}_{-i})\right]\geq p\tfrac{D}{n}(a^{k}-a^{k-1})(a^{k}-c)$
    $>\tfrac{D}{2n\Delta}(1-a^{1})(a^{1}-c)(a^{k}-a^{k-1})(a^{k}-c)$
    $\geq\frac{1}{2}u_{i}(a^{1}_{i},a^{1}_{-i})\cdot\frac{1}{\Delta}\min_{k}(a^{k}-a^{k-1})(a^{k}-c)$
    $>\frac{1}{2}u_{i}(a^{1},a^{1}_{-i})\cdot\frac{1}{\Delta}\min_{k}(a^{k}-a^{k-1})^{2}>0$
    In the first line, we ignored instances where all opponents might have played $a^{k-1}$ and used the inequality we derived above to show uniqueness of the NE. In the last line we used $a^{k}-c>a^{k-1}-c\geq a^{k}-a^{k-1}$.
Thus, for all mixed strategies of the opponents within the given support, there always exists an action $a_{i}^{\prime}$ for which the utility is strictly increased. As a consequence, $a^{k}$ is not stable under $R_{i}$ and the supports cannot be correlated rationalizable. ∎
Proposition A.7.
For a symmetric Bertrand competition game with linear demand, $\gamma<\beta$, and action space $\mathcal{A}=\{a^{1},\dots,a^{m}\}$ where $0\leq a^{1}\leq\dots,<a^{k^{*}}\leq p^{*}<a^{k^{*}+1}\dots<a^{m}$ with $p^{*}=\frac{\alpha+\beta c}{2\beta-\gamma}$, all correlated rationalizable sets are subsets of $\{a^{k^{*}},a^{k^{*}+1}\}$ for every player $i$.
Proof.
We will use the Nash equilibrium of the continuous game $p^{*}=\tfrac{\alpha+\beta c}{2\beta-\gamma}$ as a reference point and show that only the neighboring points of $p^{*}$, i.e., $a^{k^{*}}$ or $a^{k^{*}+1}$ can be stable under $R$. Using that the utility $u_{i}(a_{i},a_{-i})$ is linear in $a_{-i}$, we can write the difference in the utility between some action $a$ and $a^{\prime}$ for agent $i$ as
$\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a^{\prime},a_{-i})-u_{i}(a,a_{-i})\right]$ $=u_{i}(a^{\prime},\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[a_{-i}\right])-u_{i}(a,\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[a_{-i}\right])$
$=(\alpha-\beta a^{\prime}+\gamma\bar{a}_{-i})(a^{\prime}-c)-(\alpha-\beta a+\gamma\bar{a}_{-i})(a-c)$
$=(a^{\prime}-a)(\alpha+\beta c-\beta(a^{\prime}+a)+\gamma\bar{a}_{-i}),$
where $\bar{a}_{-i}:=\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}[\tfrac{1}{n-1}\sum_{j\neq i}a_{j}]$ is the expected average action of the opponents. Using the definition of $p^{*}$, we obtain
$u_{i}(a^{\prime},x_{-i})-u_{i}(a,x_{-i})=(a^{\prime}-a)\left(\beta(p^{*}-a^{\prime})+\beta(p^{*}-a)-\gamma(p^{*}-\bar{a}_{-i})\right).$ (9)
We show that deviating from an action $a\notin\{a^{k^{*}},a^{k^{*}+1}\}$ to one of these actions is always better if $a$ is the lowest or highest action of any support. Therefore, an action outside of $\{a^{k^{*}},a^{k^{*}+1}\}$ can never be correlated rationalizable. We have to consider two different cases: Given an action profile $a$
  1. (1)
    let agent $i$ be the one with the highest support up to action $a$ with $a>a^{k^{*}+1}>p^{*}$:
    $u_{i}(a^{k^{*}+1},x_{-i})-u_{i}(a,x_{-i})$ =(9)${{=}}(a^{k^{*}+1}-a)\left(\beta(p^{*}-a^{k^{*}+1})+\beta(p^{*}-a)-\gamma(p^{*}-\bar{a}_{-i})\right)$ )
    $\geq(a^{k^{*}+1}-a)(\beta(p^{*}-a^{k^{*}+1})+(\beta-\gamma)(p^{*}-a))$
    $\geq(a^{k^{*}+1}-a^{k^{*}+2})\beta(p^{*}-a^{k^{*}+1})>0$ (10)
  2. (2)
    let agent $i$ be the one with the lowest support up to action $a$ with $a<a^{k^{*}}\leq p^{*}$:
    $u_{i}(a^{k^{*}},x_{-i})-u_{i}(a,x_{-i})$ =(9)${{=}}(a^{k^{*}}-a)\left(\beta(p^{*}-a^{k^{*}})+\beta(p^{*}-a)-\gamma(p^{*}-\bar{a}_{-i})\right)$ )
    $\geq(a^{k^{*}}-a)(\beta(p^{*}-a^{k^{*}})+(\beta-\gamma)(p^{*}-a))$
    $\geq(a^{k^{*}}-a^{k^{*}-1})(\beta-\gamma)(p^{*}-a^{k^{*}-1})>0$ (11)
In summary, this shows that for any choice of, potentially mixed, opponent strategy with support of at most/at least $a$, there is an action that is better than $a$. Thus, $a$ is not stable under $R_{i}$ and cannot be part of any correlated rationalizable set. ∎
A.4 Convergence with Mean-based Algorithms
Recall that we defined the series of correlated rational response sets $\mathcal{A}^{0},\dots,\mathcal{A}^{k},\dots$ as $\mathcal{A}^{k}=R(\mathcal{A}^{k-1})$ for all $k\geq 1$ and $\mathcal{A}^{0}=\mathcal{A}$. Let $K$ be the number for which $\mathcal{A}^{K-1}_{i}$ is a strict subset of $\mathcal{A}^{K}_{i}$ for some $i$ and $\mathcal{A}^{K}=\mathcal{A}^{K+1}=\bar{\mathcal{A}}$. Because of remarks A.5 and A.1 this $K$ always exists. To prove theorem 4.1, we show that mean-based algorithms converge along this series of support sets with high probability, approaching $\mathcal{A}^{K}=\bar{\mathcal{A}}$ in the limit. Also, let $m$ be the maximum cardinality of any action space $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$.
A.4.1 Step 1: Individual Players Play Correlated Rational Responses
In the first step, we reason that individual players with mean-based algorithms will choose correlated rational responses. To do so, we look at the history of past actions and rewards $H_{t-1}$ of all players. If the opponents have mainly selected actions within $\mathcal{A}^{k}_{-i}$, we show that the player is likely to choose actions within $\mathcal{A}^{k+1}_{i}=R_{i}(\mathcal{A}^{k})$.
We denote the empirical frequency of opponents ($-i$) selecting actions in $\mathcal{A}^{k}_{-i}$ up to time $t$ by
$P_{i,t}(\mathcal{A}^{k}):=\frac{1}{t}\sum_{s=1}^{t}\mathbb{I}\left[\forall j\neq i:~{}a_{j,s}\in\mathcal{A}^{k}_{j}\right].$
Now we can state the proposition:
Proposition A.8.
Given $k\in\{0,\dots,K-1\}$ and $\gamma_{t}<\frac{\delta}{2}$ where $\delta$ is the competition constant of the game. If $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large for history $H_{t-1}$, then $\mathbb{P}_{i,t}(a|H_{t-1})\leq\gamma_{t}$ for all $a\not\in\mathcal{A}^{k+1}_{i}$.
By $\mathbb{P}_{i,t}(a)$ we denote the probability that player $i$ with $\gamma_{t}$-mean-based algorithm chooses action $a$ at time $t$.
Proof.
We say that the empirical frequency $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large if
$P_{i,t-1}(\mathcal{A}^{k})\geq\tilde{p}:=\frac{1}{2}\left(\frac{\Delta u}{\delta+\Delta u}+1\right)\in(0,1)$ (12)
where $\delta$ is the competition constant of the game and $\Delta u$ is the maximum absolute utility difference between any two actions of player $i$ for any opponent profile $a_{-i}$. Let $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ be an action which increases player $i$’s utility by at least $\delta$ compared to playing action $a\not\in\mathcal{A}^{k+1}_{i}$ if opponents play a mixed strategy with support on $\mathcal{A}^{k}_{-i}$. Consider the counterfactual advantage of action $a^{\prime}$ over $a$:
$\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a)$ $:=\frac{1}{t-1}\sum_{s=1}^{t-1}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})$
$=\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\not\in\mathcal{A}_{-i}^{k}\end{subarray}}\underbrace{u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})}_{\geq-\Delta u}+\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\in\mathcal{A}_{-i}^{k}\end{subarray}}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})$
$\geq(1-P_{i,t-1}(\mathcal{A}^{k}_{-i}))\cdot(-\Delta u)+P_{i,t-1}(\mathcal{A}^{k})\cdot\underbrace{\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a^{\prime},a_{-i})-u_{i}(a,a_{-i})\right]}_{>\delta}$
$>(1-P_{i,t-1}(\mathcal{A}^{k}))(-\Delta u)+\delta P_{i,t}(\mathcal{A}^{k})=P_{i,t-1}(\mathcal{A}^{k})(\delta+\Delta u)-\Delta u$
$\geq\tilde{p}(\delta+\Delta u)-\Delta u=\frac{1}{2}(2\Delta u+\delta)-\Delta u=\frac{\delta}{2}>\gamma_{t}$
The third line interprets the sum over all times for which the opponents play $a_{-i}\in\mathcal{A}^{k}_{-i}$ as an expectation over a joint action distribution $x_{-i}\in\Delta(\mathcal{A}^{k}_{-i})$ with support on $\mathcal{A}^{k}_{-i}$. We choose $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ such that it performs better than $a$ by a margin of at least $\delta$. For $a\not\in\mathcal{A}^{k+1}_{i}$, this action exists for any such mixed strategy, according to Equation (8). The last line implies that the probability $\mathbb{P}_{i,t}(a)$ that the mean-based algorithm chooses actions $a\not\in\mathcal{A}^{k+1}_{i}$ is small ($\leq\gamma_{t}$). ∎
A.4.2 Step 2: High-probability Induction
In the second step, we use the insights from Propositions A.8 to make a statement about all players. Our inductive argument is based on the following observation. Suppose all agents played actions in $\mathcal{A}^{k}$ at time $t$ with high frequency. Then, each of the competitors will tend to play actions within $\mathcal{A}^{k+1}$ in the next step, thus increasing $P_{i,t-1}(\mathcal{A}^{k})$ further (recall that $\mathcal{A}^{k+1}\subseteq\mathcal{A}^{k}$). At the same time, $P_{i,t-1}(\mathcal{A}^{k+1})$ will also increase. Eventually, $P_{i,t^{\prime}-1}(\mathcal{A}^{k+1})$ will exceed some threshold at time $t^{\prime}$ with high probability. This induces the next step towards $\mathcal{A}^{k+2}$, and so on, until the algorithms reach $\mathcal{A}^{K}$.
Instead of looking at $P_{i,t-1}(\mathcal{A}^{k})$ for individual players, we will now move on to looking at the frequency at which all players (not only the opponents) played within $\mathcal{A}^{k}$. We capture that this frequency is high by a series of events
$A_{k}:=\left[\frac{1}{T_{k}}\sum_{t=1}^{T_{k}}\mathbb{I}\left[\exists i:~{}a_{i,t}\not\in\mathcal{A}^{k}_{i}\right]\leq\beta\right]=\left[\frac{1}{T_{k}}\sum_{t=1}^{T_{k}}\mathbb{I}\left[\forall i:~{}a_{i,t}\in\mathcal{A}^{k}_{i}\right]\geq 1-\beta\right].$
These events are defined for some small $\beta>0$ and for a corresponding series of time steps $T_{0}<\dots<T_{k}<\dots<T_{K}$. In words, they capture that only in rare cases (fraction $\beta$ of all timesteps up to $T_{k}$), any agent chooses an action outside $\mathcal{A}^{k}$. Equivalently, this means that in most cases (fraction $1-\beta$), all agents played actions from $\mathcal{A}^{k}$.
We note the meaning of the event $A_{K}$ with respect to time-average convergence. For $\beta\to 0$, this event captures that all agents played within $\mathcal{A}^{K}$, i.e., the set of rationalizable actions, for a fraction of the time close to 1. If $\beta\to 0$ as $T_{K}\to\infty$, this implies time-average convergence. The following proposition makes a statement about the probability that this will happen, thus proving theorem 4.1.
Proposition A.9.
Assume that agents follow a $\gamma_{t}$-mean-based algorithm. For every sufficiently small $\beta>0$, there exists a rational constant $c>1$ such that
$\mathbb{P}(A_{K})\geq 1-\sum_{k=1}^{K}\sum_{j=1}^{J}\exp\left(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\right),$
where $T_{k}^{j}=c^{j+Jk}T_{0}$ and $T_{k}^{0}=T_{k}$ with a sufficiently large $T_{0}\in\mathbb{N}$ and $J=\lceil\log_{c}\tfrac{2}{\beta}\rceil$.
Moreover, $\mathbb{P}(A_{K})$ approaches 1 as $T_{0}\to\infty$.
Proof.
For a sufficiently small $\beta$, i.e., $0<\beta<1-\tilde{p}$ (as defined in Eq. 12), we choose the constant $c\in\mathbb{Q}$ such that $1<c<\tfrac{1-\beta}{\tilde{p}}$ and a sufficiently large $T_{0}$ such that $\gamma_{t}<\min(\tfrac{\delta}{2},\tfrac{\delta}{2\Delta u})$ for all $t\geq T_{0}$, where $\delta$ is the competition constant of the game (see Section A.2) and $\Delta u$ is the maximium absolute utility difference between any two action profiles. Additionally, (for technical reasons) we choose $T_{0}$ such that $T^{j}_{k}=c^{j+Jk}T_{0}\in\mathbb{N}$ for all $j=0,\dots,J$ and $k=0,\dots,K-1$.
Before we can prove the result, we have to introduce some sub-events we use in the following:
$B_{k}^{j}$ $:=\left[\frac{1}{T^{j}_{k}-T^{j-1}_{k}}\sum_{t=T^{j-1}_{k}+1}^{T^{j}_{k}}\mathbb{I}[\exists i:~{}a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]\leq\frac{\beta}{2}\right],$ (13)
$A_{k}^{j}$ $:=(A_{0},\dots,A_{k},B_{k}^{1},\dots,B_{k}^{j}).$ (14)
Note that $B_{K}^{j}$ is a somehow stricter version of the events $A_{k}$ since $\tfrac{\beta}{2}<\beta$ and $\mathcal{A}^{k+1}\subseteq\mathcal{A}^{k}$. Using these definitions and some intermediate results, which we will prove afterwards, we get:
$\mathbb{P}(A_{K})$ $\geq\mathbb{P}(A_{K},A_{K-1},\dots,A_{1},A_{0})$
$=\underbrace{\mathbb{P}(A_{0})}_{=1}\cdot\prod_{k=1}^{K}\mathbb{P}(A_{k+1}|A_{k},\dots,A_{0})$ (Chain Rule)
$\geq\prod_{k=1}^{K}\mathbb{P}(A_{k}^{J}|A_{k},\dots,A_{0})$ (Lemma A.1)
$=\prod_{k=1}^{K}\mathbb{P}(B_{k}^{1},\dots,B_{k}^{J}|A_{k},\dots,A_{0})$ (Equation 14)
$=\prod_{k=1}^{K}\prod_{j=1}^{J}\mathbb{P}(B_{k}^{j}|A_{k},\dots,A_{0},B_{k+1}^{1},\dots,B_{k+1}^{j-1})$ (Chain Rule)
$=\prod_{k=1}^{K}\prod_{j=1}^{J}\mathbb{P}(B_{k}^{j}|A_{k}^{j-1})$ (Equation 14)
$\geq\prod_{k=1}^{K}\bigg{[}1-\sum_{j=1}^{J}\exp\bigg{(}\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg{)}\bigg{]}$ (Lemma A.2)
$\geq 1-\sum_{k=1}^{K}\sum_{j=1}^{J}\exp\bigg{(}\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg{)}$ (Weierstrass Product Inequality)
Since $(T^{j}_{k}-T^{j-1}_{k})=(c-1)c^{j-1+Jk}T_{0}$ the term in the last line converges to $1$ for $T_{0}\rightarrow\infty$. ∎
In the remaining part of this subsection, we will show the necessary steps used in the proof of Proposition A.9.
Lemma A.1.
$A_{k}^{J}$ implies $A_{k+1}$, which means $\mathbb{P}(A_{k+1})\geq\mathbb{P}(A_{k}^{J})$
Proof.
If $A_{k}^{J}$ holds we know that $B^{1}_{k},\dots,B^{J}_{k}$ hold, which means that the agents have played actions outside of $\mathcal{A}^{k+1}$ for a fraction of at most $\frac{\beta}{2}$ during $\{T_{k}^{0}+1,\dots,T_{k}^{1}\},\dots,\{T_{k}^{J-1}+1,\dots,T_{k}^{J}\}$. Recall that $T_{k}^{0}=T_{k}$ and $T_{k}^{J}=T_{k+1}$. Using this, we can bound the relevant sum for $A_{k+1}$:
$\frac{1}{T_{k+1}}\sum_{t=1}^{T_{k+1}}\mathbb{I}[\exists i:a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]$ $=\frac{1}{T_{k+1}}\sum_{t=1}^{T_{k}}\mathbb{I}[\exists i:a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]+\frac{1}{T_{k+1}}\sum_{t=T_{k}+1}^{T_{k+1}}\mathbb{I}[\exists i:a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]$
$\leq\frac{T_{k}}{T_{k+1}}+\frac{1}{T_{k+1}}\sum_{j=1}^{J}(T_{k}^{j}-T_{k}^{j-1})\frac{\beta}{2}=\frac{T_{k}}{T_{k+1}}+\frac{T_{k+1}-T_{k}}{T_{k+1}}\frac{\beta}{2}$
$=c^{-J}+(1-c^{-J})\frac{\beta}{2}\leq c^{-\log_{c}\left(\frac{2}{\beta}\right)}+\frac{\beta}{2}=\beta.$
Therefore $A_{k}^{J}$ implies $A_{k+1}$. ∎
Lemma A.2.
$\mathbb{P}(B_{k}^{j}|A_{k}^{j-1})\geq 1-\exp\bigg{(}\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg{)},\qquad\qquad j=1,\dots,J.$
Proof.
Recall that $B_{k}^{j}$ defined in Equation (13) is the event, that agents play only a small fraction outside of $\mathcal{A}^{k+1}$ during the time steps $\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}$ given some $j\in J$. To get a bound on the probability, we first introduce random variables that count if an agent plays outside this set:
$X_{t}:=\mathbb{I}\left[\exists i:~{}a_{i,t}\not\in\mathcal{A}^{k+1}_{i}\right],\qquad t\in\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}.$
We can show that the expectation of $X_{t}$ conditioned on a history $H_{t-1}$ such that $A_{k}^{j-1}$ holds can be bounded by
$\operatorname{\mathbb{E}}[X_{t}|A_{k}^{j-1},H_{t-1}]$ $=\mathbb{P}(\exists i:~{}a_{i,t}\not\in\mathcal{A}^{k+1}_{i}|A_{k}^{j-1},H_{t-1})$
$\leq n\cdot\mathbb{P}(a_{i,t}\not\in\mathcal{A}^{k+1}_{i}|A_{k}^{j-1},H_{t-1})$ (Union Bound Agents)
$\leq n\cdot m\cdot\mathbb{P}(a_{i,t}=a\notin A^{k+1}_{i}|A_{k}^{j-1},H_{t-1})$ (Union Bound Actions)
$\leq n\cdot m\cdot\gamma_{t}\leq n\cdot m\cdot\gamma_{T_{k}^{j-1}}.$ (Proposition A.8)
Note that Proposition A.8 is applicable because of Lemma A.3 (below). Recall that $m$ was the number of elements in the largest action set, so this bound holds for any player. Using this upper bound, we can use the random variables $X_{t}$ to define
$Z_{t}:=\sum_{s=T_{k}^{j-1}+1}^{t}\left(X_{s}-nm\gamma_{T_{k}^{j-1}}\right),\qquad t\in\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}.$ (15)
With $Z_{T_{k}^{j-1}}:=0$ and under the condition of $A_{k}^{j-1}$, the value of $Z_{t}$ decreases with $t$ in expectation as we always subtract an upper bound of the expectation of each summand $X_{s}$, i.e., $\operatorname{\mathbb{E}}[Z_{t}|Z_{t-1},A_{k}^{j-1},H_{t-1}]\leq Z_{t-1}$. This makes $Z_{t}$ a super-martingale and allows us to apply Azuma’s inequality.
First, we choose $T_{0}$ sufficiently large, such that $\gamma_{T_{k}^{j-1}}\leq\beta\tfrac{1}{4nm}$, where $n$ is the number of agents and $m$ is the maximum size of any action space. This gives us $|Z_{t}-Z_{t-1}|=|X_{t}-nm\gamma_{T_{k}^{j-1}}|\leq 1$ since $X_{t}\in\{0,1\}$ and $0\leq nm\gamma_{T_{k}^{j-1}}\leq\beta\tfrac{nm}{4nm}\leq 1$. By Azuma’s inequality we get for all $\varepsilon>0$
$\mathbb{P}(Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}\geq\epsilon|A_{k}^{j-1},H_{t-1})\leq\exp\bigg{(}\frac{-\epsilon^{2}}{2(T^{j}_{k}-T^{j-1}_{k})}\bigg{)}$
[…]
Now we choose $\varepsilon=\frac{\beta}{4}(T^{j}_{k}-T^{j-1}_{k})$ such that we end up with the event described by $B_{k}^{j}$:
$Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}$ $<\varepsilon$
$\Leftrightarrow\quad\sum_{t=T_{k}^{j-1}+1}^{T_{k}^{j}}\left(\mathbb{I}\left[\exists i:~{}a_{i,t}\not\in\mathcal{A}^{k+1}_{i}\right]-nm\gamma_{T_{k}^{j-1}}\right)$ $<\frac{\beta}{4}(T^{j}_{k}-T^{j-1}_{k})$
$\Leftrightarrow\quad\frac{1}{T^{j}_{k}-T^{j-1}_{k}}\sum_{t=T_{k}^{j-1}+1}^{T_{k}^{j}}\mathbb{I}\left[\exists i:~{}a_{i,t}\not\in\mathcal{A}^{k+1}_{i}\right]$ $<nm\gamma_{T_{k}^{j-1}}+\frac{\beta}{4}$
Since $\gamma_{T_{k}^{j-1}}\leq\beta\tfrac{1}{4nm}$, the right hand side is less or equal to $\frac{\beta}{2}$ and satisfies thereby $B_{k}^{j}$. Finally, we have
$\mathbb{P}(B_{k}^{j}|A_{k}^{j-1})\geq\mathbb{P}(Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}\geq\tfrac{\beta}{4}(T^{j}_{k}-T^{j-1}_{k})|A_{k}^{j-1},H_{t-1})\geq 1-\exp\bigg{(}\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg{)}.$
To apply Proposition A.8 in the proof of Lemma A.2, we need to make sure that the assumptions hold. This is ensured by the following lemma.
Lemma A.3.
If $A_{k}^{j-1}$ holds, the condition $P_{t-1}^{i}(\mathcal{A}^{k})>\tilde{p}$ holds for all $i\in\mathcal{N}$, $t\in\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}$, and any $j=1,\dots,J$.
Proof.
Recall that given $\tilde{p}$ as defined in Eq.(12), the constants $\beta$ and $c$ in the previous proofs were chosen such that $0<\beta<1-\tilde{p}$ and $1<c<\tfrac{1-\beta}{\tilde{p}}$.
$P^{i}_{t-1}(\mathcal{A}^{k})$ $=\frac{1}{t-1}\sum_{s=1}^{t-1}\mathbb{I}[\forall j\neq i:a_{j,s}\in\mathcal{A}^{k}_{j}]$
$=\frac{1}{t-1}\bigg{(}\sum_{s=1}^{T_{k}}\mathbb{I}[\forall j\neq i:a_{j,s}\in\mathcal{A}^{k}_{j}]+\sum_{s=T_{k}+1}^{T_{k}^{j-1}}\mathbb{I}[\forall j\neq i:a_{j,s}\in\mathcal{A}^{k}_{j}]+\sum_{s=T_{k}^{j-1}+1}^{t-1}\mathbb{I}[\forall j\neq i:a_{j,s}\in\mathcal{A}^{k}_{j}]\bigg{)}$
$\geq\frac{1}{t-1}\bigg{(}\underbrace{\sum_{s=1}^{T_{k}}\mathbb{I}[\forall j:a_{j,s}\in\mathcal{A}^{k}_{j}]}_{\text{bounded using }A_{k}}+\underbrace{\sum_{s=T_{k}+1}^{T_{k}^{j-1}}\mathbb{I}[\forall j:a_{j,s}\in\mathcal{A}^{k}_{j}]}_{\text{bounded using }(B_{k}^{1},\dots,B_{k}^{j-1})}+\underbrace{\sum_{s=T_{k}^{j-1}+1}^{t-1}\mathbb{I}[\forall j:a_{j,s}\in\mathcal{A}^{k}_{j}]}_{\geq 0}\bigg{)}$
$\geq\frac{1}{t-1}\big{(}T_{k}\cdot(1-\beta)+(T_{k}^{j-1}-T_{k})(1-\tfrac{\beta}{2})\big{)}$
$>\frac{1}{t-1}T_{k}^{j-1}(1-\beta)\geq\frac{T_{k}^{j-1}}{T_{k}^{j}}(1-\beta)=\tfrac{1-\beta}{c}>\tilde{p}.$
In the last line, we used $1-\tfrac{\beta}{2}>1-\beta$ and $t-1<T_{k}^{j}$. ∎
Appendix B Additional Experiments
B.1 Online Optimization Algorithms
In what follows, we provide a brief summary of the online optimization algorithms used in our paper.
B.1.1 $\epsilon$-Greedy Bandit
One of the simplest algorithms is the $\epsilon$-greedy exploration strategy. With its internal belief, it keeps track of the empirical mean reward $\bar{u}_{t}(a)$ of each explored action $a$. With probability $1-\epsilon$, it selects the most promising action in the current time step, assuming that each action has been chosen at least once. With probability $\epsilon$, it chooses an arbitrary action with uniform probability.
$\pi_{t}(a_{t})=\begin{cases}1-\epsilon\cdot\frac{K-1}{K}&a_{t}=\arg\max_{a\in\mathcal{A}}\bar{u}_{t-1}(a)\\ \frac{\epsilon}{K}&\text{else}\end{cases}$ (16)
B.1.2 UCB Algorithms
Like $\epsilon$-greedy, the UCB algorithms estimate and update the empirical mean reward of all their actions. They balance exploration and exploitation by choosing arms according to a heuristic, the upper confidence bound (UCB) of estimated rewards. This value can be computed by adding an exploration bonus $\rho_{t}$ to the empirical mean of past rewards $\bar{u}_{t}$:
$\text{UCB}_{t}(a)=\bar{u}_{t}(a)+\rho_{t}(a)$ (17)
The exploration bonus may depend on past rewards, $u_{s}(a_{s}),~{}s\leq t$, the number of times an action was selected, $n_{t}(a)$, and the time, $t$. In each step, the algorithm chooses the action $a_{t}$ which maximizes $\text{UCB}_{t-1}(a)$. Similar to Hansen et al. (2021), we use the UCB-Tuned algorithm (Auer et al. (2002)) in our experiments. This algorithm is a refined version of UCB1 (Auer et al. (2002)) with an exploration bonus
$\rho_{t}^{UCB-Tuned}(a)=\sqrt{\frac{\ln n}{n_{t}(a)}\min\left\{\dfrac{1}{4},V_{t}(a)\right\}}$
where $n$ is the number of plays so far and $V_{t}(a)$ is a variance estimate
$V_{t}(a)=\left(\frac{1}{n_{t}(a)}\sum_{\begin{subarray}{c}1\leq s\leq t:\\ a_{s}=a\end{subarray}}u_{s}(a)^{2}\right)-\bar{u}_{t}(a)^{2}+\sqrt{\frac{2\ln t}{n_{t}(a)}}.$
Unlike Hansen et al. (2021), our implementation of the algorithm does not eliminate any actions that it deems unpromising, and it resolves ties of the UCB heuristic randomly. The algorithm has vanishing regret in the stochastic multiarm bandit problem under mild assumptions (see, e.g., Lattimore and Szepesvári (2020)) but may suffer linear regret in the adversarial setting.
B.1.3 Exp3 Algorithms
Exp3 ("Exponential-weight algorithm for Exploration and Exploitation") is an algorithm introduced by Auer et al. (1995), which was designed for the adversarial model. This means that we do not make any assumptions about the distribution of rewards. In contrast, the rewards could also be chosen by an adversary who has full information about the setting and the agent’s algorithm and who chooses rewards in each round.
The Exp3 algorithm’s action-selection rule assigns weights to each arm, and its update rule adjusts these weights based on the observed rewards. Following the slightly modified version by Lattimore and Szepesvári (2020), our implementation of the Exp3 algorithm samples its actions $a_{t}$ from
$P_{t}(a)=(1-\epsilon_{t})\cdot\frac{\exp\left(\eta\sum_{s=1}^{t-1}\hat{u}_{s}(a)\right)}{\sum_{\hat{a}\in\mathcal{A}}\exp\left(\eta\sum_{s=1}^{t-1}\hat{u}_{s}(\hat{a})\right)}+\frac{\epsilon_{t}}{\lvert\mathcal{A}\rvert}.$ (18)
Here, $\epsilon_{t}$ is an additional exploration rate that encourages the algorithm to choose arbitrary actions from time to time and $\hat{u}_{s}(a)$ is an estimate of the reward. Let $P_{s}(a)$ be the probability that the algorithm chooses action $a$ at time $s$, given the history until then. Assuming rewards are bounded within $[0,1]$, we choose
$\hat{u}_{s}(a)=1-\frac{\mathbb{I}\{a_{s}=a\}}{P_{s}(a)}\cdot(1-u_{s}).$ (19)
The Exp3 algorithm is known to be no-regret for some choices of $\epsilon_{t}$ in the adversarial setting (Auer et al. (1995), Lattimore and Szepesvári (2020)) and mean-based for a specific choice of exploration rate $\epsilon_{t}$ and learning rate $\eta_{t}$ (Braverman et al. 2018). Our implementation relies on a fixed exploration and learning rate for simplicity.
B.1.4 Thompson Sampling
Thompson Sampling is a Bayesian approach to multi-armed bandit algorithms that was named after William R. Thompson based on his papers from 1933 and 1935. The heuristic maintains probability distributions over the rewards of every possible action and uses these to determine the next action.
More specifically, Thompson sampling assumes that the rewards of each action $a\in\mathcal{A}$ are sampled independently from a probability distribution $\mathbb{P}(u_{t}=r|\theta_{a})$. [^5] This also generalizes to probability densities. For simplicity, we stick to the discrete formulation in this introduction. Bayes’ rule allows us to maintain a posterior distribution to update our beliefs on the (unknown) parameters $\theta_{a}$, and consequently on the distribution of rewards of each action:
$\mathbb{P}(\theta_{a}|\{u_{s}\}_{s=1}^{t},\theta_{a}^{0})=\frac{\mathbb{P}\left(\left.\{u_{s}\}_{1\leq s\leq t:a_{s}=a}\right|\theta_{a}\right)\cdot\mathbb{P}(\theta_{a}|\theta_{a}^{0})}{\mathbb{P}\left(\{u_{s}\}_{1\leq s\leq t:a_{s}=a}\right)}$
Here, $\mathbb{P}(\theta_{a}|\theta_{a}^{0})$ is a prior distribution over reward parameters for action $a$, and ($\mathbb{P}\left(\left.\{u_{s}\}_{1\leq s\leq t:a_{s}=a}\right|\theta_{a}\right)$ is the probability of rewards $u_{s}$ at iterations for which the selected action was $a$.
In each iteration, Thompson sampling draws parameter samples $\theta_{a}$ for all actions from their corresponding posterior distribution, introducing randomness. Subsequently, the action with the highest expected reward, based on $\theta_{a}$, is chosen. After receiving reward feedback, the posterior distribution of the chosen action is updated according to Bayes’ rule.
While Thompson (1933) described the approach for two actions with Bernoulli distributions, our setting has $K$ actions with continuous, bounded rewards. The Bayesian approach can also be used in this setting; a generalization that has been described, e.g., by Russo et al. (2020). Our implementation of Thompson sampling assumes Gaussian-distributed rewards [^6] We note that this does not reflect that the rewards are bounded. The approach is not severely limited by this. and poses a Gaussian distribution over the parameters as conjugated prior.
B.2 Bertrand Duopolies
We also ran experiments with another variant of the UCB algorithms described by Auer et al. (2002), UCB1, and with Q-Learning (QL). Our implementation of QL uses an $\epsilon$-greedy exploration strategy with an exponentially decaying exploration rate $\epsilon$, which seems to be a necessary condition for the evolution of supra-competitive pricing. For all experiments where QL was involved, we iterated 3 Mio. steps (instead of 250,000), and we provided the actions of both players from the previous iteration as the state.
Figure 9 shows the price-collusion indices of all algorithm configurations. We observe that, in line with what we observed in our main experiments, collusion mostly evolves when two algorithms of the same type are paired. The combination of UCB-T and UCB1 deviates from this pattern and displays a significant level of collusion. It is likely that these two algorithms can learn to "coordinate" (unconsciously) because of their similarity, which shows that algorithms need not share the exact same implementation and parameters in order to collude.
Figure 9: Price-collusion indices in duopoly O3 (logit demand).
We display the price-collusion index by agent as an average value over all runs. To determine the index, we use the median price over the last 1000 iterations.
For the Q-learning experiment, Figure 10 shows the evolution of collusion over the experiment iterations. We observe that collusion evolves late and abruptly, which can likely be attributed to the exploration rate becoming sufficiently small.
Figure 10: Evolution of collusion with QL.
We display the price-collusion index of a QL-duopoly over time. To determine the index, we use the median price over 1000 iterations.
B.3 Collusion Results by Demand Function
In the main part of the paper, we averaged the collusion indices of the duopolies for all demand functions. Here, we provide the results for the individual demand functions. Figures 11 and 12 display the price-collusion indices for all five games. We find that supra-competitive prices evolve consistently with the UCB-T algorithm throughout all settings, with particularly pronounced collusion in the standard demand setting (O1). Likewise, prices remain mostly competitive for the remaining algorithm combinations.
(a) O1.
(b) O2.
(c) O3.
Figure 11: Price-collusion indices in symmetric games.
We visualize the price-collusion indices for the symmetric games and several combinations of algorithms. The displayed numbers are the numeric values of the collusion indices; the values in brackets display their corresponding standard deviation. Averages and standard deviations were computed over two agents and 10 runs.
(a) O2’.
(b) O3’.
Figure 12: Price-collusion indices in asymmetric games.
We visualize the price-collusion indices for the asymmetric games and several combinations of algorithms. The displayed numbers are the numeric values of the collusion indices; the values in brackets display their corresponding standard deviation. Averages and standard deviations were computed over two agents and 10 runs.