### v2

# Online Optimization Algorithms in Repeated Price Competition: Equilibrium Learning and Algorithmic Collusion

Abstract
This paper examines whether widely used online learning algorithms in pricing can independently reach competitive outcomes or instead foster tacit collusion. This issue has drawn considerable attention from competition regulators as algorithmic pricing becomes more common in digital markets. Understanding when such algorithms lead to equilibrium prices or to supra-competitive prices is critical for buyers, sellers, and policymakers.
We study the behavior of multi-armed bandit algorithms in repeated price competition. These algorithms only observe profits from the chosen prices, making them realistic models of automated pricing. Our formal analysis shows that an important class of online learning algorithms, called mean-based algorithms, reliably converges to Nash equilibrium in Bertrand competition. This finding is notable because, generally, online learning algorithms do not guarantee convergence. We also run extensive numerical experiments with different bandit algorithms, confirming that most widely used algorithms, including those not mean-based, converge to equilibrium. We observe supra-competitive prices only in specific cases where all sellers implement the same symmetric version of certain algorithms, such as UCB or Q-learning, and this effect diminishes as the number of competitors increases.
Our results highlight that the risk of algorithmic collusion in competitive markets is often overstated. For most practical implementations of bandit algorithms, sellers’ prices converge to competitive levels. Only under very specific and symmetric setups do prices remain above competitive benchmarks, and this effect diminishes with more competitors. These insights support regulators concerned with consumer welfare and managers considering algorithmic pricing tools. They suggest that while vigilance is warranted, fears of widespread algorithm-driven collusion may be exaggerated.
Contents
1 Introduction
Algorithmic pricing, where prices are determined by software agents, is becoming increasingly prevalent in online retail markets. chen2016empirical estimated for 2015 that algorithms were involved in the pricing of approximately one-third of the roughly 1600 best-selling products on the Amazon marketplace. In 2018, the average product price on Amazon was reported to change once every ten minutes. [^1] https://www.businessinsider.com/amazon-price-changes-2018-8 Since then, an industry has emerged that specializes in automated pricing software with over 130 companies offering a variety of 185 pricing algorithms in 2021 (calzolari_pricing_2024).
From the point of view of a single seller, algorithmic pricing aims to solve an online optimization problem, where a firm’s actions are the prices it sets, and the objective is the (average) profit it wants to maximize over time. When entering the market, sellers have little information about competitors’ costs or the demand of buyers at different prices and need to find out over time which prices maximize profit.
A key challenge is deciding whether to prioritize short-term revenue (by exploiting a known price with a high payoff) or invest in discovering better long-term pricing strategies (through exploration of different prices). Online optimization algorithms are designed to effectively balance this trade-off, and they scale effectively across large sets of products (bubeck2011introduction). In online exchanges, such optimization algorithms have access to bandit feedback. This means that after performing an action (setting a price), an agent learns the value of the objective function (his profit) for this specific action. This (multi-armed) bandit model in online optimization is particularly suitable for algorithmic pricing applications, which was recognized early on. Bandit algorithms were already suggested for pricing by rothschild1974two long before digital platform markets emerged. Today, there is an extensive academic literature on multi-armed bandit algorithms for pricing (trovo2015multi, den2015dynamic, bauer2018optimal, mueller2019low, elreedy2021novel, taywade_multi-armed_2023, qu24) and there are many resources by practitioners on how to implement bandit algorithms for pricing. [^2] https://towardsdatascience.com/dynamic-pricing-with-multi-armed-bandit-learning
-by-doing-3e4550ed02ac
, https://www.griddynamics.com/blog/dynamic-pricing-algorithms
Online optimization algorithms target online problems where agents optimize against a stochastic process that is unknown and independent of their actions. Game-theoretical problems are different because the actions of one player impact the objectives of the others. In games, the Nash equilibrium (NE) describes a state where no agent wants to deviate unilaterally. Unlike the optimum in single-player optimization, this state may bring suboptimal payoffs to all agents. In spite of a long literature on learning in games, we don’t have a comprehensive theory about which algorithms converge to a Nash equilibrium in which types of games (fudenberg1999TheoryLearningGames, young2004strategic, cesa2006prediction).
Algorithmic pricing is a game-theoretical setting. An online retail market where multiple sellers compete via prices in a market for homogeneous goods is naturally modeled as a Bertrand competition (bertrand1883theorie). This established oligopoly model allows for different assumptions about consumer demand (e.g., all-or-nothing, linear, or logit demand models), and there is a long literature characterizing the prices that emerge in equilibrium. In a Bertrand competition with homogeneous products and all-or-nothing demand, the equilibrium price equals marginal cost, which leads to the maximization of consumer welfare.
Equilibrium analysis assumes rational agents that play their equilibrium strategy from the start. In practice, however, competitors lack information about each other’s costs and the demand model, which are both needed to calculate such a price in advance. Additionally, changes in demand and supply over time would require players to recalculate their equilibrium strategy.
A growing number of articles shows that optimization and learning algorithms can lead to supra-competitive prices higher than the Nash equilibrium of the stage game waltman2008q, calvano_algorithmic_2019, abada_artificial_2023, klein_autonomous_2021, abada_collusion_2024, brown_competition_2023.333Note that by the Folk Theorem, such a supra-competitive price could still be an equilibrium in an infinitely repeated game, if players have sufficient patience (fudenberg1991game). The phenomenon is commonly referred to as algorithmic collusion (den2023mathematical), and it raised substantial concerns among regulators (oecd2017) as it decreases consumer welfare. Meanwhile, some states start to regulate the use of pricing algorithms, especially if they are used jointly between competitors (aguiar-curry_ab-325_2025, vinson__elkins_llp_california_nodate). abada2024algorithmic provide an up-to-date treatment of the subject, an overview of relevant literature, and policy implications. They define algorithmic collusion as persistent supra-competitive outcomes produced by learning algorithms without human design to produce those outcomes, and we adhere to this definition.
The discussion on algorithmic collusion draws largely on numerical experiments of certain learning algorithms in a repeated Bertrand competition with a fixed set of sellers and a specific demand model. While real-world environments might be more complex, this model enables an analysis of convergence to the static Nash equilibrium. In particular, analysts can perform comparative statics based on the knowledge of costs and demand functions that are not available in empirical work. One might argue that if algorithmic collusion does not arise in repeated play in a Bertrand pricing game, it is unlikely to emerge in more complex environments where demand and supply change over time. There are a number of environments where algorithmic collusion was shown experimentally. Most articles focus on Q-learning (Calvano.2020b, calvano_algorithmic_2021-1, klein_autonomous_2021). But hansen_frontiers_2021 recently found evidence for supra-competitive prices with Upper Confidence Bound (UCB, auer2002bandit), a multi-armed bandit algorithm, as well. Reward-punishment schemes that might be the result of reinforcement learning algorithms with large state spaces can be ruled out in these environments (lambin2024less). We argue that it is important to study not only properties of algorithms but also to consider the structure of the games at hand when analyzing the convergence to equilibrium.
Decades of research on learning in games (young2004strategic, foster1997CalibratedLearningCorrelated) has shown that convergence of learning algorithms to a Nash equilibrium typically requires strong assumptions. It is well-known that there are games that cannot be learned via any uncoupled learning dynamics (hart2006stochastic, milionis2022nash). Also, random normal-form games can lead to cycles or even chaotic dynamics (sanders2018prevalence). The Bertrand competition has more structure that we exploit in this paper to show the convergence of large classes of algorithms relevant to algorithmic pricing.
1.1 Contributions
Prior studies on algorithmic pricing and algorithmic collusion analyzed specific algorithms in a repeated Bertrand competition with specific demand models. We make two main contributions: First, we prove that an important class of online optimization algorithms, mean-based algorithms, converge to correlated rationalizable strategies (brandenburger1987rationalizability), a solution concept that contains correlated equilibria (aumann1987correlated) and Nash equilibria. Then, we show that in symmetric Bertrand competition with all-or-nothing or linear demand and discrete actions, mean-based algorithms converge almost surely to actions that are close to their Nash equilibria. This is because the set of correlated rationalizable strategies is equivalent or close to the set of Nash equilibria of these games. The connection between mean-based algorithms and correlated rationalizable strategies was unknown, and the finding adds to the literature on learning in games. The result has real-world relevance because it provides proof that widely used algorithms such as Exponential Weights (Exp3) converge to a Nash equilibrium in important models of pricing competition. Note that such convergence results of bandit algorithms to Nash equilibrium are rare in the literature on learning in games.
Not all online optimization algorithms suitable for algorithmic pricing are mean-based. In a second contribution, we provide extensive experiments with a variety of multi-armed bandit algorithms that have been suggested for algorithmic pricing and show that algorithmic collusion is an exception limited to symmetric installations of UCB algorithms and Q-learning among a small number of sellers. We report experiments on Exp3, $\varepsilon$-greedy, UCB, and Thompson sampling, showing that UCB leads to high prices while the other algorithms converge to the Nash equilibrium. This holds for all-or-nothing, linear, and logit demand models for symmetric and asymmetric model parameterizations. In experiments in which two or more different algorithms (including UCB) are combined, we find that prices converge to the Nash equilibrium quickly. Algorithmic collusion with UCB is also much reduced if there are more than three sellers. Overall, these results indicate that for the important class of online optimization algorithms, algorithmic collusion is less of a concern.
1.2 Organization of the Paper
In the next section, we discuss related work on online optimization, algorithmic collusion, and learning in games. In Section 3, we introduce the relevant game-theoretical solution concepts and the Bertrand competition. Section 4 provides the central theoretical result of our paper on the convergence of mean-based algorithms to equilibrium, before we describe our experimental results in Section 5. Section 6 provides conclusions and an outlook.
2 Related Work
In what follows, we provide an overview of the relevant literature. We briefly discuss online optimization algorithms, summarize the literature on algorithmic collusion, and the key results from the literature on learning in games relevant to this paper.
2.1 Online Optimization
Online optimization is concerned with making sequential decisions in an unknown environment with the goal of optimizing a performance metric over time. Let us briefly introduce the basic model of online optimization. At each stage $t=1,2,\dots$ the agent chooses a strategy $x_{t}\in\mathcal{X}$ from some set and gets a utility $u_{t}(x_{t})$. In algorithmic pricing, this action could be the price or the quantity. It is usually assumed that $\mathcal{X}$ is a closed convex subset of some vector space and $u_{t}$ are convex functions (shalev-shwartzOnlineLearningOnline2011).
Algorithms are commonly analyzed in two models, the adversarial model and the stochastic model. In the stochastic model, the objective goal is to minimize the expected regret over the distribution of the input, which is drawn independently and identically from some underlying distribution. In the adversarial model, the input can be chosen by an adversary. A standard performance measure of algorithms generating a sequence of strategies $x_{t}$ in both models is the notion of (external) regret.
Definition 1 (No-regret Algorithm).
The (external) regret denotes the difference between the aggregated utilities after $T$ stages and the best fixed action in hindsight and is defined by
$\text{Reg}(T)=\max_{x\in\mathcal{X}}\sum_{t=1}^{T}u_{t}(x)-u_{t}(x_{t}).$ (1)
We say that an algorithm has no regret if the regret $\text{Reg}(T)$ grows sublinearly. This means that an algorithm performs at least as well as the best fixed action in hindsight.
There are different types of no-regret online algorithms. One can distinguish such algorithms based on the feedback available to the agents. Some noteworthy feedback types include:
  • Bandit feedback: The algorithm only receives partial feedback about the performance of its decisions. Typically, the feedback consists of a scalar reward or cost signal associated with the chosen action, but it does not reveal the potential rewards or costs of the other actions.
  • Gradient feedback: The algorithm receives feedback about the first-order derivatives (e.g., gradients) of the cost or payoff function with respect to the chosen action. Algorithms that have access to first-order oracles are often more efficient and allow for the implementation of fast algorithms. While such feedback is typically not available to learning algorithms in the field, one can compute gradient feedback when the goal is to develop equilibrium solvers rather than to mimic real-world interactions.
  • Full feedback: The algorithm receives feedback on the reward of all its possible actions or is granted access to the entire reward function. This information is given independent of which action was selected, thus providing access to counterfactual information.
We will largely focus on bandit feedback, which is a natural algorithm design principle for algorithmic pricing. Sellers set a price, and they learn their profit for this price in the next period. They often have no or only incomplete information about the demand model or the strategies used by all other sellers. This is particularly true on large online retail platforms where there are many substitutes for a good. We also include experiments with full feedback, and our theoretical results cover both bandit-feedback and full-feedback algorithms.
Exponential Weights, Follow-the-Perturbed-Leader (FTPL), Online Gradient Descent, or Online Mirror Descent are all algorithms that satisfy the no-regret property (which does not make any assumptions on the feedback model). Exponential Weights (or the Exp3 variant) uses bandit feedback and updates the weights associated with each action based on the cumulative reward observed for that action. The algorithm then chooses actions with probabilities proportional to their weights.
Exp3 is also a mean-based algorithm (braverman_selling_2018), a property of online optimization algorithms that will play an important role in our analysis. Informally, if the mean reward of action $a$ is significantly larger than the mean reward of action $b$, the learning algorithm will choose action $b$ with negligible probability. Apart from Exp3, also Multiplicative Weights Update (MWU), and FTPL are mean-based. We will show that such algorithms converge to a Nash equilibrium in versions of the Bertrand competition.
Reinforcement learning (RL) algorithms have also been applied to algorithmic pricing (rana2014real, kastius2022dynamic, deng2024algorithmic). RL with states can be appropriate in environments where demand systematically depends on state variables, such as the day of the week or customer history (abada_artificial_2023). However, these algorithms are typically sample-inefficient: as the state and action space grow, they require extensive training rounds and adapt slowly to changing market conditions. This makes them less suitable for dynamic pricing environments, where rapid adaptation is crucial. By contrast, online learning algorithms with bandit feedback can adjust more quickly and scale better. Prior work on algorithmic collusion has often applied Q-learning, a simple RL technique, but with the state space restricted to the current price, it essentially resembles an $\varepsilon$-greedy bandit algorithm (see also lambin2024less, who shows collusion can arise even in stateless Q-learning). Based on these arguments and discussions with practitioners, we therefore focus on online optimization algorithms with bandit feedback in our work, while reporting Q-learning benchmarks for comparison in the online supplement (Section B.6).
2.2 Algorithmic Collusion
According to harrington_developing_2018, “[c]ollusion is when firms use strategies that embody a reward–punishment scheme which rewards a firm for abiding by the supracompetitive outcome and punishes it for departing from it.” While this definition states that a reward-punishment scheme is a necessary mechanism for genuine collusion, abada_collusion_2024 argue that “[h]arm to consumers comes from supracompetitive prices, not the policy functions that produce those prices nor the learning algorithm that produces the policy function.” For this paper, we follow the latter consideration when referring to ”algorithmic collusion”. As such, ”collusive” outcomes are identified by supra-competitive prices without focusing on the internal workings of the algorithms.
Algorithmic collusion has drawn substantial attention from the research community and from policymakers, as we outlined in the introduction. Most of this literature analyzes specific algorithms such as Q-learning for specific model variations, i.e., Bertrand oligopolies with standard all-or-nothing demand, with linear, or logit demand (bertrand1883book). Calvano.2020 analyze a Bertrand competition with logit demand and constant marginal cost. They find that Q-learning agents under self-play can lead to supra-competitive prices higher than the Nash equilibrium. A related sequential move pricing duopoly environment with linear demand (instead of the simultaneous move Bertrand model in Calvano.2020) was analyzed by klein_autonomous_2021, who also found collusion with Q-learning agents. asker_impact_2024 detect in their experiments on Bertrand competition with standard (all-or-nothing) demand that collusion depends on specifics of the Q-learning algorithm (e.g., synchronous vs. asynchronous updating). lambin2024less analyzes a two-staged exploration scheme that helps explain how collusion can sometimes form with Q-learning under self-play, and schaefer_emergence_2023 empirically derives a probability boundary for the evolution of cooperation in a repeated Prisoner’s dilemma. While there are measures that can be implemented against collusion, e.g., on a platform-level (johnson_platform_2023), these articles agree on the potential risk to consumer welfare posed by algorithmic sellers, although some raise doubts about the purposeful coordination between the algorithms.
In contrast, abada_collusion_2024 analyze Q-learning in Bertrand oligopolies and show that Q-learning algorithms with sufficiently large $\epsilon$-greedy exploration show no collusion. den_boer_artificial_2022 provide a detailed analysis of the inner workings of Q-learning and reveal that there is no immediate reason to believe that Q-learning would lead to collusion easily. In addition, eschenbaum2022robust criticize the claim that algorithms can be trained offline to successfully collude online in different market environments. The authors find that collusion breaks down when collusive reinforcement learning policies are extrapolated from a training environment to the market.
Most of this experimental literature on algorithmic collusion is based on Q-learning, although there is no evidence that this algorithm is particularly important or widespread for algorithmic pricing. hansen_frontiers_2021 analyze the price levels that arise in a duopoly setting where UCB (”Upper Confidence Bound”) agents determine prices. They run a series of experiments where symmetric UCB algorithms interact simultaneously in a Bertrand economy competition with linear demand functions. The agents observe a perturbed estimate of their revenues, which are a result of their prices and the corresponding demand. hansen_frontiers_2021 find that, under sufficiently small reward noise, agents eventually explore prices in a correlated manner, giving rise to supra-competitive outcomes. [^4] Note that in this literature review, we focus on algorithms that were not designed to collude (meylahn_learning_2022). douglas_naive_2024 provide more details on these results by analyzing UCB and an $\epsilon$-greedy bandit algorithm in a repeated Prisoner’s dilemma. They argue that deterministic algorithms (UCB) always converge to collusion while randomized algorithms ($\epsilon$-greedy) never do so in this simple game.
A more extensive discussion of the literature is provided by abada2024algorithmic. They find that in the settings studied in the literature, sets that include Q-learning and non-Q-learning algorithms do not seem to constitute persistent sets of supracompetitive algorithms. We provide a rationale for this by analyzing the properties of these games and analyze a variety of bandit algorithms that have been reported for algorithmic pricing.
In summary, the literature on algorithmic collusion largely focuses on specific algorithms (mostly Q-learning) that are employed in specific types of oligopoly models. The same type of algorithm is used in a symmetric model. We analyze a variety of multi-armed bandit algorithms competing against the same, but also against different algorithms. We analyze standard all-or-nothing demand, linear, and logit demand models in a Bertrand competition.
2.3 Learning in Games
The literature on algorithmic collusion is inherently tied to that on learning in games. The latter has a long history and asks what type of equilibrium behavior (if any) may arise in the long run of a process of learning and adaptation, in which agents are trying to maximize their payoff while learning about the actions of other agents through repeated interactions (fudenberg1999TheoryLearningGames). Cournot’s study of duopoly (cournot1838recherches) already introduced a Nash equilibrium and a particular learning process. Cournot’s model of best reply dynamics appears unrealistic as a model for algorithmic pricing, because players would need a lot of information about competitors. Later, brown1951IterativeSolutionGames suggested fictitious play, which converges to equilibrium distributions for any two-player, finite-strategy, zero-sum game. However, shapley1964TopicsTwopersonGames established that it can lead to cycles and that the frequency distribution does not necessarily converge. Many learning algorithms have been developed, ranging from iterative best-response to first-order online optimization algorithms in which agents follow their utility gradient in each step (mertikopoulos2019learning, bichler2023soda).
In this general context, it is well known that the dynamics of learning agents do not always converge to a Nash equilibrium (milionis2022nash, vlatakis2020no): they may cycle, diverge, or be chaotic, even in zero-sum games, where the Nash equilibrium is tractable (mertikopoulos2018cycles, bailey2018multiplicative). While there is no comprehensive characterization of games that are “learnable” and one cannot expect that uncoupled dynamics lead to Nash equilibrium in all games (hart2003uncoupled, milionis2022nash), there are some important results regarding learners. A classical result is that the class of no-regret learning algorithms converges to the so-called [coarse] correlated equilibrium ([C]CE) of a game fudenberg1999TheoryLearningGames, a result that has recently been extended to Markov games for an algorithm called ”V-learning” (jin_v-learningsimple_2024). [C]CEs are supersets of Nash equilibria. However, CCEs can also contain dominated strategies and are a rather weak solution concept. Convergence guarantees to the stronger class of CEs exist for no-internal-regret algorithms, like the scheme introduced by foster_regret_1999. These schemes, until this point, lack the simplicity of most no-external-regret algorithms. Our results provide a complementary guarantee for mean-based algorithms.
Less is known about conditions of games in which learning algorithms converge to a Nash equilibrium. In a landmark paper, monderer1996potential introduced the class of potential games and showed that Cournot oligopolies with linear price or cost functions are potential games. Potential games are guaranteed to have at least one pure Nash equilibrium. Importantly, it was shown that algorithms such as best and better response dynamics or fictitious play converge to a Nash equilibrium in these games. More recently, convergence of MWU (palaiopanos2017multiplicative) and Exp3 (cohen2017learning) in potential games was also shown.
Another important property of a game is supermodularity (vives1999oligopoly, topkis1998supermodularity, milgrom1990rationalizability). In supermodular games, the best response of each player is positively correlated with the strategies chosen by other players. milgrom1990rationalizability showed that Bertrand oligopoly games are (log-)supermodular if each firm’s elasticity of demand is a decreasing function of its competitors’ prices. Later, milgrom1991adaptive established that adaptive learning algorithms converge to the unique NE in a Cournot duopoly model and Bertrand oligopoly models with linear and logit demand. Adaptive learning algorithms include best-response dynamics and fictitious play. In the bandit algorithms that we analyze, an agent can only observe their profit for a particular action, but not necessarily the strategies of all other players, as is necessary for these earlier equilibrium-finding algorithms.
mertikopoulos2024unified develops a general stochastic approximation template that unifies many online learning algorithms (gradient, multiplicative weights, optimistic, bandit, etc.) and analyzes their convergence properties across broad classes of games. Our paper complements this by focusing on repeated Bertrand competition, showing how such bandit dynamics manifest in price competition with direct implications for market design and regulation.
It is important to note the distinction between learning algorithms that have access to full payoff information (as in fudenberg1999TheoryLearningGames) and bandit feedback, where players only observe realized payoffs from their chosen actions; while convergence results are more established in the former case, our work contributes to the latter, where such guarantees are harder to obtain. Our work also differs from the extensive literature on dynamic pricing in the monopolist setting (see, e.g., den2015dynamic), which focuses on single-seller learning of demand, whereas we study learning dynamics in competitive multi-seller environments where strategic interaction is central
Finally, some recent papers aim to find an algorithm that, when employed by all agents, yields a Nash equilibrium. They require a common and coordinated exploration scheme. yang_competitive_2024 introduce an algorithm that learns the Nash equilibrium in a variety of Bertrand settings, based on a common exploration scheme that all participants need to follow. goyal_learning_2023 restrict their attention to the multinomial logit-demand case and show that the Nash equilibrium can be found with a specialized, decentralized online learning algorithm. Similar to our work, wang_learning_2022 focus on rationalizability in (coarse) correlated equilibria. They provide a number of algorithms that are able to achieve this under limited coupling requirements.
3 Model
In the following, we will first introduce the necessary notation and definitions before we define different types of Bertrand competition models.
3.1 Basic Definitions and Notation
We begin with the concepts of a finite normal-form game, a Nash equilibrium, and a symmetric game. A normal-form game is a representation in game theory that defines the strategies available to each player, their corresponding payoffs, and the resulting outcomes in a simultaneous and strategic interaction. Formally, we have the following definition.
Definition 2 ((Finite) Normal-form Game).
A normal-form game with $n$ players can be described as a tuple $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$, with a finite set of players $i\in\mathcal{N}=\{1,\dots,n\}$, an action space $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ for all $i\in\mathcal{N}$, and payoff or utility functions $u=(u_{1},\dots,u_{n})$ with $u_{i}:\mathcal{A}\rightarrow\mathbb{R}$. We call the game a finite normal-form game if the action spaces are also finite, i.e., $|\mathcal{A}_{i}|<\infty$.
Please note the slight abuse of notation in the subscripts: Previously, we used indices to refer to time (e.g., $u_{t}(\dots)$”); now, we mostly differentiate players (e.g., $u_{i}(\dots)$”). In the following, we will use letters $s$, $t$, and $T$ whenever we consider time, and $i$, $j$, and $n$ whenever we refer to players.
In a normal-form game, all agents $i=1,\dots,n$ submit their actions $a_{i}$ simultaneously to form an action profile $a=(a_{1},\dots,a_{n})$. We often abbreviate the profile by $(a_{i},a_{-i})$ where $a_{-i}=(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n})$. Similarly, we will sometimes index parts of combined spaces as follows: $\mathcal{A}_{-i}=\mathcal{A}_{1}\times\dots\times\mathcal{A}_{i-1}\times\mathcal{A}_{i+1}\times\dots\times\mathcal{A}_{n}$. The agents may also play a distribution over actions, known as a mixed strategy, denoted by $x_{i}\in\Delta(\mathcal{A}_{i})$, where $\Delta(\mathcal{A}_{i})$ is the probability simplex over $\mathcal{A}_{i}$. For any joint distribution $x\in\Delta(\mathcal{A})$ over action tuples, the expected utility of player $i$ is $u_{i}(x)=\operatorname{\mathbb{E}}_{a\sim x}\left[u_{i}(a_{i},a_{-i})\right]$. Similarly, we will write $u_{i}(a_{i},x_{-i})=\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a_{i},a_{-i})\right]$ for the expected utility of some action $a_{i}$ of player $i$ against randomizing opponents.
Textbook versions of the Bertrand competition assume symmetry, which allows for clean results. In contrast, our main result and our experiments do not rely on this assumption.
A Nash equilibrium is a situation in a strategic interaction where each player’s strategy is optimal given the strategies chosen by all other players, and no player has an incentive to unilaterally deviate from their chosen strategy.
Definition 3 (Nash Equilibrium (NE)).
In a normal-form game $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$, a strategy profile $x^{*}=(x_{1}^{*},\ldots,x_{n}^{*})$ is a Nash equilibrium if, for every player $i\in\mathcal{N}$, we have:
$u_{i}(x_{i}^{*},x_{-i}^{*})\geq u_{i}(x_{i},x_{-i}^{*}),\quad\forall x_{i}\in\Delta\mathcal{A}_{i}.$ (2)
It is well-known that computing Nash equilibria is PPAD-hard in the worst case (daskalakis2009complexity). There are specific subsets of normal-form games where the utilities have more structure. This allows us to find an equilibrium with faster iterative algorithms. Potential games (monderer1996potential) and supermodular games (topkis1979submodular_games, milgrom1990rationalizability) have received significant attention.
Definition 4 (Potential Game).
A game $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$ is a potential game if there exists a potential function $\phi:\mathcal{A}\rightarrow\mathbb{R}$ such that for every player $i\in\mathcal{N}$ and every pair of actions $a_{i},a^{\prime}_{i}\in\mathcal{A}_{i}$
$u_{i}(a_{i},a_{-i})-u_{i}(a^{\prime}_{i},a_{-i})=\phi(a_{i},a_{-i})-\phi(a^{\prime}_{i},a_{-i}),\quad\forall a_{-i}\in\mathcal{A}_{-i}.$ (3)
It is known that best response dynamics, better response dynamics, Fictitious Play, Replicator Dynamics, and simultaneous gradient ascent all converge in potential games (fudenberg1999TheoryLearningGames, monderer1996potential, sandholm2010population, swenson2018best). cohen2017learning could also show convergence of bandit algorithms such as Exp3 to a Nash equilibrium in potential games.
Bertrand competitions with some demand models are supermodular games, which were introduced by topkis1979submodular_games. Supermodular games are defined on complete lattices and require supermodular utility functions (see, for instance, (milgrom1990rationalizability)). Since we only consider one-dimensional action spaces $\mathcal{A}\subset\mathbb{R}$, the utility functions trivially satisfy the supermodularity condition, and we can use the componentwise ordering to define a complete lattice on $\mathcal{A}_{i}$ and $\mathcal{A}_{-i}$. Then, the definition of supermodular games reduces to the following.
Definition 5 (Supermodular Game).
A normal-form game $\mathcal{G}=\left(\mathcal{N},\mathcal{A},u\right)$ with $\mathcal{A}_{i}\subset\mathbb{R}$ is called a supermodular game if it satisfies the following conditions for each player $i\in\mathcal{N}$:
  1. (1)
    $u_{i}$ is order upper semi-continuous in $a_{i}$ and continuous in $x_{-i}$ and has a finite upper bound.
  2. (2)
    $u_{i}$ has increasing differences in $a_{i}$ and $a_{-i}$, i.e., for any two action profiles $a=(a_{i},a_{-i})$ and $a^{\prime}=(a_{i}^{\prime},a_{-i}^{\prime})$ such that $a_{i}^{\prime}\geq a_{i}$ and $a_{j}^{\prime}\geq a_{j}$for all $j\neq i$, the following inequality holds:
    $u_{i}(a_{i}^{\prime},a_{-i}^{\prime})-u_{i}(a_{i},a_{-i}^{\prime})\geq u_{i}(a_{i}^{\prime},a_{-i})-u_{i}(a_{i},a_{-i}).$ (4)
If we restrict ourselves to finite normal-form games, i.e., discrete actions, then we only have to check the increasing differences property (2) and can ignore the continuity requirements (1).
It is known that pure strategy Nash equilibria exist in supermodular games. The set of actions surviving iterated strict dominance has a greatest and least element, and both are Nash equilibria (levin2006solution). Consequently, if a game has a unique Nash equilibrium, then strict elimination of dominated strategies finds this equilibrium. However, the elimination of dominated strategies is not a learning method that can be used independently by all players of the game because it requires complete information about the payoff matrix of a game. Such information is not available in algorithmic pricing. Therefore, we will focus on uncoupled algorithms that can be used independently by the agents and that only learn from bandit feedback after each round.
3.2 Bertrand Competition
The Bertrand pricing game or Bertrand competition (bertrand1883theorie) is an economic model that describes the revenue of a firm according to the price it sets for its product and the resulting demand. The firm’s action $a_{i}\in\mathcal{A}_{i}$ is the price for a good it wants to produce. All firms produce the same good, and they all take action simultaneously. Firms compete for demand with their prices and affect each other’s revenues: The demand depends on all agents’ actions and is decreasing in the agent’s own price. As is common, we assume that the cost functions are linear in the demand. In summary, the utility can be described by
$u_{i}(a_{i},a_{-i})=d_{i}(a)\cdot(a_{i}-c_{i}).$ (5)
3.2.1 Demand Models
We consider different demand models. The standard or all-or-nothing demand (also called Bertrand demand) is similar to an auction, where only the highest bidder gets the item. In this case, the firm with the lowest price gets all the demand. Different from auctions, where the valuation of an item does not depend on its price, the demand in this competition is additionally linearly decreasing in the agent’s price. If multiple firms offer the lowest price, the demand is shared equally. This leads to a demand function given by
$d_{i}(a_{i},a_{-i})=\begin{cases}\frac{D}{n_{min}}(1-a_{i})&\text{if }i\in\arg\min_{j\in\mathcal{N}}a_{j}\\ 0&\text{else}\end{cases},$ (standard)
where $D>0$ is the maximum total demand, and $n_{min}:=\lvert\arg\min_{j\in\mathcal{N}}a_{j}\rvert$ is the number of firms with the lowest price.
Another demand model that was used in Calvano.2020 is the multinomial or logit demand. In this setting, the demand is split between the $n$ agents/goods and some outside good (indexed with 0) according to
$d_{i}(a_{i},a_{-i})=\dfrac{\exp\left(\frac{\alpha_{i}-a_{i}}{\mu}\right)}{\exp\left(\frac{\alpha_{0}}{\mu}\right)+\sum_{j=1}^{n}\exp\left(\frac{\alpha_{j}-a_{j}}{\mu}\right)}.$ (logit)
The parameters $\alpha_{i}>0$ capture different product quality indices and $\mu>0$ models a product differentiation between the goods, i.e., if $\mu\rightarrow 0$, the goods are perfect substitutes. (Please note the difference between $\alpha$ (”alpha”) and $a$ in the formula.)
And finally, we also introduce the linear demand that was considered in hansen_frontiers_2021. The agents’ demand function is given by
$d_{i}(a_{i},a_{-i})=\alpha_{i}-\beta_{i}a_{i}+\tfrac{\gamma}{n-1}\sum_{j\neq i}a_{j}.$ (linear)
Similar to the standard model, the demand decreases with rate $\beta_{i}>0$ in the price of the agent, but it also increases with rate $\gamma>0$ in the sum of the prices of all other agents. We assume that the effect of the own price is greater than the effect of the average opponents’ prices, i.e., $\beta_{i}>\gamma$.
In a standard Bertrand model with homogeneous products and symmetric firms, prices equal marginal costs in equilibrium. However, this changes with asymmetries or product differentiation. With asymmetric costs between firms, the Nash equilibrium typically involves the low-cost firm pricing at or just below the marginal cost of the high-cost firm. With product differentiation (as in logit demand), equilibrium prices are typically above marginal costs due to firms having some market power. Independent of the demand model, we will use the Nash equilibrium as a baseline for the learning outcome.
3.2.2 Game-Theoretical Properties
Let us now discuss some properties that are already known about Bertrand pricing models. The referenced propositions can be found in the online appendix. milgrom1991adaptive showed that Bertrand competition with linear and logistic demand and continuous actions are (log-)supermodular games. As indicated earlier, iterated elimination of dominated strategies converges to a unique Nash equilibrium in this model. However, this is not the case for standard all-or-nothing demand, as we show (see Online Supplement, Proposition 3). Potential games have even stronger properties, as we have seen. Apart from best- and better-response dynamics, and fictitious play, the class of no-regret learning converges to equilibrium in potential games. We show that the Bertrand competition with a linear demand function is not only supermodular, it is also a potential game (see Online Supplement, Proposition 4). For Bertrand competitions with standard all-or-nothing and logit demand, we can show that these games are, in general, not potential games (see Online Supplement, Proposition 5 and 6). An overview of these concepts for the different utility models is given in Table 1.
3.3 Alternative Solution Concepts
We explore alternative solution concepts to the Nash equilibrium. While the latter is computationally hard to find, there exist simple iterative algorithms that can identify strategy profiles corresponding to these solution concepts. In some games, the alternatives coincide with the Nash equilibrium, which is what we will leverage in our convergence proof for mean-based algorithms.
3.3.1 (Coarse) Correlated Equilibria
Apart from Nash equilibria (NE), Correlated Equilibria (CE) and Coarse Correlated Equilibria (CCE) have received significant attention in the literature on learning in games. It is well known that $NE\subseteq CE\subseteq CCE$.
Definition 6 (Correlated Equilibrium).
A joint distribution $\sigma\in\Delta(\mathcal{A})$ is a correlated equilibrium if for every player $i\in\mathcal{N}$ and for all actions $a_{i},a_{i}^{\prime}\in\mathcal{A}_{i}$,
$\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i},a_{-i})\mid a_{i}]\geq\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i}^{\prime},a_{-i})\mid a_{i}].$
Definition 7 (Coarse Correlated Equilibrium).
A joint distribution $\sigma\in\Delta(\mathcal{A})$ is a coarse correlated equilibrium if for every player $i\in\mathcal{N}$ and for all actions $a_{i}^{\prime}\in\mathcal{A}_{i}$,
$\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i},a_{-i})]\geq\mathbb{E}_{a\sim\sigma}[u_{i}(a_{i}^{\prime},a_{-i})].$
Both CE and CCE capture a scenario in which a mediator recommends actions to the players, based on a joint action profile distribution, such that no player wants to deviate. For CEs, deviation may depend on the recommended action, while the deviation for CCEs does not take this information into account. The corresponding CCE constraints form a polytope, and therefore, individual CCEs can be computed via a linear program. It was shown that algorithms with no internal regret converge to a CE and those with no external regret to a CCE (foster1997CalibratedLearningCorrelated). Popular bandit algorithms such as Exp3 are no-external-regret algorithms. However, CCEs are a weak solution concept, and they can contain dominated strategies (viossat2013no). Also, the set of CCEs can be a very large superset of the set of Nash equilibria. In Example 1, we computed different CCEs for Bertrand competition models with linear and all-or-nothing demand.
Example 1 (Coarse Correlated Equilibria in Bertrand Competitions).
We consider Bertrand competitions with 2 players, action spaces $\mathcal{A}=\{0.1,0.2,\dots,0.9\}$, cost parameters $c_{1}=c_{2}=0$, and two different demand functions, namely linear demand ($\alpha_{i}=0.48,\,\beta_{i}=0.9,\,\gamma=0.6$ for $i\in\mathcal{N}$) and all-or-nothing demand ($D=1$). The computed CCEs are visualized in Figure 1.
(a) Linear Demand
(b) All-or-Nothing Demand
Figure 1: Coarse Correlated Equilibria for Different Bertrand Competitions
In the figures, we visualize exemplary coarse correlated equilibria and highlight the Nash equilibria (boxes with black frames). The CCEs were computed by solving an LP where the constraints come from the definition of a CCE, and the objective was chosen such that the Wasserstein distance to the Nash equilibria is maximized, i.e., we put more weight on action profiles further away from the equilibrium profiles.
For the linear demand model, the CCEs contain only actions near the Nash equilibrium. However, for the all-or-nothing demand model, we can find CCEs that contain dominated actions. This is why the guarantee that no-regret learners converge to a CCE is not sufficient in a Bertrand competition with all-or-nothing demand.
3.3.2 Serially Undominated Set
Another important solution concept that can be computed by means of a simple, iterative algorithm is that of a serially undominated set. For this, let us introduce different types of dominance relationships.
Definition 8 (Dominated Actions).
An action $a_{i}\in\mathcal{A}_{i}$ is
  • strongly dominated, if there exists another action $\hat{a}\in\mathcal{A}_{i}$ such that $u_{i}(\hat{a},a_{-i})>u_{i}(a_{i},a_{-i})$
  • strictly dominated, if there exists a mixed strategy $\hat{x}\in\Delta(\mathcal{A}_{i})$ such that $u_{i}(\hat{x},a_{-i})>u_{i}(a_{i},a_{-i})$
for all opponents’ action profiles $a_{-i}\in\mathcal{A}_{-i}$.
If we iteratively remove dominated actions, we end up with a serially undominated set. From (milgrom1990rationalizability), we know that, in supermodular games, the iterative procedure of removing strongly dominated strategies leads to such a serially undominated set where the maximum and minimum elements are component strategies of pure Nash equilibria. This set includes all Nash equilibria, correlated equilibria, and rationalizable actions (bernheim_rationalizable_1984, pearce_rationalizable_1984).
Iteratively crossing out strictly dominated actions, also known as Iterated Strict Dominance (ISD), leads to a smaller set of actions, which we will call strictly serially undominated set (SSU). A game is said to be strict dominance solvable if ISD yields a unique strategy profile. In finite games, this strategy profile is necessarily the unique Nash equilibrium (fudenberg1991game, Kolpin2009Strict). Note that in order to perform ISD, we need complete information about the payoff matrix. This assumption is too strong for algorithmic pricing, which is why we will focus on bandit algorithms later. In finite games, there is a close relationship between actions that survive ISD and correlated rationalizable actions, which we will explore in the following subsection.
3.3.3 Correlated Rationalizable Set
Our main convergence statement is centered on the concept of correlated rationalizability (brandenburger1987rationalizability). These are actions that are a best response against any joint mixed strategy of the opponents that is also correlated rationalizable.
Definition 9 (Correlated Rationalizable Set (CR)).
The product set of actions $\bar{\mathcal{A}}=\bar{\mathcal{A}}_{1}\times\dots\times\bar{\mathcal{A}}_{n}$, $\bar{\mathcal{A}}_{i}\subseteq\mathcal{A}_{i}$, is correlated rationalizable if for each $i\in\mathcal{N}$ and $a_{i}\in\bar{\mathcal{A}}_{i}$
$\exists~\hat{x}_{-i}\in\Delta(\bar{\mathcal{A}}_{-i}),~\forall~a_{i}^{\prime}\in\mathcal{A}_{i}:~u_{i}(a_{i},\hat{x}_{-i})\geq u_{i}(a_{i}^{\prime},\hat{x}_{-i}).$
This means that an action $a_{i}\in\mathcal{A}_{i}$ is correlated rationalizable if it maximizes the player’s utility for some probability distribution $\hat{x}_{-i}$ with support on $\bar{\mathcal{A}}_{-i}$. Note the recursive nature of the definition, which ensures that only those actions are in the support of $\hat{x}_{-i}$ that are correlated rationalizable themselves.
Different from the definition of rationalizable actions (bernheim_rationalizable_1984, pearce_rationalizable_1984), correlated rationalizability allows the mixed strategies $\hat{x}_{-i}\in\hat{\mathcal{A}}_{-i}$ to be dependent/correlated probability distributions. The set of rationalizable actions is a subset of the set of correlated rationalizable actions, but the subset relation may not be strict. bernheim_rationalizable_1984 has shown that all Nash equilibria are rationalizable. As a consequence, all Nash equilibria are also correlated rationalizable. In finite games, iterated strict dominance and correlated rationalizability give the same solution set (levin2006solution). While this set (SSU) can be computed with ISD having the complete payoff matrix available, we will show that uncoupled mean-based algorithms also find specific instances of the correlated rationalizable set.
3.3.4 Summary
We summarize the relation between different solution concepts for games in Figure 2. It is well known that $NE\subseteq CE\subseteq CCE$ in finite games. The relationship between correlated rationalizable sets (or strictly serially undominated sets) and [coarse] correlated equilibria is more complex. First, we can show that, similar to rationalizable actions, actions contained in the support of correlated equilibria are also correlated rationalizable.
Proposition 1.
Given a finite normal-formal game, any action profile that is in the support of a correlated equilibrium is also in the correlated rationalizable set.
While it is known that CCE may contain dominated strategies (viossat2013no), we can show that actions supported by a CCE are not a superset of correlated rationalizable actions. To show this, we use the equivalence of CR and SSU.
Proposition 2.
There exists a finite-normal form game where an action from the strictly serially undominated set is not in the support of any CCE.
Please refer to the online appendix for both proofs.
NECECR=SSUCCE
Figure 2: Relation Between Solution Concepts
The displayed relations are between the support sets of the solution concepts.
CCE = Coarse Correlated Equilibrium, SSU = Strictly Serially Undominated, CR = Correlated rationalizable, CE = Correlated Equilibrium, NE = Nash equilibrium.
As indicated earlier, algorithms that have no (external) regret (see Equation (1)) converge to a CCE, while algorithms with no internal regret (such as regret matching) converge to CEs (foster1997CalibratedLearningCorrelated). ISD computes the strictly serially undominated set (SSU) and, thus, also the correlated rationalizable set (CR), but this algorithm requires full access to the payoff matrix.
Table 1 summarizes properties of games that are satisfied by Bertrand competition with different demand models (all-or-nothing, linear, and logit demand) and algorithms that are known to converge to a Nash equilibrium if certain of the game properties hold.
Game properties Demand Model Convergence to NE
All-or-Nothing Linear Logit
Potential ✗  (Prop. 5) ✓  (Prop. 4) ✗  (Prop. 6) FP, BR, Exp3
Supermodular & unique NE ✗  (Prop. 3) ✓  (MR90) ✓  (MR90) ISD, FP, BR (MR90)
[…]
Table 1: Overview of Games and Properties.
Fictitious play (FP), best response (BR), Exponential Weights (Exp3), iterated strict dominance (ISD), mean-based (MB); (MR90) refers to milgrom1990rationalizability.
1 The set of CR is almost unique. We show that only two neighboring actions per agent can be in CR. Depending on the discretization, either one or both are NE. 2 Numerically, we observe two neighboring or one unique NE in the game, depending on the discretization. And only the actions of the NE are contained in CR.
The Bertrand competition with linear demand is a potential game for which we know that even bandit algorithms such as Exp3 converge. However, other demand types do not lead to a potential game as we showed earlier. milgrom1990rationalizability showed that Bertrand competition with continuous actions and linear demand is supermodular and has a unique Nash equilibrium, while with logit demand, it is log-supermodular. We don’t know of convergence proofs for bandit algorithms, but we know that ISD, Fictitious Play, and best response dynamics find the unique Nash equilibrium in such supermodular games. ISD also finds correlated rationalizable sets, because they are equivalent to serially undominated sets. Our main analytical contribution is to show that mean-based algorithms, with Exp3 as the leading example, converge to the correlated rationalizable set. Importantly, in a Bertrand competition with all-or-nothing or linear demand, this set is either a singleton, meaning it must be the Nash equilibrium, or consists of only two adjacent actions.
4 Convergence of Mean-Based Algorithms
Let us now focus on mean-based algorithms. These are online optimization algorithms that pick actions with low average rewards with low probability (braverman_selling_2018, deng2022nash).
Definition 10 (Mean-based Algorithm).
Let $\alpha_{t}(a)$ be the average reward of action $a\in\mathcal{A}$ in the first t1t-1 rounds: $\alpha_{t}(a)=\frac{1}{t-1}\cdot\sum_{s=1}^{t-1}u_{s}(a)$. An algorithm is $\gamma_{t}$-mean-based if, for any $a\in\mathcal{A}$, whenever there exists $a^{\prime}\in\mathcal{A}$ such that $\alpha_{t}(a^{\prime})-\alpha_{t}(a)>\gamma_{t}$, the probability that the algorithm picks $a$ at round $t$ is at most $\gamma_{t}$. An algorithm is mean-based if it is $\gamma_{t}$-mean-based for some decreasing sequence $(\gamma_{t})_{t=1}^{\infty}$ such that $\gamma_{t}\rightarrow 0$ as $t\rightarrow\infty$.
There are many no-regret algorithms, such as Exp3, Multiplicative Weights, and the Follow-the-Perturbed-Leader algorithm, which are also mean-based. We know from kolumbus_auctions_2022 that if a mean-based algorithm converges to a CCE, then the CCE is co-undominated, which avoids outcomes with dominated strategies as illustrated in Figure 1(b). deng2022nash and feng_convergence_2021 analyzed the convergence of mean-based algorithms in first-price and second-price auctions, and they found convergence in the complete-information games. We focus on Bertrand pricing competitions, which are different due to the various demand models. In what follows, we provide proof that mean-based algorithms converge to strategy profiles in the correlated rationalizable set.
We analyze convergence in a time-average and in a last-iterate sense (anagnostides2024convergence). While the former measures the fraction of times that all players have chosen actions according to a reference profile or reference set of actions, the latter makes a statement about the current strategy of the players, i.e., their current distribution over action selections.
Definition 11 (Time-average Convergence).
A sequence of action profiles $\{a_{s}\}_{s=1}^{t}$ converges in time-average to a set of actions $\bar{\mathcal{A}}\subseteq\mathcal{A}$ if $\lim_{t\to\infty}\frac{1}{t}\sum_{s=1}^{t}\mathbb{I}[\forall i:a_{s,i}\in\bar{\mathcal{A}}]=1$. The sequence converges in time-average to a single action profile $\bar{a}$ if $\bar{\mathcal{A}}=\{\bar{a}\}$.
With this definition at hand, we can state our central convergence theorem. For brevity of notation, we restrict ourselves to an informal formulation and provide a mathematically thorough version in our online appendix, Section A.4.8. The proof, building on techniques introduced by feng_convergence_2021 and deng2022nash, along with the referenced propositions below, can also be found in Online Appendix A.4.
Theorem 1 (Informal version).
If all players use mean-based algorithms in a finite normal-form game, their empirical frequency of play converges almost surely to the set of correlated rationalizable actions (time-average convergence). This also holds if players $i$ observe $U_{it}$ at time $t$, which is a stochastic version of the average reward $u_{i}$, as long as the random variables $U_{it}(a_{i},a_{-i})$ are bounded, sampled i.i.d. for each pair $(i,t)$, and have expected value $u_{i}$. Moreover, agents may enter the game at different points in time $\tau_{i}$.
So far, we have shown that players will eventually play correlated rationalizable actions in almost every round. But we can also say something about the actual (mixed) strategy profile of the players, i.e., about last-iterate convergence. In the general setting of the previous theorem, we can only show that the support of the mixed strategies is within the correlated rationalizable actions in the limit. In the special case, where the correlated rationalizable actions correspond to the unique Nash equilibrium, e.g., in the all-or-nothing demand setting as described in Proposition 9, we get last-iterate convergence to the Nash equilibrium. We formalize this result in the corollary below and provide its proof in Online Appendix A.4.9.
Corollary 1.
In games with a unique pure Nash equilibrium $a^{*}$ being the only correlated rationalizable action, mean-based algorithms converge almost surely to the Nash equilibrium in last-iterate, i.e., $\lim_{t\to\infty}x_{t}=\bar{x}$, where $x_{t}$ is the sequence of mixed strategy profiles and $\bar{x}_{i}=\mathbf{1}_{a_{i}^{*}}$.
Our analysis leads to a corollary that is of independent interest and beyond the analysis of Bertrand competition.
Corollary 2.
Mean-based algorithms converge to strategy profiles in the serially undominated set.
This follows from the fact that serially undominated sets are equivalent to correlated rationalizable sets of strategy profiles. Next, we focus on the class of Bertrand competition games. Let us present our second main result informally.
Informal theorem (Informal theorem).
In symmetric Bertrand competition games with standard or linear demand and discrete actions, mean-based algorithms converge almost surely to actions that are close to their Nash equilibria.
For this, we just need to show that the correlated rationalizable sets are close to the set of Nash equilibria for the two demand models, which we do in Online Supplement A.3.
5 Experimental Analysis
In our experimental analysis, we analyze the behavior of various bandit algorithms in Bertrand oligopolies with different demand models. The focus variables of our experiments are the converged prices and the convergence behavior. We focus on scenarios that our theoretical contribution does not cover. We use algorithms that are well-known but (with the exception of Exp3) not known to be mean-based and simulate their behavior in a variety of Bertrand oligopolies.
In our numerical experiments, most algorithms consistently reach Nash equilibrium prices within a few iterations, indicating that even without tens of thousands of repeated interactions and without the mean-based property, algorithmic collusion is rarely established. By including different demand functions and symmetric as well as asymmetric environments, we show that our findings are robust across a wide range of treatments.
5.1 Selected Algorithms
In our study, we include $\epsilon$-greedy, UCB, Thompson Sampling, and Exp3, which are widely used and cited in the literature (bubeck2011introduction). All bandit algorithms share similarities. First, an action is selected according to some potentially random action-selection rule that is based on past experiences. Then, the agent receives a corresponding reward feedback with which it subsequently updates its internal state. We describe this generic scheme in the Algorithm 1 below. All algorithms that we analyze follow this generic scheme while implementing distinct rules for selecting actions and updating their beliefs. Details of the algorithms can be found in online supplement B.1. Typically, the algorithms allow for different variations, and each description describes a family of algorithms.
Algorithm 1 Bandit Algorithm Scheme
Initialization: Algorithm parameters $\theta$
for $t=1,2,\dots$ do
  $a_{t}\leftarrow$ get_action($\theta$, $t$) // Agents choose action.
  $u_{t}\leftarrow$ environment($a_{t}$) // Agents submit actions and observe own utility.
  $\theta\leftarrow$ update($\theta$, $a_{t}$, $u_{t}$, $t$) // Agents update their strategy.
end for
5.2 Experiment Setup
Algorithmic collusion can lead to supra-competitive prices between the Nash equilibrium prices and the monopoly prices. Prices that are not supra-competitive will be called competitive in our results. We demonstrate that even in treatments where no analytical convergence results are available, convergence to competitive prices can be observed consistently with many bandit algorithms. We characterize supra-competitive prices via the following metrics. Let $(\bar{a}_{i})_{i=1}^{n}$ be the largest component strategies of any pure Nash equilibrium, and let $(a^{M}_{i})_{i=1}^{n}$ be the actions played in the monopoly. We define the price-collusion index of player $i$ and the profit-collusion index of all players by
$CI_{price,i}(a_{i})=\frac{a_{i}-\bar{a}_{i}}{a^{M}_{i}-\bar{a}_{i}},~i\in[n]\qquad\text{and}\qquad CI_{profit}(a)=\frac{\sum_{i=1}^{n}(u_{i}(a)-u_{i}(\bar{a}))}{\sum_{i=1}^{n}(u_{i}(a^{M})-u_{i}(\bar{a}))}.$ (6)
A price-collusion index of zero captures that player $i$ plays $\bar{a}_{i}$, and thus is non-collusive. Similarly, an index value of one means that the agent plays its monopoly price, indicating perfect collusion. Analogously, the profit-collusion index represents the fraction of monopolistic payoff surplus that all agents could achieve in total.
Let us briefly discuss the choice of metrics. In essence, we want to capture a deviation from competitive play (Nash equilibrium) to cooperation. A Nash equilibrium naturally lends itself as the lower bound of our metric’s values, but the choice of the upper deserves discussion. As noted by loots_datadriven_2023, multiple notions of collusive prices are reasonable, and joint profit maximization might not always be the adequate choice. In line with previous research (Calvano.2020), we nonetheless select these prices as our upper reference points, because in our analyzed settings, joint profit maximization is beneficial to all firms. Being a joint maximum, it is Pareto-optimal for sellers, too. As indicated by loots_datadriven_2023, joint profit maximization also exhibits the highest price increase and the worst effect on consumer welfare, on average.
In the following, we introduce the treatment variables of our experiments, such as the oligopoly configurations and the bandit algorithms, as well as the experiment setup. We then state the conclusions drawn from our experiments and present the empirical evidence supporting them.
In our experiments, we vary the games that the agents compete in, the algorithms they use, and the number of competitors they face. An experiment is formed by a unique combination of these factors. Table 2 provides an overview of the treatment and focus variables. Below, we describe the individual treatments in more detail.
Treatment variables Variations
Oligopoly model Demand models (standard, linear, logit), Symmetric/asymmetric parameterization
Algorithms UCB-T, Exp3-$\epsilon$, $\epsilon$-Greedy, TS, (UCB1, Q-Learning, MWU)
Number of competitors 2 - 10
Focus variables Description
Prices Prices after convergence
Convergence speed Time to convergence and convergence behavior
Table 2: Treatment and Focus Variables of the Experiments
Oligopoly Model.
Our oligopolies model the demand and the costs that firms face. They are based on the games described in Section 3.2. We implemented models with standard, linear, and logit demand functions. Depending on the parameter choices, the utility functions are the same for all agents (symmetric game) or not (asymmetric game). Our experiments are based on the following instances of Bertrand competition. The first three configurations are symmetric games with standard demand (O1), linear demand (O2), and logit demand (O3). O2 and O3 are inspired by Calvano.2020 and hansen_frontiers_2021, respectively. Likewise, we define asymmetric counterparts for duopolies with linear demand (O2’) and logit demand (O3’). The asymmetric settings are a step towards more realistic scenarios where firms might face different costs and demand structures. The parameterization of the models is summarized in Table 3. The price ranges are chosen to span the Nash equilibrium prices and the prices that maximize the joint profit, including a small margin. The agents can choose from 21 evenly distributed prices within this range. For example, an agent in O3 can select actions from the set $\{1.00,1.05,1.10,\dots,1.95,2.00\}$. Other discretizations, and in particular asymmetric discretizations for the players, did not substantially change our results in preliminary experiments.
Model Demand Costs $(c_{1},c_{2})$ Other parameters Price range Nash Monopoly
O1 standard $(0.0,0.0)$ - $[0.05,1.00]$ $(0.05,0.05)$ $(0.48,1.00)$
O2 linear $(0.0,0.0)$ $\alpha=0.48$, $\beta=0.9$, $\gamma=0.6$ $[0.00,1.00]$ $(0.40,0.40)$ $(0.80,0.80)$
O3 logit $(1.0,1.0)$ $a_{1}=a_{2}=2.0$, $a_{0}=0.0$, $\mu=0.25$ $[1.00,2.00]$ $(1.50,1.50)$ $(1.90,1.90)$
O2’ linear $(0.0,0.2)$ $\alpha=0.48$, $\beta=0.9$, $\gamma=0.6$ $[0.00,1.00]$ $(0.45,0.50)$ $(0.80,0.90)$
O3’ logit $(0.5,1.0)$ $a_{1}=1.5$, $a_{2}=2.0$, $\mu=0.25$, $a_{0}=0.0$ $[1.00,2.00]$ $(0.95,1.48)$ $(1.40,1.93)$
Table 3: Parameterization of the Bertrand oligopoly models.
The (pure) Nash equilibria and monopoly prices are stated for the discretized game. In case of multiple equilibria, the maximum prices are taken.
Algorithms.
We investigate the behavior of four widespread bandit algorithms. First, we evaluate a variant of the UCB-Tuned (UCB-T) algorithm that resolves ties randomly and does not eliminate actions. Next, we run experiments with the simple $\epsilon$-greedy algorithm ($\epsilon$-Greedy) and an Exp3 (Exp3-$\epsilon$) algorithm with fixed exploration rate $\epsilon$. We note that this version of Exp3 is not mean-based. Known mean-based versions of Exp3 suffer from a very slowly decaying exploration rate, which makes them impractical for experiments and applications. Finally, we implemented an algorithm (TS) that performs Thompson Sampling with Gaussian priors, likelihoods, and posteriors. We note that our experiment results are stable against reasonable modifications of the parameters. In one result on mean-based algorithms, we also analyze Multiplicative Weights Update (MWU) and the mean-based version of Exp3 introduced by braverman_selling_2018. In the online appendix, we also report results on Q-learning to provide a comparison. However, given the multitude of hyperparameters available in this algorithm, we restrict our analysis to a few experiments replicating prior results. We also compare these results to those of the algorithms in the main part of the article.
Number of Competitors.
Some of our empirical results are based on the number of firms that compete in an oligopoly. The symmetric games (O1, O2, O3) can be easily extended to an arbitrary number of players. We investigate numbers between two and ten for an oligopoly based on O2.
Throughout our experiments, we keep certain experimental settings constant. We discretize the action interval for each game into 21 equidistant points, including the boundaries of the intervals described above. We run each experiment ten times with random seeds from 0 to 9. Fixing the seeds allows us to reproduce results despite the random nature of the algorithms. Each run consists of 250,000 iteration steps during which the agents first submit their actions simultaneously and independently to the market. Then, the demands are evaluated, and the agents observe their respective rewards. The agents update their beliefs at the end of each step. We found that almost all of our experiments converged within the provided time frame.
5.3 Results
Let us now report the main results of our experimental analysis.
5.3.1 Convergence Speed with Mean-based Algorithms
Let us first focus on markets with purely mean-based algorithms, for which we have a convergence guarantee. Speed of convergence is unknown, however. We consider the mean-based versions of Multiplicative Weights and Exp3 that have been described by braverman_selling_2018. Figure 3 shows the evolution of the price collusion index over the course of the experiments for both algorithms. We observe that both converge to very low index values. While the MWU algorithm reaches the final outcome within a few iterations, the Exp3 algorithm requires significantly more iterations. This is to be expected as the mean-based algorithm has access to full feedback while Exp3 only receives information about the action that it sampled. In the remainder, we use a different version of Exp3, which is faster than its mean-based counterpart.
(a) Multiplicative Weights
(b) Exp3 (mean-based)
Figure 3: Evolution of Prices During Training for Mean-based Algorithms
The experiments are based on the symmetric environment with linear demand (O2). The figures display the collusion indices based on the charged prices (running medians over 1000 steps) for ten runs of the experiment in thin lines. The thick line shows the average of all runs and thus indicates the overall trend.
5.3.2 Algorithmic Collusion in Duopolies
Next, we analyze duopolies with different algorithms that are not mean-based anymore. We first report the prices at the end of the experiments in two-player games. After 250,000 iterations, the algorithms consistently settle at a price level (see next section), so we can assume convergence. In Online Supplement B.2, we provide an analysis of the convergence speed. Our first result makes a statement about the collusion levels agents eventually reach.
Result 1.
Supra-competitive pricing only evolves with the UCB algorithm under self-play. Other algorithms consistently price close to the Nash equilibrium prices. In particular, combinations of diverse algorithms rarely display collusive behavior. This happens for all demand models (standard, linear, logit) and for symmetric and asymmetric settings.
We visualize these results in Figure 4, which displays the profit- and price-collusion indices found in our experiments. We average the indices over settings, runs, and - if applicable - over agents. In settings with two identical competing algorithms (diagonal entries), we only observe collusive behavior with the UCB-T algorithm. The combined settings of any two different algorithms show collusion indices close to zero, with a combination of UCB-T and $\epsilon$-Greedy resulting in only slightly supra-competitive prices. We observe collusion both in terms of prices and profit, meaning that colluding algorithms indeed benefit from their non-competitive play.
Our experiments also demonstrate that collusion with Bandit algorithms does not develop in a stable manner. While we could observe collusion with UCB-T for all demand functions, and in symmetric as well as asymmetric games, these experiments usually suffer from a high variance between runs, making the result unpredictable in advance. In contrast, non-colluding algorithms performed consistently in all environments and all runs.
(a) Price-collusion indices
(b) Profit-collusion indices
Figure 4: Collusion in Duopolies
We visualize the collusion indices, based on a) the median prices and b) the mean profits in the last 1000 steps, of several combinations of algorithms. The displayed numbers are the numeric values of the collusion indices; the values in brackets display their corresponding standard deviation. Averages and standard deviations were computed over two agents, ten runs, and the five settings O1 - O3’.
While experiments with two players in simplified oligopolies are useful to extract the essence of algorithmic interaction, we also investigated extensions of our games with more players and richer demand functions. We provide our results in Online Supplement B.
6 Conclusions
The discussion of algorithmic collusion is largely based on the experimental results of specific algorithms and versions of the Bertrand competition. In general, learning algorithms do not converge to an equilibrium. However, we prove that important algorithms do so in repeated Bertrand pricing competition with all-or-nothing and linear demand models. The convergence of mean-based algorithms to correlated rationalizable strategies is of independent interest and might be useful for the analysis of different games as well. The result closes a gap in the literature on learning in games, which largely focused on algorithms with no internal or external regret, which converge to correlated or coarse correlated equilibria, respectively. Mean-based algorithms such as Exp3 only require bandit feedback after each round, a realistic assumption for algorithms used in algorithmic pricing. The fact that correlated rationalizable strategy profiles coincide with the set of Nash equilibria in Bertrand competition games is an important insight, and it shows that such games can be learned even with very simple algorithms relevant in practice.
We also show experimentally that when (not mean-based) online optimization algorithms are used in combination, algorithmic collusion is mostly not sustained in experiments. If sellers use UCB or Q-learning symmetrically, they might learn prices higher than the Nash equilibrium. A combination of different algorithms, however, rarely develops collusion. Such insights are important for regulators when they want to identify collusion in online markets. While multi-armed bandit algorithms are an important and widely used class of algorithms used for algorithmic pricing, one cannot guarantee that sellers don’t use other algorithms. In future research, it will be valuable to analyze and classify alternative algorithms and models. The properties that we showed for the Bertrand competition will also be useful for subsequent studies.
Appendix A Proofs
A.1 Proofs of Game-Theoretical Properties
Proposition 3.
The Bertrand competition with all-or-nothing demand is not supermodular.
Proof.
To prove that the Bertrand-competition with standard all-or-nothing demand is not supermodular, we show that the property of increasing differences, i.e., $u_{i}(a_{i}^{\prime},a_{-i}^{\prime})-u_{i}(a_{i},a_{-i}^{\prime})\geq u_{i}(a_{i}^{\prime},a_{-i})-u_{i}(a_{i},a_{-i})$ with $a_{i}^{\prime}\geq a_{i}$ and $a_{j}^{\prime}\geq a_{j}$ for all $j\neq i$, is not satisfied. To that end, we consider two cases in the simple $n=2$ player case:
For $a_{2}<a_{1}<a_{2}^{\prime}<a_{1}^{\prime}$, the inequality is violated:
$u_{1}(a_{1}^{\prime},a_{2}^{\prime})-u_{1}(a_{1},a_{2}^{\prime})$ $=0-u_{1}(a_{1},a_{2}^{\prime})<0$
$u_{1}(a_{1}^{\prime},a_{2})-u_{1}(a_{1},a_{2})$ $=0-0=0$
For $a_{1}<a_{2}<a_{1}^{\prime}<a_{2}^{\prime}$ with $1-(a_{1}-c_{1})>a_{1}^{\prime}$, the inequality is satisfied:
$u_{1}(a_{1}^{\prime},a_{2}^{\prime})-u_{1}(a_{1},a_{2}^{\prime})$ $=(1-a_{1}^{\prime})(a_{1}^{\prime}-c_{1})-(1-a_{1})(a_{1}-c_{1})>0$
$u_{1}(a_{1}^{\prime},a_{2})-u_{1}(a_{1},a_{2})$ $=0-u_{1}(a_{1},a_{2})<0$
The condition $1-(a_{1}-c_{1})>a_{1}^{\prime}$ simply means that $a_{1}$ is closer to $c_{1}$ than $a_{1}^{\prime}$ is to 1, which is necessary to ensure that $u_{1}(a_{1}^{\prime},a_{2}^{\prime})>u_{1}(a_{1},a_{2}^{\prime})$. Since the inequalities is strictly satisfied and violated, we cannot define a different order on the opponent’s action set such that the increasing differences property is satisfied. Therefore, the game cannot be supermodular. Note that this argument also works for continuous actions. ∎
Proposition 4.
The Bertrand competition with linear demand is a potential game.
Proof.
To show that the Bertrand competition with linear demand is a potential game, we prove that a function $\phi:\mathcal{X}\rightarrow\mathbb{R}$ satisfies the definition of a potential. Our choice of $\phi$ is similar to the potential function of the Cournot setting in slade1994 and is given by
$\phi(a)=\sum_{k}\alpha_{k}a_{k}-\sum_{k}\beta_{k}a_{k}(a_{k}-c_{k})+\tfrac{\gamma}{n-1}\sum_{k<l}a_{k}a_{l}.$
It is easy to check that this function satisfies the condition of an exact potential:
$\phi(a_{i},a_{-i})-\phi(a^{\prime}_{i},a_{-i})$ $=\alpha_{i}[a_{i}-a^{\prime}_{i}]-\beta_{i}[a_{i}(x_{i}-c_{i})-a^{\prime}_{i}(a^{\prime}_{i}-c_{i})]$
$\quad+\tfrac{\gamma}{n-1}\sum_{l>i}[a_{i}-a^{\prime}_{i}]a_{l}+\tfrac{\gamma}{n-1}\sum_{k<i}a_{k}(a_{i}-a^{\prime}_{i}).$
Note that in the sum over $k,l$, we have the two cases where either $k=i$ or $l=i$, which leads to the two sums above. We can combine them and get:
$=\alpha_{i}(a_{i}-a^{\prime}_{i})-\beta_{i}[a_{i}(a_{i}-c_{i})-a^{\prime}_{i}(a^{\prime}_{i}-c_{i})]+\tfrac{\gamma}{n-1}\sum_{j\neq i}a_{j}(a_{i}-a^{\prime}_{i})$
$=d_{i}(a_{i},a_{-i})\cdot(a_{i}-c_{i})-d_{i}(a^{\prime}_{i},a_{-i})\cdot(a^{\prime}_{i}-c_{i})=u_{i}(a_{i},a_{-i})-u_{i}(a^{\prime}_{i},a_{-i}).$
This argument works for both continuous and discrete actions. ∎
Proposition 5.
The Bertrand competition with all-or-nothing demand is not a potential game.
Proof.
To show that the Bertrand competition with standard demand is not a potential game, it is sufficient to show that there exists an example, where the closed path property is not satisfied [monderer1996potential, Theorem 2.8]. For simplicity, we focus on the two player version, but the argument can be extended to more agents. Consider the actions $0<a_{1}<a_{2}<a_{1}^{\prime}<a_{2}^{\prime}<1$ and the closed path $\gamma$ of length 4 given by the action profiles $(a_{1},a_{2})\rightarrow(a_{1}^{\prime},a_{2})\rightarrow(a_{1}^{\prime},a_{2}^{\prime})\rightarrow(a_{1},a_{2}^{\prime})\rightarrow(a_{1},a_{2})$. We know that our game is a potential game if and only if $I(\gamma,u)=0$. In this example, $I(\gamma,u)$ is given by
$I(\gamma,u)$ $:=[u_{1}(a_{1}^{\prime},a_{2})-u_{1}(a_{1},a_{2})]+[u_{2}(a_{1}^{\prime},a_{2}^{\prime})-u_{2}(a_{1}^{\prime},a_{2})]$
$\quad+[u_{1}(a_{1},a_{2}^{\prime})-u_{1}(a_{1}^{\prime},a_{2}^{\prime})]+[u_{2}(a_{1},a_{2})-u_{2}(a_{1},a_{2}^{\prime})]$
Using the definition of the standard demand model, we get
$=[0-D(1-a_{1})(a_{1}-c_{1})]+[0-D(1-a_{2})(a_{2}-c_{2})]$
$\quad+[D(1-a_{1})(a_{1}-c_{1})-D(1-a_{1}^{\prime})(a_{1}^{\prime}-c_{1})]+[0-0]$
$=-D(1-a_{2})(a_{2}-c_{2})-D(1-a_{1}^{\prime})(a_{1}^{\prime}-c_{1})$
If we further assume that both agents bid above their marginal costs, i.e., $c_{1}<a_{1}<1$ and $c_{2}<a_{2}$ the last line is strictly negative and thereby $I(\gamma,u)\neq 0$, which implies that the game cannot be a potential game. ∎
Proposition 6.
The Bertrand competition with logit demand is in general not a potential game.
Proof.
Since the utility function in the Bertrand competition with logit demand is twice differentiable, we can use that the game is a potential game if and only if $\tfrac{\partial^{2}u_{i}}{\partial a_{i}\partial a_{j}}=\tfrac{\partial^{2}u_{j}}{\partial a_{i}\partial a_{j}}$ [monderer1996potential, Theorem 4.5]. Again, we focus on a simplified version to show that the game is in general not a potential game. Let us assume that we have $n=2$ agents with costs $c_{1},c_{2}>0$ and $\alpha_{0}=\alpha_{1}=\alpha_{2}=\alpha$, and $\mu_{0}=\mu_{1}=\mu_{2}=\mu$. If we compute the second derivatives we get
$\quad\frac{\partial^{2}u_{1}(a_{1},a_{2})}{\partial a_{1}\partial a_{2}}=\frac{\partial^{2}u_{2}(a_{1},a_{2})}{\partial a_{1}\partial a_{2}}$
$\Leftrightarrow\quad$ $\mu\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+\left(-a_{1}+c_{1}\right)\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+2\left(a_{1}-c_{1}\right)e^{\frac{-2a_{1}-a_{2}+3\alpha}{\mu}}$
$=\mu\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+\left(-a_{2}+c_{2}\right)\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+2\left(a_{2}-c_{2}\right)e^{\frac{-a_{1}-2a_{2}+3\alpha}{\mu}}$
$\Leftrightarrow\quad$ $\left(a_{1}-c_{1}\right)\left(-\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+2e^{\frac{-2a_{1}-a_{2}+3\alpha}{\mu}}\right)$
$=\left(a_{2}-c_{2}\right)\left(-\left(e^{\frac{\alpha}{\mu}}+e^{\frac{-a_{1}+\alpha}{\mu}}+e^{\frac{-a_{2}+\alpha}{\mu}}\right)e^{\frac{-a_{1}-a_{2}+2\alpha}{\mu}}+2e^{\frac{-a_{1}-2a_{2}+3\alpha}{\mu}}\right)$
It is easy to see that if both agents bid for instance $c_{1}$, only the left side is equal to zero while the right side is generally not. And even if agents are symmetric with respect to the costs, i.e., $c_{1}=c_{2}=c$, and assume that agents bid $a_{1}=2c$ and $a_{2}=3c$ we get
$\Rightarrow\quad$ $c\left(e^{\frac{3c}{\mu}}+3e^{\frac{c}{\mu}}-3\right)e^{\frac{3\alpha-8c}{\mu}}=0,\text{ which is obviously not true in general.}$
Therefore the condition from the theorem is not satisfied and the Bertrand compition with logit demand is in general not a potential game. ∎
A.2 Remarks on Correlated Rationalizability
In this section, we define a series of rational response sets, $(\mathcal{A}^{i})_{i=1}^{K}$, together with a game-dependent number, the competition constant $\delta$, that will play an essential role in our proof of convergence.
Definition 12 (Correlated Rational Responses).
[…]
$R_{i}(\hat{\mathcal{A}})=\{a_{i}\in\mathcal{A}_{i}~|~\exists~\hat{x}_{-i}\in\Delta(\hat{\mathcal{A}}_{-i}),~\forall~a_{i}^{\prime}\in\mathcal{A}_{i}:~u_{i}(a_{i},\hat{x}_{-i})\geq u_{i}(a_{i}^{\prime},\hat{x}_{-i})\}.$
[…]
$\forall~\hat{x}_{-i}\in\Delta(\hat{\mathcal{A}}_{-i}),~\exists~a_{i}^{\prime}\in\mathcal{A}_{i}:\quad u_{i}(a_{i},\hat{x}_{-i})<u_{i}(a_{i}^{\prime},\hat{x}_{-i}).$ (7)
[…]
$\forall k\geq 0,~a_{i}\not\in R_{i}(\mathcal{A}^{k}):\quad\forall~\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i}),~\exists~a_{i}^{\prime}\in\mathcal{A}_{i}:~u_{i}(a_{i}^{\prime},\hat{x}_{-i})-u_{i}(a_{i},\hat{x}_{-i})\geq\delta.$
Proposition 7.
The competition constant $\delta>0$ exists for any finite normal-form game in which the set of correlated rationalizable actions is a strict subset of the action spaces: $\bar{\mathcal{A}}\subset\mathcal{A}$.
Proof.
Let $k\geq 0$ be an iteration such that $\exists a_{i}\not\in R_{i}(\mathcal{A}^{k})$. Since $a_{i}$ is not a correlated rational response, we know that
$\forall~\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i}),~\exists~a_{i}^{\prime}\in\mathcal{A}_{i}:~u_{i}(a_{i}^{\prime},\hat{x}_{-i})-u_{i}(a_{i},\hat{x}_{-i})>0.$
This is equivalent to stating that
$\forall~\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i}),\qquad g(\hat{x}_{-i})>0,$
where we define $(\hat{x}_{-i}):=\max_{a_{i}^{\prime}\in\mathcal{A}_{i}}u_{i}(a_{i}^{\prime},\hat{x}_{-i})-u_{i}(a_{i},\hat{x}_{-i}). gg$ is the maximum of a finite number of linear functions (the utility differences), so it is real-valued and continuous. Additionally, the space $\Delta(\mathcal{A}_{-i}^{k})$ is compact, which implies that the minimum of $g$ on this space is attained according to the extreme value theorem. Since we know that $g(\hat{x}_{-i})>0$, the minimum must be strictly positive, yielding
$\forall~\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i}),\qquad g(\hat{x}_{-i})\geq\delta_{k},$
for some $\delta_{k}>0$. This again implies that
$\forall~\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i}),~\exists~a_{i}^{\prime}\in\mathcal{A}_{i}:~u_{i}(a_{i}^{\prime},\hat{x}_{-i})-u_{i}(a_{i},\hat{x}_{-i})\geq\delta_{k}.$
Now take $\delta=\min_{\{k\geq 0\text{ s.t. }\exists a_{i}\not\in R_{i}(\mathcal{A}^{k})\}}\{\delta_{k}\}>0$, a minimum over a finite number of values, to obtain the result. ∎
Remark 1.
The sets of correlated rationalizable actions $\bar{\mathcal{A}}$ are fix points under $R(\cdot)$, which follows directly from Definition 9.
Remark 2.
The operator $R(\cdot)$ may cycle for some games and some inputs, meaning that after some number of iterations greater than 1, the same set is obtained: $(R\circ\dots\circ R)(\hat{\mathcal{A}})=\hat{\mathcal{A}}$. However, this can only happen if there are some correlated rational responses to $\hat{\mathcal{A}}$ that are not within $\hat{\mathcal{A}}$. If we start applying $R$ at the full action sets $\mathcal{A}^{0}$ as defined above, we will never end up in such a cyclic relation. Instead, we will reach a fixed point of $R$ that is a correlated rationalizable set. The following proposition makes this statement more formal.
Proposition 8.
As before, let $\mathcal{A}^{0}=\mathcal{A}=\mathcal{A}_{1}\times\dots\times\mathcal{A}_{n}$ and $\mathcal{A}^{k}=R(\mathcal{A}^{k-1})$, $k=1,\dots,K$, define the sequence of correlated rational responses. The sets $\mathcal{A}^{k}_{i}$ are (non-strict) subsets of their predecessors $\mathcal{A}^{k-1}_{i}$: $\mathcal{A}^{k}_{i}\subseteq\mathcal{A}^{k-1}_{i}$.
Proof.
We prove this statement by induction.
Induction start, $k=1$:
Since $\mathcal{A}^{0}=\mathcal{A}$, we get $\mathcal{A}^{1}\subseteq\mathcal{A}^{0}$.
Induction step, $k\to k+1$:
Assume that $\mathcal{A}^{k^{\prime}}\subseteq\mathcal{A}^{k^{\prime}-1}$ holds for all $k^{\prime}\leq k$. We show that this also holds for $k+1$, i.e., $\mathcal{A}_{i}^{k+1}\subseteq\mathcal{A}_{i}^{k}$. Assume, towards contradiction, that $\mathcal{A}^{k+1}_{i}$ contains an element $a\not\in\mathcal{A}^{k}_{i}$. For this element, there is a mixed opponent strategy $\hat{x}_{-i}\in\Delta(\mathcal{A}^{k}_{-i})$ such that $a$ maximizes $u_{i}(a,\hat{x}_{-i})$. Since $\mathcal{A}^{k}_{-i}\subseteq\mathcal{A}^{k-1}_{-i}$, $\hat{x}_{-i}\in\Delta(\mathcal{A}^{k-1}_{-i})$, and thus $a$ must also be in $\mathcal{A}^{k}_{i}$, leading to contradiction.
Finally, we also provide proof for Proposition 1 and 2.
Proof of Proposition 1.
Consider an action $a_{i}\in\mathcal{A}_{i}$ that is supported in some CE, meaning that there is some probability distribution $p$ on the cartesian product of action sets $\mathcal{A}=\mathcal{A}_{1}\times\dots\times\mathcal{A}_{n}$ such that for all $a_{i}^{\prime}\in\mathcal{A}_{i}$
$\operatorname{\mathbb{E}}_{a\sim p}[u_{i}(a)|a_{i}]\geq\operatorname{\mathbb{E}}_{a\sim p}[u_{i}(a_{i}^{\prime},a_{-i})|a_{i}].$
The existence of $p$ implies that there exists a joint distribution $\hat{x}_{-i}$ on the opponents’ actions $\mathcal{A}_{-i}$, namely the posterior distribution of $p$ given $a_{i}$, such that for all $a_{i}^{\prime}$
$u_{i}(a_{i},\hat{x}_{-i})\geq u_{i}(a_{i}^{\prime},\hat{x}_{-i}).$
This shows that $a_{i}$ is a correlated rational response to $\hat{x}_{-i}$. Because this holds for all players in a CE, all players can only choose correlated rational responses. Mutually correlated rational responses show that the actions are indeed correlated rationalizable. ∎
Proof of Proposition 2.
We show that not every strictly serially undominated action is supported by some CCE. To that end we consider the following matrix game with two agents $\mathcal{N}=\{1,2\}$ and two actions $\mathcal{A}_{1}=\mathcal{A}_{2}=\{a_{1},a_{2}\}$ each. The utilities are given by the payoff matrices
$P_{1}=\begin{pmatrix}0&2\\ 2&1\end{pmatrix},\quad P_{2}=\begin{pmatrix}2&0\\ 0&0\end{pmatrix}$,
where $u_{i}(a_{k},a_{l})=(P_{i})_{kl}$ with $i\in\mathcal{N}$ and $k,l\in\{1,2\}$. It is easy to see that no action is strictly dominated. Therefore the set of serially undominated actions is the whole action set. We will now show that there is no CCE $\sigma=(\sigma_{11},\sigma_{12},\sigma_{21},\sigma_{22})\in\Delta(\mathcal{A})$ where the action $a_{1}$ of agent 1 has a positive probability, i.e., we show that $\sigma_{11}=\sigma_{12}=0$, even though $a_{1}$ is undominated. The constraints for the CCE are given by the following constraints (see Definition 7):
$\sigma_{21}(2-0)+\sigma_{22}(1-2)$ $\geq 0,$
$\sigma_{11}(0-2)+\sigma_{12}(2-1)$ $\geq 0,$
$\sigma_{12}(0-2)+\sigma_{22}(0-0)$ $\geq 0,$
$\sigma_{11}(2-0)+\sigma_{21}(0-0)$ $\geq 0.$
Since $\sigma$ has to be a probability measure over $\mathcal{A}$, we also have $\sigma_{11}+\sigma_{12}+\sigma_{21}+\sigma_{22}=1$ and $\sigma_{11},\sigma_{12},\sigma_{21},\sigma_{22}\geq 0$ From the third constraint we immediately get $\sigma_{12}\leq 0$ and thereby $\sigma_{12}=0$. Using this in the second constraint gives us $\sigma_{12}\geq 2\sigma_{11}$. Therefore we have $\sigma_{11}=\sigma_{12}=0$, which means that action 1 of agent 1 is not supported in any CCE even though it is a undominated action $\Rightarrow$ $SSU\not\subseteq CCE$. ∎
A.3 Correlated Rationalizable Set in Bertrand Oligopolies
Proposition 9.
In the symmetric Bertrand competition game with standard demand and action space $\mathcal{A}_{i}=\{a^{1},\dots,a^{m}\}$, $i=1,\dots,n$, with $c<a^{1}<\dots<a^{m}\leq 1$ and $a^{k-1}-(a^{k}-a^{k-1})\geq c$, the set of correlated rationalizable actions is $\bar{\mathcal{A}}_{i}=\{a^{1}\}$ for every player, which is equivalent to the unique Nash equilibrium of the game.
Proof.
First, we show that $a_{i}^{*}=a^{1},\,\forall i\in\mathcal{N}$ is the unique pure Nash equilibrium. It is easy to see that it is a NE, since it is only possible to deviate to a higher action which yields a utility of 0 instead of $\tfrac{D}{n}(1-a^{1})(c-a^{1})>0$. It remains to show that there is no other Nash equilibrium. Other candidates have to be symmetric, i.e., $a_{i}^{*}=a^{k},\,\forall i\in\mathcal{N}$ for some $k\in\{2,\dots,m\}$, since otherwise there is at least one agent pricing higher than some other agents, which means that the player has a utility of zero which could be strictly improved by deviating to $a^{1}$. Now assuming that all agents play $a^{k}$, we want to show that deviating to $a^{k-1}$ strictly increases an agent’s utility, showing that $a_{i}^{*}=a^{k}$ for all $i\in\mathcal{N}$ cannot be a NE.
$u_{i}(a^{k-1},a^{k}_{-i})-u_{i}(a^{k},a^{k}_{-i})$ $=D(1-a^{k-1})(a^{k-1}-c)-\tfrac{D}{n}(1-a^{k})(a^{k}-c)$
$\geq\tfrac{D}{n}\left[(1-a^{k-1})(2a^{k-1}-2c)-(1-a^{k-1})(a^{k}-c)+(a^{k}-a^{k-1})(a^{k}-c)\right]$
$=\tfrac{D}{n}\left[(1-a^{k-1})(a^{k-1}-(a^{k}-a^{k-1})-c)+(a^{k}-a^{k-1})(a^{k}-c)\right]$
$\geq\tfrac{D}{n}(a^{k}-a^{k-1})(a^{k}-c)>0.$
In the last step, we use the assumption on the discretization, i.e., $(a^{k-1}-(a^{k}-a^{k-1})-c)\geq 0$. Therefore, we have uniqueness of our NE.
Next, we want to show that no other action than $a^{1}$ can be correlated rationalizable. Towards contradiction, assume we have supports $\hat{\mathcal{A}}_{1},\dots,\hat{\mathcal{A}}_{n}$ for all players where at least one player can play an action unequal $a^{1}$. We show that there must be some player $i$ and one action for which Equation (7) holds, meaning that the support will change under $R_{i}$. W.l.o.g. let player $i$ support actions up to $a^{k},k\in\{2,\dots,m\}$, the highest action of any support. (There could be multiple such players.) The other players have a joint distribution $x_{-i}$ with support only on actions up to $a^{max}=a^{k}$. Let us define $\Delta:=\max_{a}u_{i}(a)$ and $p\in[0,1]$, which denotes the probability that all opponents play $a^{k}$ under $x_{-i}$. Depending on $x_{-i}$ (in particular $p$), we will show that either $a^{1}$ or $a^{k-1}$ can be chosen to get inequality (7).
  1. (i)
    If $p\leq\frac{1}{2\Delta}(1-a^{1})(a^{1}-c)$, we choose $a_{i}^{\prime}=a^{1}$:
    $\operatorname{\mathbb{E}}_{-i}\left[u_{i}(a^{1},a_{-i})-u_{i}(a^{k},a_{-i})\right]$ $=p\left[u_{i}(a^{1},a^{k}_{-i})-u_{i}(a^{k},a^{k}_{-i})\right]+(1-p)\left[u_{i}(a^{1},a^{1}_{-i})+\dots\right]$
    $\geq p\tfrac{D}{n}\left[(n-1)(1-a^{1})(a^{1}-c)-(1-a^{k})(a^{k}-c)\right]+\tfrac{D}{n}(1-a^{1})(a^{1}-c)$
    $>-p\tfrac{D}{n}\Delta+\tfrac{D}{n}(1-a^{1})(a^{1}-c)$
    $>\tfrac{D}{2n}(1-a^{1})(a^{1}-c)=\frac{1}{2}u_{i}(a^{1},a^{1}_{-i})>0.$
    For the last line we used that $(n-1)(1-a^{1})(a^{1}-c)>0$, $(1-a^{k})(a^{k}-c)\geq-\Delta$, and $p<\frac{1}{2\Delta}(1-a^{1})(a^{1}-c)$.
  2. (ii)
    If $p>\frac{1}{2\Delta}(1-a^{1})(a^{1}-c)$, we choose $a_{i}^{\prime}=a^{k-1}$:
    $\operatorname{\mathbb{E}}_{-i}\left[u_{i}(a^{k-1},a_{-i})-u_{i}(a^{k},a_{-i})\right]$ $\geq p\left[u_{i}(a^{k-1},a^{k}_{-i})-u_{i}(a^{k},a^{k}_{-i})\right]\geq p\tfrac{D}{n}(a^{k}-a^{k-1})(a^{k}-c)$
    $>\tfrac{D}{2n\Delta}(1-a^{1})(a^{1}-c)(a^{k}-a^{k-1})(a^{k}-c)$
    $\geq\frac{1}{2}u_{i}(a^{1}_{i},a^{1}_{-i})\cdot\frac{1}{\Delta}\min_{k}(a^{k}-a^{k-1})(a^{k}-c)$
    $>\frac{1}{2}u_{i}(a^{1},a^{1}_{-i})\cdot\frac{1}{\Delta}\min_{k}(a^{k}-a^{k-1})^{2}>0.$
    In the first line, we ignored instances where all opponents might have played $a^{k-1}$ and used the inequality we derived above to show uniqueness of the NE. In the last line we used $a^{k}-c>a^{k-1}-c\geq a^{k}-a^{k-1}$.
Thus, for all mixed strategies of the opponents within the given support, there always exists an action $a_{i}^{\prime}$ for which the utility is strictly increased. As a consequence, $a^{k}$ is not stable under $R_{i}$ and the supports cannot be correlated rationalizable. ∎
Proposition 10.
For a symmetric Bertrand competition game with linear demand, $\gamma<\beta$, and action space $\mathcal{A}=\{a^{1},\dots,a^{m}\}$ where $0\leq a^{1}\leq\dots,<a^{k^{*}}\leq p^{*}<a^{k^{*}+1}\dots<a^{m}$ with $p^{*}=\frac{\alpha+\beta c}{2\beta-\gamma}$, all correlated rationalizable sets are subsets of $\{a^{k^{*}},a^{k^{*}+1}\}$ for every player $i$.
Proof.
We will use the Nash equilibrium of the continuous game $p^{*}=\tfrac{\alpha+\beta c}{2\beta-\gamma}$ as a reference point and show that only the neighboring points of $p^{*}$, i.e., $a^{k^{*}}$ or $a^{k^{*}+1}$ can be stable under $R$. Using that the utility $u_{i}(a_{i},a_{-i})$ is linear in $a_{-i}$, we can write the difference in the utility between some action $a$ and $a^{\prime}$ for agent $i$ as
$\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a^{\prime},a_{-i})-u_{i}(a,a_{-i})\right]$ $=u_{i}(a^{\prime},\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[a_{-i}\right])-u_{i}(a,\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[a_{-i}\right])$
$=(\alpha-\beta a^{\prime}+\gamma\bar{a}_{-i})(a^{\prime}-c)-(\alpha-\beta a+\gamma\bar{a}_{-i})(a-c)$
$=(a^{\prime}-a)(\alpha+\beta c-\beta(a^{\prime}+a)+\gamma\bar{a}_{-i}),$
where $\bar{a}_{-i}:=\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}[\tfrac{1}{n-1}\sum_{j\neq i}a_{j}]$ is the expected average action of the opponents. Using the definition of $p^{*}$, we obtain
$u_{i}(a^{\prime},x_{-i})-u_{i}(a,x_{-i})=(a^{\prime}-a)\left(\beta(p^{*}-a^{\prime})+\beta(p^{*}-a)-\gamma(p^{*}-\bar{a}_{-i})\right).$ (8)
We show that deviating from an action $a\notin\{a^{k^{*}},a^{k^{*}+1}\}$ to one of these actions is always better if $a$ is the lowest or highest action of any support. Therefore, an action outside of $\{a^{k^{*}},a^{k^{*}+1}\}$ can never be correlated rationalizable. We have to consider two different cases: Given an action profile $a$
  1. (i)
    let agent $i$ be the one with the highest support up to action $a$ with $a>a^{k^{*}+1}>p^{*}$:
[…]
    $\geq(a^{k^{*}+1}-a)(\beta(p^{*}-a^{k^{*}+1})+(\beta-\gamma)(p^{*}-a))$
    $\geq(a^{k^{*}+1}-a^{k^{*}+2})\beta(p^{*}-a^{k^{*}+1})>0.$ (9)
  2. (ii)
    let agent $i$ be the one with the lowest support up to action $a$ with $a<a^{k^{*}}\leq p^{*}$:
[…]
    $\geq(a^{k^{*}}-a)(\beta(p^{*}-a^{k^{*}})+(\beta-\gamma)(p^{*}-a))$
    $\geq(a^{k^{*}}-a^{k^{*}-1})(\beta-\gamma)(p^{*}-a^{k^{*}-1})>0.$ (10)
In summary, this shows that for any choice of, potentially mixed, opponent strategy with support of at most/at least $a$, there is an action that is better than $a$. Thus, $a$ is not stable under $R_{i}$ and cannot be part of any correlated rationalizable set.
A.4 Convergence with Mean-based Algorithms
In this section, we finally provide the proof for our main statement: the convergence of mean-based algorithms to correlated rationalizable actions. Before we start with the formal derivation, we briefly discuss the steps of our proof.
A.4.1 Proof Sketch
We want to show that all agents play only correlated rationalizable actions with probability close to 1 as time goes to infinity. Towards this result, we do the following:
Step 1.
(Proposition 11): We show that every mean-based agent tends to choose correlated rational responses to their competitors’ past actions with high probability. In the context of Bertrand competition games, this means that an individual agent undercuts the opponents’ prices if the prices are too high and increases prices if the prices are too low.
Step 2.
(Proposition 12): By an inductive argument, we find that agents will settle on correlated rationalizable actions. Assume that all players have been playing a subset of the action space for a prolonged time. Every mean-based player will thus select actions that are correlated rational responses to their competitors’ past actions. This will continue for an extended number of steps, which changes the set of past actions. As a result, the players adapt their strategy again by selecting correlated rational responses to the new history. By noting that such a behavior of repeated correlated rational responses will lead to a set of correlated rationalizable responses, we can establish our result. Since mean-based algorithms do not ”exploit” all of the time but sometimes choose ”exploratory” actions at random, we must rely on probabilistic arguments to show that this happens with probability one.
Step 3.
(Propositions 14 and 15): Finally, we extend these results by providing variants of Proposition 11 that extend the results to noisy or stochastic payoffs and staggered entry.
A.4.2 Reminder on Notation
Recall that we defined the series of correlated rational response sets $\mathcal{A}^{0},\dots,\mathcal{A}^{k},\dots$ as $\mathcal{A}^{k}=R(\mathcal{A}^{k-1})$ for all $k\geq 1$ and $\mathcal{A}^{0}=\mathcal{A}$. Let $K$ be the number for which $\mathcal{A}^{K-1}_{i}$ is a strict subset of $\mathcal{A}^{K}_{i}$ for some $i$ and $\mathcal{A}^{K}=\mathcal{A}^{K+1}=\bar{\mathcal{A}}$. Because of remarks 8 and 1 this $K$ always exists. To prove Theorem 1, we show that mean-based algorithms converge along this series of support sets with high probability, approaching $\mathcal{A}^{K}=\bar{\mathcal{A}}$ in the limit. Also, let $m$ be the maximum cardinality of any action space $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$.
A.4.3 Step 1: Individual Players Play Correlated Rational Responses
In the first step, we reason that individual players with mean-based algorithms will choose correlated rational responses. To do so, we look at the history of past actions and rewards $H_{t-1}$ of all players. If the opponents have mainly selected actions within $\mathcal{A}^{k}_{-i}$, we show that the player is likely to choose actions within $\mathcal{A}^{k+1}_{i}=R_{i}(\mathcal{A}^{k})$.
We denote the empirical frequency of opponents (i-i) selecting actions in $\mathcal{A}^{k}_{-i}$ up to time $t$ by
$P_{i,t}(\mathcal{A}^{k}):=\frac{1}{t}\sum_{s=1}^{t}\mathbb{I}\left[\forall j\neq i:~a_{j,s}\in\mathcal{A}^{k}_{j}\right].$
Now we can state the proposition:
Proposition 11.
Given $k\in\{0,\dots,K-1\}$ and $\gamma_{t}<\frac{\delta}{2}$ where $\delta$ is the competition constant of the game. If $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large for history $H_{t-1}$, then $\mathbb{P}_{i,t}(a|H_{t-1})\leq\gamma_{t}$ for all $a\not\in\mathcal{A}^{k+1}_{i}$.
By $\mathbb{P}_{i,t}(a)$ we denote the probability that player $i$ with $\gamma_{t}$-mean-based algorithm chooses action $a$ at time $t$.
Proof.
We say that the empirical frequency $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large if
$P_{i,t-1}(\mathcal{A}^{k})\geq\tilde{p}:=\frac{1}{2}\left(\frac{\Delta u}{\delta+\Delta u}+1\right)\in(0,1),$ (11)
where $\delta$ is the competition constant of the game and $\Delta u$ is the maximum absolute utility difference between any two actions of player $i$ for any opponent profile $a_{-i}$. Let $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ be an action which increases player $i$’s utility by at least $\delta$ compared to playing action $a\not\in\mathcal{A}^{k+1}_{i}$ if opponents play a mixed strategy with support on $\mathcal{A}^{k}_{-i}$. Consider the counterfactual advantage of action $a^{\prime}$ over $a$:
$\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a)$ $=\frac{1}{t-1}\sum_{s=1}^{t-1}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})$
$=\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\not\in\mathcal{A}_{-i}^{k}\end{subarray}}\underbrace{u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})}_{\geq-\Delta u}+\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\in\mathcal{A}_{-i}^{k}\end{subarray}}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})$
$\geq(1-P_{i,t-1}(\mathcal{A}^{k}_{-i}))\cdot(-\Delta u)+P_{i,t-1}(\mathcal{A}^{k})\cdot\underbrace{\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a^{\prime},a_{-i})-u_{i}(a,a_{-i})\right]}_{>\delta}$
$>(1-P_{i,t-1}(\mathcal{A}^{k}))(-\Delta u)+\delta P_{i,t-1}(\mathcal{A}^{k})=P_{i,t-1}(\mathcal{A}^{k})(\delta+\Delta u)-\Delta u$
$\geq\tilde{p}(\delta+\Delta u)-\Delta u=\frac{1}{2}(2\Delta u+\delta)-\Delta u=\frac{\delta}{2}>\gamma_{t}.$
The third line interprets the sum over all times for which the opponents play $a_{-i}\in\mathcal{A}^{k}_{-i}$ as an expectation over a joint action distribution $x_{-i}\in\Delta(\mathcal{A}^{k}_{-i})$ with support on $\mathcal{A}^{k}_{-i}$. We choose $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ such that it performs better than $a$ by a margin of at least $\delta$. For $a\not\in\mathcal{A}^{k+1}_{i}$, this action exists for any such mixed strategy, according to Equation (7). The last line implies that the probability $\mathbb{P}_{i,t}(a)$ that the mean-based algorithm chooses actions $a\not\in\mathcal{A}^{k+1}_{i}$ is small ($\leq\gamma_{t}$). ∎
A.4.4 Step 2: High-probability Induction
In the second step, we use the insights from Propositions 11 to make a statement about all players. Our inductive argument is based on the following observation. Suppose all agents played actions in $\mathcal{A}^{k}$ at time $t$ with high frequency. Then, each of the competitors will tend to play actions within $\mathcal{A}^{k+1}$ in the next step, thus increasing $P_{i,t-1}(\mathcal{A}^{k})$ further (recall that $\mathcal{A}^{k+1}\subseteq\mathcal{A}^{k}$). At the same time, $P_{i,t-1}(\mathcal{A}^{k+1})$ will also increase. Eventually, $P_{i,t^{\prime}-1}(\mathcal{A}^{k+1})$ will exceed some threshold at time $t^{\prime}$ with high probability. This induces the next step towards $\mathcal{A}^{k+2}$, and so on, until the algorithms reach $\mathcal{A}^{K}$.
Instead of looking at $P_{i,t-1}(\mathcal{A}^{k})$ for individual players, we will now move on to looking at the frequency at which all players (not only the opponents) played within $\mathcal{A}^{k}$. We capture that this frequency is high by a series of events
$A_{k}:=\left[\frac{1}{T_{k}}\sum_{t=1}^{T_{k}}\mathbb{I}\left[\exists i:~a_{i,t}\not\in\mathcal{A}^{k}_{i}\right]\leq\beta\right]=\left[\frac{1}{T_{k}}\sum_{t=1}^{T_{k}}\mathbb{I}\left[\forall i:~a_{i,t}\in\mathcal{A}^{k}_{i}\right]\geq 1-\beta\right].$
These events are defined for some small $\beta>0$ and for a corresponding series of time steps $T_{0}<\dots<T_{k}<\dots<T_{K}$. In words, they capture that only in rare cases (fraction $\beta$ of all timesteps up to $T_{k}$), any agent chooses an action outside $\mathcal{A}^{k}$. Equivalently, this means that in most cases (fraction $1-\beta$), all agents played actions from $\mathcal{A}^{k}$.
We note the meaning of the event $A_{K}$ with respect to time-average convergence. For $\beta\to 0$, this event captures that all agents played within $\mathcal{A}^{K}$, i.e., the set of rationalizable actions, for a fraction of the time close to 1. If $\beta\to 0$ as $T_{K}\to\infty$, this implies time-average convergence. The following proposition makes a statement about the probability that this will happen.
Proposition 12.
Assume that all agents follow a $\gamma_{t}$-mean-based algorithm with respect to a monotonously decreasing sequence $\gamma_{t}$ such that $\gamma_{t}\to 0$ as $t\to\infty$. For every sufficiently small $\beta>0$, there exists a rational constant $c>1$ such that
$\mathbb{P}(A_{K})\geq 1-\sum_{k=1}^{K}\sum_{j=1}^{J}\exp\left(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\right),$
where $T_{k}^{j}=c^{j+Jk}T_{0}$ and $T_{k}^{0}=T_{k}$ with a sufficiently large $T_{0}\in\mathbb{N}$ and $J=\lceil\log_{c}\tfrac{2}{\beta}\rceil$.
Moreover, $\mathbb{P}(A_{K})$ approaches 1 as $T_{0}\to\infty$.
Proof.
For a sufficiently small $\beta$, i.e., $0<\beta<1-\tilde{p}$ (as defined in Eq. 11), we choose the constant $c\in\mathbb{Q}$ such that $1<c<\tfrac{1-\beta}{\tilde{p}}$ and a sufficiently large $T_{0}$ such that $\gamma_{t}<\min\{\tfrac{\delta}{2},\frac{\beta}{4nm}\}$ for all $t\geq T_{0}$, where $\delta$ is the competition constant of the game (see Section A.2) and $\Delta u$ is the maximium absolute utility difference between any two action profiles. Additionally, (for technical reasons) we choose $T_{0}$ such that $T^{j}_{k}=c^{j+Jk}T_{0}\in\mathbb{N}$ for all $j=0,\dots,J$ and $k=0,\dots,K-1$.
Before we can prove the result, we have to introduce some sub-events we use in the following:
$B_{k}^{j}$ $:=\left[\frac{1}{T^{j}_{k}-T^{j-1}_{k}}\sum_{t=T^{j-1}_{k}+1}^{T^{j}_{k}}\mathbb{I}[\exists i:~a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]\leq\frac{\beta}{2}\right],$ (12)
$A_{k}^{j}$ $:=(A_{0},\dots,A_{k},B_{k}^{1},\dots,B_{k}^{j}).$ (13)
Note that $B_{K}^{j}$ is a somehow stricter version of the events $A_{k}$ since $\tfrac{\beta}{2}<\beta$ and $\mathcal{A}^{k+1}\subseteq\mathcal{A}^{k}$. Using these definitions and some intermediate results, which we will prove afterwards, we get:
$\mathbb{P}(A_{K})$ $\geq\mathbb{P}(A_{K},A_{K-1},\dots,A_{1},A_{0})$
$=\underbrace{\mathbb{P}(A_{0})}_{=1}\cdot\prod_{k=1}^{K}\mathbb{P}(A_{k+1}|A_{k},\dots,A_{0})$ (Chain Rule)
$\geq\prod_{k=1}^{K}\mathbb{P}(A_{k}^{J}|A_{k},\dots,A_{0})$ (Lemma 1)
$=\prod_{k=1}^{K}\mathbb{P}(B_{k}^{1},\dots,B_{k}^{J}|A_{k},\dots,A_{0})$ (Equation 13)
$=\prod_{k=1}^{K}\prod_{j=1}^{J}\mathbb{P}(B_{k}^{j}|A_{k},\dots,A_{0},B_{k+1}^{1},\dots,B_{k+1}^{j-1})$ (Chain Rule)
$=\prod_{k=1}^{K}\prod_{j=1}^{J}\mathbb{P}(B_{k}^{j}|A_{k}^{j-1})$ (Equation 13)
$\geq\prod_{k=1}^{K}\bigg[1-\sum_{j=1}^{J}\exp\bigg(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg)\bigg]$ (Weierstrass Product Inequality and Lemma 2)
$\geq 1-\sum_{k=1}^{K}\sum_{j=1}^{J}\exp\bigg(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg)$ (Weierstrass Product Inequality)
Since $(T^{j}_{k}-T^{j-1}_{k})=(c-1)c^{j-1+Jk}T_{0}$, the term in the last line converges to $1$ for $T_{0}\rightarrow\infty$. ∎
In the remaining part of this subsection, we will show the necessary steps used in the proof of Proposition 12.
Lemma 1.
$A_{k}^{J}$ implies $A_{k+1}$, which means $\mathbb{P}(A_{k+1})\geq\mathbb{P}(A_{k}^{J})$
Proof.
If $A_{k}^{J}$ holds we know that $B^{1}_{k},\dots,B^{J}_{k}$ hold, which means that the agents have played actions outside of $\mathcal{A}^{k+1}$ for a fraction of at most $\frac{\beta}{2}$ during $\{T_{k}^{0}+1,\dots,T_{k}^{1}\},\dots,\{T_{k}^{J-1}+1,\dots,T_{k}^{J}\}$. Recall that $T_{k}^{0}=T_{k}$ and $T_{k}^{J}=T_{k+1}$. Using this, we can bound the relevant sum for $A_{k+1}$:
$\frac{1}{T_{k+1}}\sum_{t=1}^{T_{k+1}}\mathbb{I}[\exists i:a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]$ $=\frac{1}{T_{k+1}}\sum_{t=1}^{T_{k}}\mathbb{I}[\exists i:a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]+\frac{1}{T_{k+1}}\sum_{t=T_{k}+1}^{T_{k+1}}\mathbb{I}[\exists i:a_{i,t}\not\in\mathcal{A}^{k+1}_{i}]$
$\leq\frac{T_{k}}{T_{k+1}}+\frac{1}{T_{k+1}}\sum_{j=1}^{J}(T_{k}^{j}-T_{k}^{j-1})\frac{\beta}{2}=\frac{T_{k}}{T_{k+1}}+\frac{T_{k+1}-T_{k}}{T_{k+1}}\frac{\beta}{2}$
$=c^{-J}+(1-c^{-J})\frac{\beta}{2}\leq c^{-\log_{c}\left(\frac{2}{\beta}\right)}+\frac{\beta}{2}=\beta.$
Therefore $A_{k}^{J}$ implies $A_{k+1}$. ∎
Lemma 2.
$\mathbb{P}(B_{k}^{j}|A_{k}^{j-1})\geq 1-\exp\bigg(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg),\qquad\qquad j=1,\dots,J.$
Proof.
Recall that $B_{k}^{j}$ defined in Equation (12) is the event, that agents play only a small fraction outside of $\mathcal{A}^{k+1}$ during the time steps $\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}$ given some $j\in J$. To get a bound on the probability, we first introduce random variables that count if an agent plays outside this set:
$X_{t}:=\mathbb{I}\left[\exists i:~a_{i,t}\not\in\mathcal{A}^{k+1}_{i}\right],\qquad t\in\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}.$
We can show that the expectation of $X_{t}$ conditioned on a history $H_{t-1}$ such that $A_{k}^{j-1}$ holds can be bounded by
$\operatorname{\mathbb{E}}[X_{t}|A_{k}^{j-1},H_{t-1}]$ $=\mathbb{P}(\exists i:~a_{i,t}\not\in\mathcal{A}^{k+1}_{i}|A_{k}^{j-1},H_{t-1})$
$\leq n\cdot\mathbb{P}(a_{i,t}\not\in\mathcal{A}^{k+1}_{i}|A_{k}^{j-1},H_{t-1})$ (Union Bound Agents)
$\leq n\cdot m\cdot\mathbb{P}(a_{i,t}=a\notin\mathcal{A}^{k+1}_{i}|A_{k}^{j-1},H_{t-1})$ (Union Bound Actions)
$\leq n\cdot m\cdot\gamma_{t}\leq n\cdot m\cdot\gamma_{T_{k}^{j-1}}.$ (Proposition 11)
Note that Proposition 11 is applicable because of Lemma 3 (below). Recall that $m$ was the number of elements in the largest action set, so this bound holds for any player. Using this upper bound, we can use the random variables $X_{t}$ to define
$Z_{t}:=\sum_{s=T_{k}^{j-1}+1}^{t}\left(X_{s}-nm\gamma_{T_{k}^{j-1}}\right),\qquad t\in\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}.$ (14)
With $Z_{T_{k}^{j-1}}:=0$ and under the condition of $A_{k}^{j-1}$, the value of $Z_{t}$ decreases with $t$ in expectation as we always subtract an upper bound of the expectation of each summand $X_{s}$, i.e., $\operatorname{\mathbb{E}}[Z_{t}|Z_{t-1},A_{k}^{j-1},H_{t-1}]\leq Z_{t-1}$. This makes $Z_{t}$ a super-martingale and allows us to apply Azuma’s inequality.
First, we choose $T_{0}$ sufficiently large, such that $\gamma_{T_{k}^{j-1}}\leq\beta\tfrac{1}{4nm}$, where $n$ is the number of agents and $m$ is the maximum size of any action space. This gives us $|Z_{t}-Z_{t-1}|=|X_{t}-nm\gamma_{T_{k}^{j-1}}|\leq 1$ since $X_{t}\in\{0,1\}$ and $0\leq nm\gamma_{T_{k}^{j-1}}\leq\beta\tfrac{nm}{4nm}\leq 1$. By Azuma’s inequality we get for all $\varepsilon>0$
$\mathbb{P}(Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}\geq\epsilon|A_{k}^{j-1},H_{t-1})\leq\exp\bigg(\frac{-\epsilon^{2}}{2(T^{j}_{k}-T^{j-1}_{k})}\bigg)$
$\Leftrightarrow\quad$ $\mathbb{P}(Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}<\epsilon|A_{k}^{j-1},H_{t-1})\geq 1-\exp\bigg(\frac{-\epsilon^{2}}{2(T^{j}_{k}-T^{j-1}_{k})}\bigg).$
Now we choose $\varepsilon=\frac{\beta}{4}(T^{j}_{k}-T^{j-1}_{k})$ such that we end up with the event described by $B_{k}^{j}$:
$Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}$ $<\varepsilon$
$\Leftrightarrow\quad\sum_{t=T_{k}^{j-1}+1}^{T_{k}^{j}}\left(\mathbb{I}\left[\exists i:~a_{i,t}\not\in\mathcal{A}^{k+1}_{i}\right]-nm\gamma_{T_{k}^{j-1}}\right)$ $<\frac{\beta}{4}(T^{j}_{k}-T^{j-1}_{k})$
$\Leftrightarrow\quad\frac{1}{T^{j}_{k}-T^{j-1}_{k}}\sum_{t=T_{k}^{j-1}+1}^{T_{k}^{j}}\mathbb{I}\left[\exists i:~a_{i,t}\not\in\mathcal{A}^{k+1}_{i}\right]$ $<nm\gamma_{T_{k}^{j-1}}+\frac{\beta}{4}$
Since $\gamma_{T_{k}^{j-1}}\leq\beta\tfrac{1}{4nm}$, the right hand side is less or equal to $\frac{\beta}{2}$ and satisfies thereby $B_{k}^{j}$. Finally, we have
$\mathbb{P}(B_{k}^{j}|A_{k}^{j-1})\geq\mathbb{P}(Z_{T_{k}^{j}}-Z_{T_{k}^{j-1}}<\tfrac{\beta}{4}(T^{j}_{k}-T^{j-1}_{k})|A_{k}^{j-1},H_{t-1})\geq 1-\exp\bigg(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\bigg).$
To apply Proposition 11 in the proof of Lemma 2, we need to make sure that the assumptions hold. This is ensured by the following lemma.
Lemma 3.
If $A_{k}^{j-1}$ holds, the condition $P_{t-1}^{i}(\mathcal{A}^{k})>\tilde{p}$ holds for all $i\in\mathcal{N}$, $t\in\{T^{j-1}_{k}+1,\dots,T^{j}_{k}\}$, and any $j=1,\dots,J$.
Proof.
Recall that given $\tilde{p}$ as defined in Eq.(11), the constants $\beta$ and $c$ in the previous proofs were chosen such that $0<\beta<1-\tilde{p}$ and $1<c<\tfrac{1-\beta}{\tilde{p}}$.
$P^{i}_{t-1}(\mathcal{A}^{k})$ $=\frac{1}{t-1}\sum_{s=1}^{t-1}\mathbb{I}[\forall j\neq i:~a_{j,s}\in\mathcal{A}^{k}_{j}]$
$=\frac{1}{t-1}\bigg(\sum_{s=1}^{T_{k}}\mathbb{I}[\forall j\neq i:~a_{j,s}\in\mathcal{A}^{k}_{j}]+\sum_{s=T_{k}+1}^{T_{k}^{j-1}}\mathbb{I}[\forall j\neq i:~a_{j,s}\in\mathcal{A}^{k}_{j}]+\sum_{s=T_{k}^{j-1}+1}^{t-1}\mathbb{I}[\forall j\neq i:~a_{j,s}\in\mathcal{A}^{k}_{j}]\bigg)$
$\geq\frac{1}{t-1}\bigg(\underbrace{\sum_{s=1}^{T_{k}}\mathbb{I}[\forall j:~a_{j,s}\in\mathcal{A}^{k}_{j}]}_{\text{bounded using }A_{k}}+\underbrace{\sum_{s=T_{k}+1}^{T_{k}^{j-1}}\mathbb{I}[\forall j:~a_{j,s}\in\mathcal{A}^{k}_{j}]}_{\text{bounded using }(B_{k}^{1},\dots,B_{k}^{j-1})}+\underbrace{\sum_{s=T_{k}^{j-1}+1}^{t-1}\mathbb{I}[\forall j:~a_{j,s}\in\mathcal{A}^{k}_{j}]}_{\geq 0}\bigg)$
$\geq\frac{1}{t-1}\big(T_{k}\cdot(1-\beta)+(T_{k}^{j-1}-T_{k})(1-\tfrac{\beta}{2})\big)$
$>\frac{1}{t-1}T_{k}^{j-1}(1-\beta)\geq\frac{T_{k}^{j-1}}{T_{k}^{j}}(1-\beta)=\tfrac{1-\beta}{c}>\tilde{p}$.
In the last line, we used $1-\tfrac{\beta}{2}>1-\beta$ and $t-1<T_{k}^{j}$. ∎
A.4.5 Extension: Asymmetric mean-based algorithms
Up until this point, we assumed that all algorithms are $\gamma_{t}$-mean-based for the same sequence $\gamma_{t}$. The following corollary extends our previous results to situations where this assumption does not hold.
Corollary 3.
Assume each agent $i$, $i\in\mathcal{N}$, follows a $\gamma_{i,t}$-mean-based algorithm with respect to a monotonically decreasing sequence $\gamma_{i,t}$ (in $t$) such that $\gamma_{i,t}\to 0$ as $t\to\infty$. For every sufficiently small $\beta>0$, there exists a rational constant $c>1$ such that
$\mathbb{P}(A_{K})\geq 1-\sum_{k=1}^{K}\sum_{j=1}^{J}\exp\left(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\right),$
where $T_{k}^{j}=c^{j+Jk}T_{0}$ and $T_{k}^{0}=T_{k}$ with a sufficiently large $T_{0}\in\mathbb{N}$ and $J=\lceil\log_{c}\tfrac{2}{\beta}\rceil$.
Moreover, $\mathbb{P}(A_{K})$ approaches 1 as $T_{0}\to\infty$.
Proof.
Let $\gamma_{t}:=\max_{i\in\mathcal{N}}\gamma_{i,t}$ be the maximum $\gamma_{i,t}$ at every time step. The pointwise maximum of finitely many nonincreasing sequences is again nonincreasing, so $\gamma_{t}$ is monotonically nonincreasing with $\gamma_{t}\to 0$. Also, $\gamma_{i,t}\leq\gamma_{t}$, so each of the algorithms is also $\gamma_{t}$-mean-based according to Proposition 13 below. The result follows from the application of Proposition 12. ∎
Proposition 13.
Let ALG be a $\gamma_{t}$-mean-based algorithm. For any sequence $\gamma_{t}^{\prime}$ such that $\gamma_{t}^{\prime}\geq\gamma_{t}$ for all $t$, ALG is also $\gamma_{t}^{\prime}$-mean-based.
Proof.
We need to show that, whenever $\alpha_{t}(a^{\prime})-\alpha_{t}(a)>\gamma_{t}^{\prime}$, the algorithm selects action $a$ with probability less than or equal to $\gamma_{t}^{\prime}$. Assume $\alpha_{t}(a^{\prime})-\alpha_{t}(a)>\gamma_{t}^{\prime}$. This implies that $\alpha_{t}(a^{\prime})-\alpha_{t}(a)>\gamma_{t}$. Since ALG is $\gamma_{t}$-mean-based, we know that $\mathbb{P}(a_{t}=a)\leq\gamma_{t}$. Thus, $\mathbb{P}(a_{t}=a)\leq\gamma_{t}^{\prime}$. ∎
A.4.6 Extension: Stochastic utility
So far, we assumed that utilities are deterministic given the players’ actions. We now drop this assumption and instead model utilities at time $t$ as random variables
$U_{i,t}(a,a_{-i}):\Omega\to\mathbb{R}$
for some space $\Omega$ of possible outcomes. (The actions are assumed to be fixed here, and we define a $U_{i,t}$ for each combination of actions and time.) The randomness is considered to stem from an external source (nature), and is i.i.d. with respect to time. We assume that the expected outcome equals the original deterministic game, which we use to evaluate the correlated rationalizable actions. We also restrict our attention to bounded random variables:
$\operatorname{\mathbb{E}}[U_{i,t}(a,a_{-i})]=u_{i}(a,a_{-i}),\qquad U_{i,t}(a,a_{-i})\in[\underline{u},\overline{u}],$
where $M:=\overline{u}-\underline{u}$ denotes the maximal range. For each fixed action profile $(a_{i},a_{-i})$, the sequence $\{U_{i,t}(a_{i},a_{-i})\}_{t\geq 1}$ is assumed i.i.d. across $t$, independent of other time steps, with mean $u_{i}(a_{i},a_{-i})$.
This definition of random payoffs allows a variety of different payoff models in the oligopoly settings. For example, we can add a random, bounded, noise to the payoffs or demands. Alternatively, we could define a model where the utility function determines the probability that an item is sold. Many models are possible, as long as the analysis of the game is based on the expected utilities.
In this setting, we provide an alternative version of Proposition 11.
Proposition 14.
Given $k\in\{0,\dots,K-1\}$ and $\gamma_{t}<\frac{\delta}{4}$ where $\delta$ is the competition constant of the game. If $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large for history $H_{t-1}$, then $\mathbb{P}_{i,t}(a|H_{t-1})\leq\tilde{\gamma}_{t}:=\gamma_{t}+\exp\left(-\frac{(t-1)\delta^{2}}{32(b-a)^{2}}\right)$ for all $a\notin\mathcal{A}_{i}^{k+1}$.
Remember that, by $\mathbb{P}_{i,t}(a|H_{t-1})$, we denote the probability that player $i$ with $\gamma_{t}$-mean-based algorithm chooses action $a$ at time $t$. This alternative version can still be used with the proofs in section A.4.4 by replacing $\gamma_{t}$ with $\tilde{\gamma}_{t}$ (a monotonically decreasing sequence) and choosing $T_{0}$ sufficiently large such that $\gamma_{T_{0}}<\frac{\delta}{4}$.
Proof.
We again say that the empirical frequency $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large if
$P_{i,t-1}(\mathcal{A}^{k})\geq\tilde{p}:=\frac{1}{2}\left(\frac{\Delta u}{\delta+\Delta u}+1\right)\in(0,1)$
where $\delta$ is the competition constant of the game and $\Delta u$ is the maximum absolute utility difference between any two actions of player $i$ for any opponent profile $a_{-i}$.
Fix some $a\notin\mathcal{A}^{k}$, and let $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ be an action which increases player $i$’s utility by at least $\delta$ in expectation compared to playing action $a$ if opponents play a mixed strategy with support on $\mathcal{A}^{k}_{-i}$. We want to bound $\mathbb{P}_{i,t}(a|H_{t-1})$.
Consider the counterfactual advantage of action $a^{\prime}$ over $a$,
$\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a):=\frac{1}{t-1}\sum_{s=1}^{t-1}U_{i,s}(a^{\prime},a_{-i,s})-U_{i,s}(a,a_{-i,s}),$
and let
$D:\quad\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a)\geq\frac{1}{t-1}\sum_{s=1}^{t-1}[u_{i}(a^{\prime},a_{-i,1})-u_{i}(a,a_{-i,s})]-\frac{\delta}{4}$
denote the event that it does not undercut the expected advantage by more than $\delta/4$. In the following, we will derive our desired bound by splitting the probability $\mathbb{P}_{i,t}(a|H_{t-1})$ into
$\mathbb{P}_{i,t}(a|H_{t-1})=\mathbb{P}_{i,t}(a,\neg D|H_{t-1})+\mathbb{P}_{i,t}(a,D|H_{t-1})\leq\mathbb{P}_{i,t}(\neg D|H_{t-1})+\mathbb{P}_{i,t}(a|D,H_{t-1})$
and bounding the two probability terms.
Let us start with the first term. The random variables $U_{i,s}(a^{\prime},a_{-i,s})-U_{i,s}(a,a_{-i,s})$, $s=1,\dots,t-1$, have expectation $u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})$, are independent, and are bounded by M:=baM:=b-a from above and M=ab-M=a-b from below. By applying Hoeffding’s inequality, we find that
$\mathbb{P}_{i,t}(\neg D|H_{t-1})$ $=\mathbb{P}_{i,t}\left(\sum_{s=1}^{t-1}U_{i,s}(a^{\prime},a_{-i,s})-U_{i,s}(a,a_{-i,s})-\sum_{s=1}^{t-1}[u_{i}(a^{\prime},a_{-i,1})-u_{i}(a,a_{-i,s})]\leq-(t-1)\frac{\delta}{4}\right)$
$\leq\exp\left(-\frac{2(t-1)^{2}\frac{\delta^{2}}{16}}{(t-1)(2M)^{2}}\right)=\exp\left(-\frac{(t-1)\delta^{2}}{32M^{2}}\right)$
Now, we bound the second term. Again, we consider the counterfactual utility advantage of $a^{\prime}$ over $a$. We show that it exceeds $\gamma_{t}$, given that $D$ holds.
$\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a)$ $=\frac{1}{t-1}\sum_{s=1}^{t-1}U_{i,s}(a^{\prime},a_{-i,s})-U_{i,s}(a,a_{-i,s})$
$\geq\frac{1}{t-1}\sum_{s=1}^{t-1}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})-\frac{\delta}{4}$
$=\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\not\in\mathcal{A}_{-i}^{k}\end{subarray}}\underbrace{u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})}_{\geq-\Delta u}+\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\in\mathcal{A}_{-i}^{k}\end{subarray}}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})-\frac{\delta}{4}$
$\geq(1-P_{i,t-1}(\mathcal{A}^{k}))\cdot(-\Delta u)+P_{i,t-1}(\mathcal{A}^{k})\cdot\underbrace{\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a^{\prime},a_{-i})-u_{i}(a,a_{-i})\right]}_{>\delta}-\frac{\delta}{4}$
$>(1-P_{i,t-1}(\mathcal{A}^{k}))(-\Delta u)+\delta P_{i,t-1}(\mathcal{A}^{k})-\frac{\delta}{4}=P_{i,t-1}(\mathcal{A}^{k})(\delta+\Delta u)-\Delta u-\frac{\delta}{4}$
$\geq\tilde{p}(\delta+\Delta u)-\Delta u-\frac{\delta}{4}=\frac{1}{2}(2\Delta u+\delta)-\Delta u-\frac{\delta}{4}=\frac{\delta}{4}>\gamma_{t}$
The second line makes use of the event $D$, which allows us to relate the random variables with their expected outcomes. The fourth line interprets the sum over all times for which the opponents play $a_{-i}\in\mathcal{A}^{k}_{-i}$ as an expectation over a joint action distribution $x_{-i}\in\Delta(\mathcal{A}^{k}_{-i})$ with support on $\mathcal{A}^{k}_{-i}$. We choose $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ such that it performs better than $a$ by a margin of at least $\delta$. For $a\not\in\mathcal{A}^{k+1}_{i}$, this action exists for any such mixed strategy, according to Equation (7). The last line implies that the probability $\mathbb{P}_{i,t}(a)$ that the mean-based algorithm chooses actions $a\not\in\mathcal{A}^{k+1}_{i}$ is small:
$\mathbb{P}_{i,t}(a|H_{t-1},D)\leq\gamma_{t}$
Combining the two bounds shows the result. ∎
Remark 3.
We can replace the assumption of a bounded support for $U_{i,t}$ by the assumption that it has a finite variance. Using Chebyshev’s inequality instead of Hoeffding’s inequality would yield a similar result that would still show the main claim.
A.4.7 Extension: Staggered Entry and Biases
In the previous sections, we have assumed that all players start at the same time $t=0$ and have made no experiences about the payoffs in advance. Now, we will show that we can drop this assumption and still show convergence. The intuition behind this is that all beliefs that have been formed up to some point will become negligible if the interaction continues sufficiently long.
Consider the following notation. We assume that all agents will have entered the competition at time $t=1$, meaning that the game will remain stable thereafter (no players leave or join after $t=0$). However, each agent $i$ might have joined the competition at an earlier point $\tau_{i}\leq 1$. We will show that the agents (employing mean-based strategies) still converge to the correlated rationalizable actions of the game that is defined by the competition present at $t=1$.
In this scenario, every agent has $t-\tau_{i}$ observations at time $t$. For $t=\tau_{i},\dots,0$, the agent observes rewards $\tilde{u}_{i,t}(\cdot,\tilde{a}_{-i,s})$ from a utility function that is potentially different from $u_{i,t}$ and that also depends on the potential actions of other competitors $\tilde{a}_{i,t}$ (these can be different competitors than at $t\geq 0$). We will assume that $\tilde{u}_{i,t}$ is bounded by the interval $c,c],c\in\mathbb{R}$. The value estimate of player $i$ for an action $a\in\mathcal{A}_{i}$ at time $t\geq\tau_{i}$ becomes
$\alpha_{i,t}(a)=\frac{1}{t-\tau_{i}}\left(\sum_{s=\tau_{i}}^{0}\tilde{u}_{i,s}(a,\tilde{a}_{-i,s})+\sum_{s=1}^{t-1}u_{i,s}(a,a_{-i,s})\right)=\frac{1}{t-\tau_{i}}\left(\tilde{b}_{i}(a)+\sum_{s=1}^{t-1}u_{i,s}(a,a_{-i,s})\right).$
We have used $\tilde{b}_{i}(a)=\sum_{s=\tau_{i}}^{0}\tilde{u}_{i,s}(a,\tilde{a}_{-i,s})$ to denote player $i$’s biases of the game that they developed during rounds $\tau_{i},\dots,0$. Note that these biases can grow with $\tau_{i}$. These biases could, alternatively, be viewed as prior knowledge that an algorithm designer injects into an algorithm based on their beliefs about the competition.
We again provide an alternative version of Proposition 11 that shows that an individual player will select a correlated rationalizable response as long as its opponents have played a certain set of actions sufficiently often. This is enough to show convergence to the set of correlated rationalizable actions as in Proposition 12.
Recall that we denote the empirical frequency of opponents (i-i) selecting actions in $\mathcal{A}^{k}_{-i}$ up to time $t$ by
$P_{i,t}(\mathcal{A}^{k}):=\frac{1}{t}\sum_{s=1}^{t}\mathbb{I}\left[\forall j\neq i:~a_{j,s}\in\mathcal{A}^{k}_{j}\right].$
Proposition 15.
Given $k\in\{0,\dots,K-1\}$ and $t\geq 1$ such that $\gamma_{(t-\tau_{i})}<\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\frac{\delta}{2}$, where $\delta$ is the competition constant of the game. If $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large for history $H_{t-1}$, then $\mathbb{P}_{i,t}(a|H_{t-1})\leq\gamma_{(t-\tau_{i})}$ for all $a\not\in\mathcal{A}^{k+1}_{i}$.
Remember that, by $\mathbb{P}_{i,t}(a|H_{t-1})$, we denote the probability that player $i$ with $\gamma_{t}$-mean-based algorithm chooses action $a$ at time $t$. This alternative version can still be used with the proofs in section A.4.4: We can replace $\gamma_{t}$ by $\tilde{\gamma}_{t}:=\max_{i\in\mathcal{N}}\gamma_{(t-\tau_{i})}$ in the proof of Proposition 12, and we choose $T_{0}$ sufficiently large such that $\gamma_{(T_{0}-\tau_{i})}<\frac{-2c\lvert\tau_{i}-1\rvert}{T_{0}-\tau_{i}}+\frac{T_{0}-1}{T_{0}-\tau_{i}}$ for all active players $i$.
Proof of Proposition 15..
We say that the empirical frequency $P_{i,t-1}(\mathcal{A}^{k})$ is sufficiently large if
$P_{i,t-1}(\mathcal{A}^{k})\geq\tilde{p}:=\frac{1}{2}\left(\frac{\Delta u}{\delta+\Delta u}+1\right)\in(0,1)$
where $\delta$ is the competition constant of the game and $\Delta u$ is the maximum absolute utility difference between any two actions of player $i$ for any opponent profile $a_{-i}$. Let $a^{\prime}\in\mathcal{A}^{k+1}_{i}$ be an action which increases player $i$’s utility by at least $\delta$ compared to playing action $a\not\in\mathcal{A}^{k+1}_{i}$ if opponents play a mixed strategy with support on $\mathcal{A}^{k}_{-i}$. Consider the counterfactual advantage of action $a^{\prime}$ over $a$:
$\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a)$ $=\frac{1}{t-\tau_{i}}\left(\tilde{b}_{i}(a^{\prime})-\tilde{b}_{i}(a)+\sum_{s=1}^{t-1}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})\right)$
$=\frac{\tilde{b}_{i}(a^{\prime})-\tilde{b}_{i}(a)}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\left(\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\not\in\mathcal{A}_{-i}^{k}\end{subarray}}\underbrace{u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})}_{\geq-\Delta u}+\frac{1}{t-1}\sum_{\begin{subarray}{c}1\leq s\leq t-1:\\ a_{-i,s}\in\mathcal{A}_{-i}^{k}\end{subarray}}u_{i}(a^{\prime},a_{-i,s})-u_{i}(a,a_{-i,s})\right)$
$\geq\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\left((1-P_{i,t-1}(\mathcal{A}^{k}_{-i}))\cdot(-\Delta u)+P_{i,t-1}(\mathcal{A}^{k})\cdot\underbrace{\operatorname{\mathbb{E}}_{a_{-i}\sim x_{-i}}\left[u_{i}(a^{\prime},a_{-i})-u_{i}(a,a_{-i})\right]}_{>\delta}\right)$
$>\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\left((1-P_{i,t-1}(\mathcal{A}^{k}))(-\Delta u)+\delta P_{i,t-1}(\mathcal{A}^{k})\right)=\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\left(P_{i,t-1}(\mathcal{A}^{k})(\delta+\Delta u)-\Delta u\right)$
$\geq\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\left(\tilde{p}(\delta+\Delta u)-\Delta u\right)=\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\left(\frac{1}{2}(2\Delta u+\delta)-\Delta u\right)$
$=\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\frac{\delta}{2}>\gamma_{(t-\tau_{i})}$
In contrast to the proof of Proposition 11, we need to take the biases into account. According to our assumption, we can bind the values for $\tilde{b}_{i}(a)$ and $\tilde{b}_{i}(a^{\prime})$ by c-c and $c$, respectively. We also gain a factor $\frac{t-1}{t-\tau_{i}}$. The series of values $\frac{-2c\lvert\tau_{i}-1\rvert}{t-\tau_{i}}+\frac{t-1}{t-\tau_{i}}\frac{\delta}{2},t\geq 0$ is strictly monotonously increasing and converges to a value of $\frac{\delta}{2}$. As $\gamma_{t}$ is monotonously decreasing and converges to zero, there exists $T\geq 0$ such that $\gamma_{t}$ falls below this bound for all $t\geq T$.
The inequality implies that $\mathbb{P}_{i,t}(a|H_{t-1})\leq\gamma_{(t-\tau_{i})}$, which is less than $\tilde{\gamma}_{t}$. ∎
A.4.8 Formal Theorem
Here, we state the formal version of the informal Theorem Theorem 1 (Informal version) in the main part, which combines all three extensions introduced above. The extensions can be combined into a single statement:
  • We can use the piecewise maximum of $\gamma_{i,t}$ to unify the mean-based bound of all players. This does not affect stochastic utility or staggered entry.
  • The bias terms of staggered entry can be included in the derivation of $\alpha_{i,t}(a^{\prime})-\alpha_{i,t}(a)$ in the proof of Proposition 14. The result, $\tilde{\gamma}_{t}$, is a scaled version of $\gamma_{t}$ with an additional term. However, it still monotonously decreases to zero, which is the only requirement for convergence used in the proof of Proposition 12.
All in all, we obtain the following main result:
Theorem 1.
Let $n$ players, $i\in\mathcal{N}$, compete in a repeated game with stochastic payoffs $U_{i,t}$ such that
$\operatorname{\mathbb{E}}[U_{i,t}(a,a_{-i})]=u_{i}(a,a_{-i}),\qquad U_{i,t}(a,a_{-i})\in[\underline{u},\overline{u}],$
where the randomness is sampled i.i.d. for each pair $(i,t)$. Assume each agent follows a $\gamma_{i,t}$-mean-based algorithm with respect to a monotonically decreasing sequence $\gamma_{i,t}$ such that $\gamma_{i,t}\to 0$ as $t\to\infty$. Then, there exists a rational constant $c>1$ such that
$\mathbb{P}\left(\frac{1}{T_{k}}\sum_{t=1}^{T_{k}}\mathbb{I}\left[\forall i:~a_{i,t}\in\bar{\mathcal{A}}_{i}\right]\geq 1-\beta\right)\geq 1-\sum_{k=1}^{K}\sum_{j=1}^{J}\exp\left(\frac{-(T^{j}_{k}-T^{j-1}_{k})\beta^{2}}{32}\right)$
for every sufficiently small $\beta>0$, where $T_{k}^{j}=c^{j+Jk}T_{0}$ and $T_{k}^{0}=T_{k}$ with a sufficiently large $T_{0}\in\mathbb{N}$ and $J=\lceil\log_{c}\tfrac{2}{\beta}\rceil$.
Moreover, the probability approaches 1 as $T_{0}\to\infty$, indicating time-average convergence to $\bar{\mathcal{A}}=\bar{\mathcal{A}}_{1}\times\dots\times\bar{\mathcal{A}}_{n}$, the set of correlated rationalizable actions of the game with rewards $u_{i}$. The result also holds if some or all of the agents have entered the game at different times $\tau_{i}\leq 1$.
A.4.9 Last-iterate convergence
Finally, we also provide the proof for Corollary 1.
Proof of Corollary 1..
From Lemma 3, we get that $P_{t-1}(\mathcal{A}^{K-1})\geq\tilde{p}$ for sufficiently large $t$. According to Proposition 11 (Prop. 14, 15), a player $i$ plays outside of $\bar{\mathcal{A}}=\mathcal{A}^{K}$ with probability at most $\gamma_{t}$ ($\tilde{\gamma}_{t}$). This means that the player plays inside $\bar{\mathcal{A}}$ with probability at least $1-\gamma_{t}$ ($1-\tilde{\gamma}_{t}$). As $t$ approaches infinity, this probability approaches 1, showing that the player converges to rationalizable actions in last-iterate, which corresponds, by assumption, to the action of the Nash equilibrium $a^{*}$. The same statement holds for all players. ∎
Appendix B Additional Experiments
In this section, we provide a brief description of our algorithms and provide additional experiments that shed further light on the algorithmic interaction in oligopoly games.
B.1 Online Optimization Algorithms
B.1.1 $\epsilon$-Greedy Bandit
One of the simplest algorithms is the $\epsilon$-greedy exploration strategy. With its internal belief, it keeps track of the empirical mean reward $\bar{u}_{t}(a)$ of each explored action $a$. With probability $1-\epsilon$, it selects the most promising action in the current time step, assuming that each action has been chosen at least once. With probability $\epsilon$, it chooses an arbitrary action with uniform probability.
$\pi_{t}(a_{t})=\begin{cases}1-\epsilon\cdot\frac{K-1}{K}&a_{t}=\arg\max_{a\in\mathcal{A}}\bar{u}_{t-1}(a)\\ \frac{\epsilon}{K}&\text{else}\end{cases}$ (15)
B.1.2 UCB Algorithms
Like $\epsilon$-greedy, the UCB algorithms estimate and update the empirical mean reward of all their actions. They balance exploration and exploitation by choosing arms according to a heuristic, the upper confidence bound (UCB) of estimated rewards. This value can be computed by adding an exploration bonus $\rho_{t}$ to the empirical mean of past rewards $\bar{u}_{t}$:
$\text{UCB}_{t}(a)=\bar{u}_{t}(a)+\rho_{t}(a)$ (16)
The exploration bonus may depend on past rewards, $u_{s}(a_{s}),~s\leq t$, the number of times an action was selected, $n_{t}(a)$, and the time, $t$. In each step, the algorithm chooses the action $a_{t}$ which maximizes $\text{UCB}_{t-1}(a)$. Similar to hansen_frontiers_2021, we use the UCB-Tuned algorithm (auer2002bandit) in our experiments. This algorithm is a refined version of UCB1 (auer2002bandit) with an exploration bonus
$\rho_{t}^{UCB-Tuned}(a)=\sqrt{\frac{\ln n}{n_{t}(a)}\min\left\{\dfrac{1}{4},V_{t}(a)\right\}}$
where $n$ is the number of plays so far and $V_{t}(a)$ is a variance estimate
$V_{t}(a)=\left(\frac{1}{n_{t}(a)}\sum_{\begin{subarray}{c}1\leq s\leq t:\\ a_{s}=a\end{subarray}}u_{s}(a)^{2}\right)-\bar{u}_{t}(a)^{2}+\sqrt{\frac{2\ln t}{n_{t}(a)}}.$
Unlike hansen_frontiers_2021, our implementation of the algorithm does not eliminate any actions that it deems unpromising, and it resolves ties of the UCB heuristic randomly. The algorithm has vanishing regret in the stochastic multiarm bandit problem under mild assumptions (see, e.g., lattimore2020bandit) but may suffer linear regret in the adversarial setting.
B.1.3 Exp3 Algorithms
Exp3 (”Exponential-weight algorithm for Exploration and Exploitation”) is an algorithm introduced by auer_gambling_1995, which was designed for the adversarial model. This means that we do not make any assumptions about the distribution of rewards. In contrast, the rewards could also be chosen by an adversary who has full information about the setting and the agent’s algorithm and who chooses rewards in each round.
The Exp3 algorithm’s action-selection rule assigns weights to each arm, and its update rule adjusts these weights based on the observed rewards. Following the slightly modified version by lattimore2020bandit, our implementation of the Exp3 algorithm samples its actions $a_{t}$ from
$P_{t}(a)=(1-\epsilon_{t})\cdot\frac{\exp\left(\eta\sum_{s=1}^{t-1}\hat{u}_{s}(a)\right)}{\sum_{\hat{a}\in\mathcal{A}}\exp\left(\eta\sum_{s=1}^{t-1}\hat{u}_{s}(\hat{a})\right)}+\frac{\epsilon_{t}}{\lvert\mathcal{A}\rvert}$. (17)
Here, $\epsilon_{t}$ is an additional exploration rate that encourages the algorithm to choose arbitrary actions from time to time and $\hat{u}_{s}(a)$ is an estimate of the reward. Let $P_{s}(a)$ be the probability that the algorithm chooses action $a$ at time $s$, given the history until then. Assuming rewards are bounded within $[0,1]$, we choose
$\hat{u}_{s}(a)=1-\frac{\mathbb{I}\{a_{s}=a\}}{P_{s}(a)}\cdot(1-u_{s}).$ (18)
The Exp3 algorithm is known to be no-regret for some choices of $\epsilon_{t}$ in the adversarial setting (auer_gambling_1995, lattimore2020bandit) and mean-based for a specific choice of exploration rate $\epsilon_{t}$ and learning rate $\eta_{t}$ [braverman_selling_2018]. Our implementation relies on a fixed exploration and learning rate for simplicity.
B.1.4 Thompson Sampling
Thompson Sampling is a Bayesian approach to multi-armed bandit algorithms that was named after William R. Thompson based on his papers from thompson_likelihood_1933 and thompson_theory_1935. The heuristic maintains probability distributions over the rewards of every possible action and uses these to determine the next action.
More specifically, Thompson sampling assumes that the rewards of each action $a\in\mathcal{A}$ are sampled independently from a probability distribution $\mathbb{P}(u_{t}=r|\theta_{a})$. [^5] This also generalizes to probability densities. For simplicity, we stick to the discrete formulation in this introduction. Bayes’ rule allows us to maintain a posterior distribution to update our beliefs on the (unknown) parameters $\theta_{a}$, and consequently on the distribution of rewards of each action:
$\mathbb{P}(\theta_{a}|\{u_{s}\}_{s=1}^{t},\theta_{a}^{0})=\frac{\mathbb{P}\left(\left.\{u_{s}\}_{1\leq s\leq t:a_{s}=a}\right|\theta_{a}\right)\cdot\mathbb{P}(\theta_{a}|\theta_{a}^{0})}{\mathbb{P}\left(\{u_{s}\}_{1\leq s\leq t:a_{s}=a}\right)}$
Here, $\mathbb{P}(\theta_{a}|\theta_{a}^{0})$ is a prior distribution over reward parameters for action $a$, and $\mathbb{P}\left(\left.\{u_{s}\}_{1\leq s\leq t:a_{s}=a}\right|\theta_{a}\right)$ is the probability of rewards $u_{s}$ at iterations for which the selected action was $a$.
In each iteration, Thompson sampling draws parameter samples $\theta_{a}$ for all actions from their corresponding posterior distribution, introducing randomness. Subsequently, the action with the highest expected reward, based on $\theta_{a}$, is chosen. After receiving reward feedback, the posterior distribution of the chosen action is updated according to Bayes’ rule.
While thompson_likelihood_1933 described the approach for two actions with Bernoulli distributions, our setting has $K$ actions with continuous, bounded rewards. The Bayesian approach can also be used in this setting; a generalization that has been described, e.g., by russo_tutorial_2020. Our implementation of Thompson sampling assumes Gaussian-distributed rewards [^6] We note that this does not reflect that the rewards are bounded. The approach is not severely limited by this. and poses a Gaussian distribution over the parameters as conjugated prior.
B.2 Convergence Speed
Aside from the converged prices, we also analyzed the training behavior of the algorithms. As with the final prices, we find significant differences between the UCB-T algorithms and the other algorithms, which we summarize as our third result.
Result 2.
Exp3-$\epsilon$, TS, and $\epsilon$-Greedy converge quickly and reliably to the Nash equilibrium prices. UCB-T combinations erratically learn to charge higher prices with large differences within and between runs.
In Figures 5 and 6, we provide the price-collusion indices for all four algorithms in symmetric duopolies with linear demand (O2). Indeed, we find that the convergence of Exp3-$\epsilon$ and TS is particularly consistent and quick. The former algorithm sets prices close to the competitive Nash equilibrium after roughly 50,000 steps, while the latter requires even fewer interactions between the agents. With $\epsilon$-Greedy, the agents arrive at competitive prices quickly, with small variations between runs. We attribute this to the fixed exploration rate of our algorithm configuration for $\epsilon$-Greedy.
(a) Exp3-$\epsilon$
(b) TS
Figure 5: Evolution of Prices During Training
The figures display the collusion indices based on the charged prices (running medians over 1000 steps) for ten runs of the experiment in thin lines. The thick line shows the average of all runs and thus indicates the overall trend.
(a) $\epsilon$-Greedy
(b) UCB-T
Figure 6: Evolution of Prices During Training
The figures display the collusion indices based on the charged prices (running medians over 1000 steps) for ten runs of the experiment in thin lines. The thick line shows the average of all runs and thus indicates the overall trend.
In stark contrast to this, the UCB-T experiments exhibit a larger variation of prices, both within runs and between runs. The variation within runs indicates that agents do not simply ”agree” on a supra-competitive price level. Instead, they erratically raise and lower prices in a seemingly random pattern. Since these agents choose actions by maximizing their heuristic (which is a non-smooth operation), their actions can vary drastically if multiple actions experience values of the heuristics at around the same level. Each run of our UCB-T experiments is different, and we observed a strong variation between the runs. We note that the only source of randomness comes from the random action selection if one or more actions have the same heuristic value. The varying outcomes make it hard for firms to reliably predict their returns when employing a UCB-T algorithm in this scenario. On average, though, we observe collusion, starting from an early point in the training.
B.3 Algorithmic Collusion with More Sellers
There may be more than two competitors in a market. There are two major questions we answer in this context: 1) Do colluding algorithms such as UCB-T eventually reach competitive prices if the number of competitors grows large enough? 2) How many non-colluding agents must be inserted in a system of colluding agents to break supra-competitive pricing? Our second result sheds light on these questions.
Result 3.
With an increasing number of competitors, collusion indices decrease. Pure UCB-T oligopolies can still achieve supra-competitive prices, but at a lower level. Adding other bandit algorithms reduces algorithmic collusion.
To answer the first question, we ran experiments with identical algorithms in the linear demand setting (O2). Figure 7(a) shows that with an increasing number of prices, UCB-T comes closer to the Nash equilibrium prices. However, price levels remain collusive in oligopolies of up to ten players. In line with our previous experiments, the other algorithms remain competitive for any number of players.
We answer the second question by configuring oligopolies where all players commit to the UCB-T algorithm except for one. The results are displayed in Figure 7(b). While the collusion levels decrease compared to the pure UCB-T oligopolies, the UCB-T algorithms are still able to coordinate on supra-competitive prices.
(a) Identical algorithms
(b) UCB-T majority
Figure 7: Profit Collusion with Increasing Number of Players
The plots visualize the price-collusion indices at the end of training for different numbers of players in a Bertrand economy with linear demand. a) shows oligopolies with identical algorithms while b) shows the combination of n1n-1 UCB-T algorithms with one different algorithm. The mean profit of the last 1000 iteration steps is averaged over all runs to form the index.
Of course, this coordination is only possible if there are sufficiently many players to support it. We analyze the competition of varying shares of UCB-T agents in a 10-player oligopoly with competing Exp3-$\epsilon$ algorithms with linear demand (O2). The result is displayed in Figure 8. We only observe collusion if the majority of agents play UCB-T. With increasing numbers, collusion increases, too.
Figure 8: Different Numbers of UCB-T Agents Against Exp3-$\epsilon$ Competitors in a 10-player Oligopoly
The plot visualizes the price-collusion index in 10-player oligopoly experiments with linear demand (O2) for an increasing share of UCB-T players. The remaining agents act according to Exp3-$\epsilon$. The mean profit of the last 1000 iteration steps is averaged over all runs to form the index.
B.4 Staggered Entry
We may not always assume that competing agents start at the same time. In this section, we complement the analytical result of Section A.4.7 with experiments. Our modified setup is as follows: We simulate the environment for 250,000 steps. In the first 20% of all iterations, only one player is in the market. Then, the second contestant enters. A logit-demand model is most appropriate here as it ensures that the single-player version of the environment exhibits bounded, positive demand. The parameters are selected as in the symmetric logit-demand duopoly described in Section 5.2.
We ran simulations with Exp3 and Thompson Sampling (TS) for 250,000 steps. The results are displayed in Figure 9. In both scenarios, the first competitor learns to charge a high price in the beginning. As soon as the second seller enters the market, the first seller immediately starts lowering prices and adapts to the increased competition. However, convergence is slow, and none of the scenarios fully recovers the competitive price within the given time frame.
(a) Exp3-$\epsilon$
(b) TS
Figure 9: Staggered-entry experiments where the second contestant entered the market at one-fifth of the iterations.
Displayed is the action averaged over ten independent runs and 1000 subsequent steps. The first contestant is displayed in black, the second entrant in red. The light grey curve represents a simultaneous-entry version of the scenario for reference. The (continuous) Nash equilibrium is located at approximately 1.47; the joint profit-maximizing prices and the monopoly price are roughly 1.92. The discrete Nash equilibrium and monopoly price are marked by dashed lines.
Depending on how long the entry of the second entrant is delayed, the two agents might not recover their equilibrium outcome due to a fixation of the first contestant. While they are a monopolist, the first player might start to reject low actions as these don’t yield a high reward in a monopolistic market. We could observe that the probability assigned to such actions becomes very small, close enough to be considered virtually equivalent to zero from a computer’s perspective. This could lead to shrinking flexibility and thus prohibit competitive play in some scenarios.
B.5 Random Coefficients Models: The Cereal Market
The examples in the main part of our paper are based on toy scenarios with a few agents and simple market structures. While they are a good choice for analyzing the algorithms’ behavior in an isolated environment, they lack the complexity of real-world markets. We investigate an additional market scenario which is based on the more complex random coefficients model introduced by berry_automobile_1993. In particular, we build on the work by nevo_practitioners_2000, nevo_measuring_2001, which analyzes the prices in the American Cereal Market in the years 1988 to 1992. This market seems particularly suited to our case as it exhibits high margins, raising the question of whether or not real firms compete as expected [nevo_measuring_2001]. We leverage the PyBLP package [conlon_best_2020, conlon_pyblp_2025], which provides dummy data for the cereal market (as used in nevo_practitioners_2000, conlon_best_2020) and code for parameter estimation and counterfactual analysis. Our random coefficients model follows the description provided in the PyBLP documentation.
The data comprises 94 markets, each characterized by a combination of city and quarter. Every market consists of five firms, each holding one to nine products with individual prices. Further increasing the complexity is the interplay between products. Due to the random coefficients model, products show more intricate substitution patterns and price elasticities. Finally, firms with multiple products can coordinate their prices to maximize their aggregated profit.
Algorithm 2 Code parts for random coefficients experiments
# Create the random coefficients model
problem = pyblp.Problem(
product_formulations=(
pyblp.Formulation("0 + prices", absorb="C(product_ids)"),
pyblp.Formulation("1 + prices + sugar + mushy"),
agent_formulation=pyblp.Formulation(
"0 + income + age + child"
), # Demographics
product_data=pd.read_csv(
pyblp.data.NEVO_PRODUCTS_LOCATION
agent_data=pd.read_csv(
pyblp.data.NEVO_AGENTS_LOCATION
# Estimate the parameters
results = problem.solve(
sigma=np.diag([0.3302, 2.4526, 0.0163, 0.2441]),
pi=[
[5.4819, 0, 0 ],
[15.8935, -1.2000, 2.6342 ],
[-0.2506, 0, 0 ],
[1.2650, 0, 0 ],
method="1s", # One-step GMM
optimization=pyblp.Optimization(
"bfgs", {"gtol": 1e-5}
# Compute market shares and profits for varying prices in a specific market
market_id = "C01Q1"
prices = np.array([…])
shares = results.compute_shares(prices=prices, market_id=market_id)
profits = results.compute_profits(
prices=prices, shares=shares, market_id=market_id
We instantiate the model, estimate its parameters based on the full dataset, and compute market shares for prices in individual markets. In essence, the code is as described in Algorithm 2.777See also https://pyblp.readthedocs.io/en/stable/_notebooks/tutorial/nevo.html. For our market simulations, every product is represented by one agent that learns to set optimal prices for this product. Eligible cooperation between the products of one firm can be implemented by sharing the reward between the respective agents (see below).
The experiments are mostly configured as described in the main part of our paper. However, there are three specific parameters that we need to choose.
  • Market ID: We can choose from a set of 94 markets, each with a unique combination of city and quarter, that all comprise 24 products. We arbitrarily select ”C01Q1” and ”C07Q2” as our choices.
  • Agents: For our experiments, we employ Exp3 and Thompson sampling.
  • Profit model: As described above, the market consists of multiple firms with one to five products each. We can either let each agent observe the aggregate profit of all products of the corresponding firm (profit model ”firm”) or let every agent only receive the profit of its own product (profit model ”product”).
The agents can choose from 21 actions between 0 and 0.25, which are prices that extend the equilibrium and monopoly prices of the continuous game. The equilibrium prices were computed by PyBLP. For the monopoly prices, we simulated a merger of all firms and evaluated the resulting optimal prices with PyBLP. We scaled the rewards based on the profits with equilibrium and monopoly prices to ensure that our algorithms’ default parameters performed reasonable. [^8] This does not change the strategic nature of the game. We ran our simulations for 500,000 steps and report the price collusion indices over time. Note that the price collusion index is now computed based on the reference prices of the continuous game, as the estimation of the discrete equilibria is out of reach for 24 players. All simulations have been repeated 5 times (with seeds 0 to 4).
The results are visualized in Figure 10 and Figure 11. In the first configuration, we observe a steep price increase in the very beginning, followed by a steady decline in prices. This indicates a consistent behavior with previous results in smaller markets. However, we also observe that convergence is much slower in market C07Q2. In general, slow convergence is to be expected as the large number of players and the complexity of the interaction make the feedback noisy and impede learning.
With Thompson sampling, our results are less promising. Interestingly, we observe a split between runs: Half of our runs resulted in high collusion indices, while the other half was much more benign. A potential reason is, again, slow convergence due to complexity. In market C07Q2 (not displayed), the convergence of Thompson sampling was slightly better, resulting in an average collusion index of about 0.25.
Lastly, the alternative profit model ”product”, where every agent is evaluated on the profit of its product alone, shows comparably quick convergence to competitive prices. Keep in mind that the reference price is still based on the assumption that a firm coordinates its prices, which means that collusion index values below zero can occur easily with this non-cooperative profit model.
(a) C01Q1, Exp3-$\epsilon$, ”firm”
(b) C07Q2, Exp3-$\epsilon$, ”firm”
Figure 10: Price collusion indices in the random coefficients model.
We report the results for markets C01Q1 and C07Q2. The collusion indices are based on median price of the previous 1000 steps and are averaged over 24 agents and five runs. The results of the individual runs are displayed in light gray.
(a) C01Q1, TS, ”firm”
(b) C01Q1, Exp3-$\epsilon$, ”product”
Figure 11: Price collusion indices in the random coefficients model.
We report the results for algorithm ”TS”, and profit models ”product”. The collusion indices are based on median price of the previous 1000 steps and are averaged over 24 agents and five runs. The results of the individual runs are displayed in light gray.
B.6 Bertrand Duopoly Experiments with More Algorithms
We also ran experiments with another variant of the UCB algorithms described by auer2002bandit, UCB1, and with Q-Learning (QL). Following Calvano.2020, our implementation of QL uses an $\epsilon$-greedy exploration strategy with an exponentially decaying exploration rate $\epsilon$. A small exploration rate towards the end seems important for the evolution of collusion with Q-Learning [barfuss_intrinsic_2023, schaefer_emergence_2023, lambin2024less]. For all experiments where QL was involved, we iterated 3 Mio. steps (instead of 250,000), and we provided the actions of both players from the previous iteration as the state. For completeness, we also repeated our experiments with pretrained Q-Learning agents, but could not see substantial differences in the final outcome.
Figure 12 shows the price-collusion indices of all algorithm configurations. We observe that, in line with what we observed in our main experiments, collusion mostly evolves when two algorithms of the same type are paired. The combination of UCB-T and UCB1 deviates from this pattern and displays a significant level of collusion. It is likely that these two algorithms can learn to ”coordinate” (unconsciously) because of their similarity, which shows that algorithms need not share the exact same implementation and parameters in order to collude.
Figure 12: Price-collusion indices in duopoly O3 (logit demand)
We display the price-collusion index by agent as an average value over all runs. To determine the index, we use the median price over the last 1000 iterations.
For the Q-learning experiment, figure 13 shows the evolution of collusion over the experiment iterations. We observe that collusion evolves late and abruptly, which can likely be attributed to the exploration rate becoming sufficiently small.
Figure 13: Evolution of collusion with QL
We display the price-collusion index of of a QL-duopoly over time. To determine the index, we use the median price over 1000 iterations.
B.7 Collusion Results by Demand Function
In the main part of the paper, we averaged the collusion indices of the duopolies for all demand functions. Here, we provide the results for the individual demand functions. Figures 14 and 15 display the price-collusion indices for all five games. We find that supra-competitive prices evolve consistently with the UCB-T algorithm throughout all settings, with particularly pronounced collusion in the standard demand setting (O1). Likewise, prices remain mostly competitive for the remaining algorithm combinations.
(a) O1.
(b) O2.
(c) O3.
Figure 14: Price-collusion indices in symmetric games
We visualize the price-collusion indices for the symmetric games and several combinations of algorithms. The displayed numbers are the numeric values of the collusion indices; the values in brackets display their corresponding standard deviation. Averages and standard deviations were computed over two agents and 10 runs.
(a) O2’.
(b) O3’.
Figure 15: Price-collusion indices in asymmetric games
We visualize the price-collusion indices for the asymmetric games and several combinations of algorithms. The displayed numbers are the numeric values of the collusion indices; the values in brackets display their corresponding standard deviation. Averages and standard deviations were computed over two agents and 10 runs.